### v1

# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

Abstract
We present a low-cost quadruped manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning (RL) purely in simulation. The system comprises 1) a hierarchical design of a high-level policy for visual-mobile manipulation following instructions, and a low-level policy for quadruped movement and limb-control, 2) a progressive policy expansion approach for solving the long-horizon task together with a teacher-student framework for efficient high-level training of the high-level visuomotor policy, and 3) a suite of techniques for minimizing sim-to-real gaps.
   With budget-friendly but limited reliability and performance hardware, and just one wrist-mounted RGB camera, the entire system fully trained in simulation achieves high success rates for long horizon tasks involving search, move, grasp, and drop-into, with fluid sim-to-real transfer in a wide variety of indoor and outdoor scenes and lighting conditions.Extensive real-world evaluations show that on the long horizon mobile manipulation tasks, our system achieves good performance when transferred to real both in terms of task success rate and execution efficiency. Finally, we discuss the necessity of our sim-to-real techniques for legged mobile manipulation, and show their ablation performance.
Figure 1: SLIM in Real. Snapshots of the same SLIM policy running in different scenes in the real world, with large scene variations, including ground properties, backgrounds, and distractors etc. These scenes are not available in simulation training. The annotations at the bottom are added for understanding task progress stages and are not part of the input to the system.
I Introduction
Until recently, robots have primarily been limited to static, highly repetitive tasks, often in industrial manufacturing. As the demand for autonomy grows, the need emerges for robots to navigate and interact with their environment – commonly referred to as mobile manipulators. Such systems typically handle primitive tasks, such as searching, grasping, and transporting objects, to achieve overarching long-horizon goals.
Several promising directions exist in training autonomous mobile manipulators: imitation learning from human demonstrations, reinforcement learning (RL) in the real world, and RL in simulation with sim-to-real transfer. Each of the directions has its pros and cons.
Imitation learning benefits from learning directly from human experts teleoperating the hardware, thus, reasonable behavior and some successes can be achieved even with a small amount of training. However, it can be expensive to collect demonstrations to cover all the combinatorial variations of the scene, the task setup, and the hardware. Reinforcement learning (RL) can automate policy optimization so minimal human effort is needed to adapt to the environment and task variations. Furthermore, the resulting policy is often more efficient than human teleoperation. However, mastering long-horizon tasks in the real world still poses a significant challenge. Directly training an RL policy in the real world can be challenging due to hardware constraints and safety issues. Having to rely on real robots, training efficiency can be low, and scalability is often limited. Additionally, a model trained in one physical setup may not generalize to another with for example different background or floor texture.
RL in simulation solves the scalability and adaptability issues by running massively parallel simulations with randomized setups. The difficulty lies in the sim-to-real gap. Shortcomings of the design often only shows up in real world testing, at the very end of the development cycle. Making things worse, several iterations of modeling, design tweaking, policy training, and real testing are often needed to close up a particular sim-to-real gap. Given such a long and often expensive pipeline of tuning steps, RL in simulation can be slow to arrive at a working prototype, and can be expensive in terms of engineering costs and server time for simulation and training. Therefore, the hope is to develop general techniques to bridge the sim-to-real gaps, so that efforts spent on developing the training pipeline and the robotic system for some tasks can be reused for training future tasks. This work focuses on training an end-to-end RL policy in simulation, and zero-shot deploying in the real world on a relatively low-cost robotic system to solve long-horizon, intricate quadruped manipulation tasks. Weak and often faulty hardware, the complexity of legged manipulation, and the long horizon are some of the factors that would compound and significantly expand the sim-to-real gap, making this a daunting task.
To the best of our knowledge, this is the first working solution of a legged manipulator fully trained with RL in simulation, for long horizon tasks, with low cost hardware.
We use a quadruped instead of a wheel base, for it can adapt to different terrains, but manipulation is also more challenging on legs, due to the unstable legged base.
This results in a large 19 DOF (12 leg joints, 6 arm joints, and 1 gripper) system, exacerbating the already difficult long-horizon RL exploration issue. To mitigate this, we utilize a hierarchical design consisting of a high-level mobile manipulation policy and low-level quadruped controller, and also rely on a teacher-student training framework similar to prior work [29, 59].
The teacher policy has access to privileged information (e.g., structured and accurate object features) that is unavailable in the real world. To tackle the difficulties of long-horizon task learning, we also incorporate task decomposition as a form of privileged information, where boundaries between subtasks are defined. The student policy is then trained to distill the teacher policy while maximizing task rewards, but with access to only standard sensor data such as camera feed and proprioceptive state. We also condition the student visuomotor policy on language instructions, allowing us to naturally command the robot to achieve different tasks.
We name our robotic system SLIM (Sim-to-Real Legged Instructive Manipulation). We reiterate that SLIM is a complete robotic system with fluid sim-to-real transfer. Unlike prior work [29, 52], SLIM learns all modules in simulation from scratch, without the need for any pretrained model or real-world finetuning. SLIM relies on pure RL, and thus it does not require a single demonstration trajectory. Our contributions are as follows:
  1. 1.
    We develop a low-cost quadruped manipulation system SLIM. To our best knowledge, SLIM is the first end-to-end system for solving long-horizon real-world mobile manipulation tasks from RL alone.
  2. 2.
    To address the long-horizon RL exploration challenge, we propose a hierarchical framework for training SLIM’s high-level mobile manipulation policy and low-level quadruped controller, and adopt a teacher-student setup for high-level training efficiency.
  3. 3.
    We identify several techniques that are crucial for SLIM to achieve successful sim-to-real transfer with inexpensive hardware, as shown in real-world ablation studies.
  4. 4.
    We conduct an extensive set of simulated and real-world experiments, achieving higher success rate and faster task completion than baselines. In the real world, we demonstrate robust task success across a wide range of indoor and outdoor environments.
II Related Work
II-A Learning-based Quadruped Locomotion
Traditionally, quadruped locomotion has been tackled through classical control methods designed to follow hand-tuned gaits on flat ground [16, 8], dynamic rigid platforms [23], discrete terrain [2], and rough terrain [11]. Though impressive, such methods require significant human engineering efforts and can be brittle to environmental changes.
More recently, there has been explosive progress in using learning-based approaches for achieving quadruped locomotion. In particular, sim-to-real reinforcement learning (RL) has arisen as a robust solution showcasing impressive feats such as parkour [60, 7, 4] and speeds of up to 3.9m/s [34]. Other works have also focused on exploiting the strengths of legged locomotion over their wheeled counterparts through utilizing the strong inherent coupling of proprioception with egocentric vision [15, 1] and active estimation of the environment [36, 27, 30]. Furthermore, minimal tuning is necessary compared to classical approaches [13] with massively parallelized simulations having been shown capable of learning gaits in just minutes [42]. Finally, data-driven legged locomotion has shown great promise in its ability to zero-shot generalize to new morphologies [12], surfaces [35, 27, 30] and agile skills [49, 26].
II-B Legged Mobile Manipulation
Significant advancements in robust locomotion have further enabled researchers to push the boundaries of legged mobile manipulation. Building upon data-driven approaches, quadrupeds have been demonstrated pushing objects with their body [25], dribble around balls [26], and manipulate objects using egocentric [48] or calve mounted grippers [28]. Still, the most popular setup has been the traditional top-mounted 6DOF manipulator design [14, 17, 29, 58, 52, 33, 37, 44]. In addition to significant workplace expansion, manipulators offer other nuanced benefits, such as assisting with balance [21] and serving as an intuitive interface for collecting human demonstrations [17]. Classical approaches to solving top-mounted legged mobile manipulation consist of designing wholebody controllers [44] or even combining locomotion learning with model-based manipulation control [33]. Others simplify the need for low-level controllers by using high-level API calls ($\dot{x},\dot{\theta}$) provided by their respective quadruped robots to achieve graspability-aware policies [58], navigational pick-and-place [52], and language-conditioned mobile manipulation [39]. Complex tasks have also been achieved by leveraging teacher-student training setups in simulation for grasping [29] and door opening [59]. Finally, to expand the robot’s workspace even further, wholebody loco-manipulation has arisen where leg joints are actuated in a way to assist manipulation [14, 29, 37].
Most similar to our work, Liu et al. [29] introduced VBC, which trains a sim-to-real visuomotor whole-body loco-manipulation policy using a high-low hierarchical model and teacher-student training setup for the object pickup task. They achieve impressive results across a diverse range of objects by leveraging whole-body manipulation, though this approach comes with several key caveats. First, VBC requires the user to manually click on the target object to perform an initial segmentation using a third-party general-purpose model. This segmentation is then used to track the object through another third-party model. Consequently, to ensure consistent segmentation, the object must always remain within the frame. This requirement inherently limits the feasible starting state distribution of VBC, as the target object must be placed within a short distance in front of the robot and in view. Finally, due to the close proximity and visibility of the object, the VBC task horizon is relatively short, ending as soon as the object is lifted. In comparison,
  1. 1.
    SLIM is a complete and self-contained system. All modules–vision, high-level visual loco-manipulation, low-level motor control–are trained within a single framework, resulting in both minimal latency and full autonomy.
  2. 2.
    SLIM is fluid and intuitive. Conditioned on language, SLIM is able to rapidly accomplish tasks in a wide variety of scenes, requiring just a user language command.
  3. 3.
    Finally, SLIM is practical. With SLIM, we demonstrate impressive real-world, long-horizon task completion for a multi-stage mobile search and manipulation task.
III Architecture Overview
[…]
Figure 2: Hierarchical Framework and Visuomotor Policy Pipeline (Left). Given a language instruction and various sensor information, the policy generates two control signals: 1) arm joint position targets and 2) a locomotion command. The first control signal is directly passed to the arm driver (passthrough), while the second signal is passed into a low-level locomotion policy to generate leg joint position targets. Low-Level and High-Level Teacher-Student Policy Training (Right). SLIM’s training pipeline can be decomposed into the following three sequential stages of policy training. First, a low-level locomotion policy is trained to track a specified locomotion command via RL. Next, a high-level teacher policy is trained with privileged information via RL to solve the long-horizon task. Finally, the student visuomotor policy is trained by distilling the teacher policy while maximizing task rewards, conditioned on visual inputs. Both the teacher and student policies use the same frozen locomotion policy from the first stage to execute their locomotion commands.
III-A Observations and Actions
Observations.
The robot observation at each time step consists of three components:
  • $\mathbf{l}$
    A tokenized instruction vector of length $L$.
  • $\mathbf{o}$
    A temporal stack of RGB images from a single wrist-mounted camera with a shape of $(N,H,W,3)$, where $N$ is the stack size, and $H$ and $W$ are the image height and width, respectively. The environment is always partially observed by the robot.
  • $\mathbf{s}_{\textrm{r}}$
    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of all joint position readings.
Privileged Observations (only available in simulation).
To achieve better learning efficiency, we will first train another policy using privileged observations, and then distill it to our robot policy conditioned on the standard observations. This guidance policy has a subtask id input $k$ in place of $\mathbf{l}$ and a lower-dimensional object observation $\mathbf{s}_{\textrm{p}}$ in place of $\mathbf{o}$:
  • $k$
    An integer in $[1,K]$ indicating which subtask the teacher is currently solving. $K$ is the maximal number of subtasks forming a long-horizon task.
  • $\mathbf{s}_{\textrm{p}}$
    A group of temporal object feature stacks, where each stack has a shape of $(N,Q_{m})$ and $0\leq m<M$ denotes the feature index. $M$ is the maximal number of task related objects in the scene. Note that object features are also partially observed and their visibility is always determined by the robot’s camera field of view.
The privileged information can be extracted or computed from the underlying simulator state.
Actions.
The robot policy outputs a triplet of actions:
  • $\mathbf{a}_{\textrm{arm}}$
    A vector in $[-z,z]^{6}$, indicating the delta changes to the arm joint positions. $z$ is the maximally allowed joint position change for a control interval.
  • $g$
    A target gripper position in $[0,1]$.
  • $\mathbf{a}_{\textrm{leg}}$
    The target quadruped joint positions in $\mathbb{R}^{12}$.
In total, a complete action $\mathbf{a}$ has 19 dimensions.
III-B Hierarchical Policy Structure
The system as a whole takes visual and instructional input and controls all joints of the quadruped and the arm, which is a daunting training task with high-dimensional input and output. We use a two-level hierarchical policy to divide the complexity of visual mobile manipulation training from legged locomotion training, similar to [29], illustrated in the left half of Fig. 2.
The high level takes visual, instructional, and proprioceptive inputs, and outputs a simple locomotion speed command and arm control command, while the low level takes proprioceptive input and outputs leg joint commands to track the locomotion command. The assumption is that quadruped locomotion control is largely independent of the high-level task semantics.
Intermediary Action.
To enable this two-level policy structure, we introduce an intermediary action as follows,
  • $\mathbf{c}$
    A vector in $\mathbb{R}^{2}$, containing the target forward and angular velocities for the quadruped.
With this, a high-level policy is defined as
$(\mathbf{l},\mathbf{o},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a}_{\text{hi}}\triangleq(\mathbf{a_{\textrm{arm}}},g,\mathbf{c})\in\mathbb{R}^{9},$
while the low-level policy is defined as
$(\mathbf{c},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a_{\textrm{leg}}}\in\mathbb{R}^{12}.$
Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$, and has a much shorter time horizon.
Low-Level.
The low-level policy is a quadruped controller that generates joint position targets for PD control to follow a task-agnostic 2D command $\mathbf{c}$ [35]. The low-level policy is trained in simulation using PPO [43] and transferred to real. An illustration of the training can be found in the first block in the right column of Fig. 2. It is worth noting that we randomly sample both command $\mathbf{c}$ and arm joint actions $\mathbf{a_{\textrm{arm}}}$ during low-level training. Since the arm mounted on the quadruped can have varying poses during the task, the low-level command following ability must generalize to various arm configurations. That is, the quadruped base has to learn to keep balance while achieving locomotion commands regardless of the arm’s current joint positions and movements. Similar to most sim2real approaches, we also randomize the simulation environment during training. Besides widely-used domain randomization parameters as in [14], we further randomize the simulated delay of each sensor and foot softness of the robot, to better adapt to variations in real deployment. Rewards design and full details of the training setup can be found in Appendix A-B.
High-Level.
The high-level policy is responsible for outputting the command $\mathbf{c}$ and the arm control $(\mathbf{a_{\textrm{arm}}},g)$, conditioned on language $\mathbf{l}$ and the current stack of RGB images $\mathbf{o}$. As demonstrated in Fig. 2, the command $\mathbf{c}$ is forwarded to the low-level policy which then follows the received command for a number of time steps, before the high-level policy outputs the next command. For the manipulator, we opt to operate in joint space (as opposed to task space) and thus no inverse kinematics is needed. Once the low-level policy is trained, it is frozen and used as a base controller by the high-level policy. In the remainder of this paper, we will only talk about the high-level policy unless otherwise stated.
III-C Teacher-Student Training
Given the challenges of an enormous exploration space implied by long-horizon tasks, we adopt a teacher-student learning framework  [10, 29, 59] for better high-level training efficiency. The overall structure of the framework is shown in the second and third blocks of the right side of Fig. 2. As shown, the teacher is trained purely with RL from privileged, structured, and low-dimensional environment data $(k,\mathbf{s_{\textrm{p}}},\mathbf{s_{\textrm{r}}})$ that can only be obtained from the simulator. After a high performing teacher policy is learned, it is frozen and used to guide the student’s policy learning via a distillation loss. As the student will eventually be deployed in the real world, its inputs are consistent with what our robot system has. Besides policy distillation, we also use an RL objective to allow the student’s behavior being shaped by the same set of task rewards used by the teacher. Later, we will show improvements of this distillation-guided RL setting, compared to pure policy distillation (Section VIII-B Distillation Only baseline.). In the next two sections, we will provide more technical details of the teacher and student.
IV The Teacher: Progressive Policy Expansion for Long-Horizon Task Learning
IV-A Long-Horizon Task Learning Challenges
Since we are targeting at solving long-horizon tasks via RL, we need an effective approach to train a policy to solve long-horizon tasks successfully. There are are few key challenges that must be addressed for solving the long horizon tasks:
  1. 1.
    Continual exploration: When learning a long-horizon task via RL, there may be a number of intermediate milestones that must be sequentially achieved. Therefore, the teacher must be able to keep exploring at its frontier despite solving an intermediate milestone stage, in pursuit of solving the full task. Without the ability of continual exploration, it is more likely to converge to a sub-optimal solution that solves the task only up to some intermediate milestones, therefore cannot solve the full task completely.
  2. 2.
    Loss of plasticity and catastrophic forgetting: As the teacher progresses into a task, it needs to properly cope with both the loss of capacity/plasticity [31, 32] of the network due to the evolving nature of the problem as well as the data imbalance during long-horizon RL task exploration and training, without impacting skills that have already been acquired (catastrophic forgetting).
IV-B Progressive Policy Expansion
Given the challenges listed above, we build upon the idea of Policy Expansion (PEX) [57] originally used in the context of offline-to-online RL. Intuitively, PEX divides the full learning process into stages (e.g. offline and online stages in [57]) and uses dedicated networks for learning within each stage. This way, the trainer allocates new network capacity for continual exploration and learning of new skills without affecting any existing skill (catastrophic forgetting).
Here we generalize this idea for addressing the challenges in long-horizon task learning as mentioned in Section IV-A, by extending it to multiple stages. More concretely, we partition the full long horizon task into a sequence of stages (subtasks), and then perform a policy expansion operation for each new encountered subtask. By doing this, it can effectively combat the challenges in long-horizon task learning, keep exploring beyond the currently solved subtasks to solve the subsequent subtasks, circumvent the loss of plasticity by leveraging the newly allocated policy network  [^1] We use the term policy network to represent all the networks that are required for learning a policy. For example, in the context of Actor-Critic formulation, it encompasses both the actor and critic networks. for learning, and avoid catastrophic forgetting by retaining the policy networks for previous subtasks. Given these progressive multi-stage expansions, we dub this strategy Progressive Policy Expansion.
Formally, we represent the full long-horizon task $\mathcal{T}$ as a composition of a set of subtasks
$\mathcal{T}=\mathcal{C}(\{\tau^{k}\}_{k=1}^{K})$ (1)
where $\mathcal{C}$ is a composition operator, assembling the set of subtasks $\{\tau^{k}\}$ into the full task $\mathcal{T}$. $K$ denotes the maximum number of subtasks. We also define a task decomposional operator as follows, mapping an input state to the corresponding subtask index:
$k=\mathcal{D}(\mathbf{s})\quad k\in\{1,2,\cdots,K\},$ (2)
where $\mathbf{s}=[\mathbf{s}_{\textrm{r}},\mathbf{s}_{\textrm{p}}]$. We use $\pi_{\rm tea}^{k}(\mathbf{a}|\mathbf{s})$ to denote the teacher sub-network that is responsible for solving its corresponding subtask $\tau^{k}$. When there is no confusion, we omit the subscript of the teacher network $\pi_{\rm tea}^{k}$ as $\pi^{k}$. Figure 3 provides graphical illustrations of the teacher sub-network structures. Each sub-network (Figure 3 left) takes low-dimensional proprioceptive and privileged observation as input and is responsible for learning to solve the corresponding subtask.
[…]
Figure 3: Teacher Network Structure. On the left, each individual teacher policy $\pi^{k}$ takes a set of privileged and proprioceptive information as input, flattening and concatenating them before passing it through an MLP. In the case of actor-critic learning, both critic and actor networks have the same architecture, with the difference in that the actor network outputs an action distribution while the critic outputs a scalar value. On the right, the full teacher network $\Pi$ is composed of multiple structurally identical networks $\{\pi^{k}\}_{k=1}^{K}$. Given an input id $k$, the $k$-th component network is activated in the computation, i.e. $\Pi[k]\equiv\pi^{k}$.
We can then solve the long horizon task $\mathcal{T}$ via Progressive PEX as follows. We initiate the exploration and learning with a single policy network $\Pi=\{\pi^{1}\}$. This policy network is responsible for learning to solve the initial subtask. Whenever a new subtask is encountered, a policy expansion operation is conducted, adding a new policy into the policy set $\Pi=\{\pi^{1},\pi^{2}\}$. By doing this progressively for the full task, we get
$\Pi$$\triangleq\{\pi^{k}\}_{k=1}^{K},$ (3)
as illustrated in Figure 3 right. A forward pass through $\Pi$ is carried out via the proper sub-policy by indexing the policy set with the task decompositional operator $\mathcal{D}$ conditioned on the current state
$\Pi(\mathbf{a}_{\text{hi}}|\mathcal{D},\mathbf{s})$ $=\pi^{\mathcal{D}(\mathbf{s})}(\mathbf{a}_{\text{hi}}|\mathbf{s}).$ (4)
There are many possible ways to implement the task decompositional operator $\mathcal{D}$. In our case, since it is used in teacher policy learning, we can leverage the fact that privileged observation can be used by the teacher policy and implement it as part of the privileged observation provided by the simulator, e.g., providing the subtask index $k$ directly in the privileged observation. Based this implementation, the training of $\Pi$ is closely related to multi-task RL [55, 45], where we sample a batch of multi-task-transitions (transitions containing the task index $k$) from the replay buffer and use each transition in the batch for training the sub-network associated with its corresponding task index. In this work, we train the teacher policy using a multi-task variant of SAC algorithm [18, 55, 45],
V The Student: Policy Distillation Guided RL
The student needs to perceive the surrounding environment through RGB images. There is no input regarding which subtask the student is currently solving, and it has to infer this information from the inputs. Because of this, the student learns only a single policy across the entire long-horizon task (no sub-task dedicated policy).
To tackle these challenges, we train the student by distilling the teacher’s multiple sub-task policies into a single task policy $\pi_{\rm stu}$. However, distillation will only give us a policy that works fine but not quite well. It could lead to the state drifting or out-of-distribution (OOD) issues when the policy is rolled out for a long horizon. There is also always an observation gap between the teacher and the student, which sometimes creates a difficulty in minimizing the distillation loss. Moreover, due to the differences in model architectures and inductive biases, the student might need a different way of mastering certain skills. Finally, distillation makes the student’s performance upper bounded by the teacher’s. Given these considerations, we choose to boost the student’s policy with RL under the same set of task rewards used by the teacher (Fig. 4).
We train the student by modifying SAC [18] to incorporate the distillation loss properly. First, we use a mixed rollout strategy to generate replay data. When rolling out a new episode, with a probability of $\beta$ we will sample from the student policy, or otherwise from the teacher policy. On one hand, we wish to use the teacher to efficiently generate high-performing transitions along the long horizon. On the other hand, we want to keep exploring with the student’s own policy. Second, following SACLite [54] we remove the entropy reward from policy evaluation. Then for policy improvement, we replace the entropy term with the distillation loss, and assign a fixed weight $\alpha$ to it:
$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{o},\mathbf{s}_{r},\mathbf{s}_{p},l,k)\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)}Q\big{(}(\mathbf{o},\mathbf{s}_{r},l),\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha KL[\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{r},\mathbf{s}_{p})||\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)]\Big{]}.\\ \end{array}$
To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform a policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion (e.g., std. for Gaussian). This makes our distillation loss the best of two worlds: imitation learning and entropy regularization.
[…]
Figure 4: A close look at the student’s architecture. Red dashed lines represent data flows that only exist for training in simulation. Solid lines denote the inference process of the student policy. For deployment, the three inputs are replaced with real-world observations.
When designing the student’s representation model, we specifically take visual sim2real gap reduction into consideration, because once after the student is trained, it will be directly deployed in real without any finetuning. While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck to reduce the visual sim2real gap (Fig. 4). The idea is to let the policy depend on as minimal visual information as possible, namely, to limit the complexity of its input vision space. This potentially leads to a better generalization ability. While there might be many choices for the vision bottleneck, from discrete latent codes [19] to canonical images [24], we choose to use a pair of segmentation and depth maps. There are mainly two reasons for this choice. First, we believe that they contain the minimal visual information required to achieve our task, while without incurring much task information loss. Second, our simulator already provides groundtruth segmentation and depth maps, as high-quality supervision for learning such a visual bottleneck. Thus when training the student, we add a supervision loss for predicting the segmentation mask and depth map from its inputs. Note that only the robot and task-related objects appearing in the instruction will be predicted in the segmentation mask. All other irrelevant object pixels will be classified as “background”. After this, the segmentation and depth maps will be fed to the downstream policy for generating actions. Using a segmentation mask for selecting interesting objects is similarly adopted by prior works [29, 52].
VI Task Implementation In Simulation
Figure 5: Task Decomposition as Privileged Information. Given a task, we decompose it into a sequential series of primitive subtasks. As shown above, the task of finding an object A, grasping it, and then dropping into an object B can be decomposed into a series of subtasks. In simulation, the subtask boundaries and transitions are well-defined, serving as a form of privileged information to the teacher. The teacher can leverage this information, and use a separate policy to learn each subtask, thereby conquering the long horizon task. This expertise is then distilled into the student, which is conditioned on both visual camera feed and language commands.
We use a long-horizon pick-and-place task as a concrete example in this work (c.f. Figure 2 left for an illustration of the task), and we will use this task as a concrete example for illustrating the task decomposition as well. Given a language instruction (e.g., “Drop the blue cube into the green basket.”), the robot needs to locate the target object, picking it up and then drop it into the target basket.
As mentioned previously in Section IV, SLIM employs a task decompositional approach paired with progressive PEX for training a teacher policy to solve the long horizon task. More concretely, for the long-horizon task under consideration, it can be naturally broken down into the following sequential steps (c.f. Figure 5):
  1. 1.
    Search: the robot should first search for the target object (the green cube) for picking up (a subtask), based on the RGB images from the gripper mounted ego-centric camera;
  2. 2.
    MoveTo: after finding the target object, the robot moves to the object until it is within the reach of the arm;
  3. 3.
    Grasp: then the robot moves the gripper towards the target object for picking it up. During the grasping process, the robot base should ideally move less compared to the MoveTo subtask to facilitate accurate grasping;
  4. 4.
    SearchWObj: after picking up the target object, the robot should keep it in the gripper and then search for the target basket; within this stage, there should be some coordination between arm and locomotion movement to facilitate search while avoid the cube to be dropped.
  5. 5.
    MoveToWObj: after finding the target basket, the robot should move towards it while keeping the object in gripper, until the target basket is within a range that is appropriate for the robot to reach for dropping the object in its gripper;
  6. 6.
    MoveGripperToWObj: the robot can now move the gripper towards the target basket until reaching an appropriate position on top of the basket for dropping;
  7. 7.
    DropInto: the robot releases its gripper and drops the cube into the basket.
In addition, we can further append an auxiliary Idle subtask, which encourage the robot to adjust its arm back to neutral position after dropping the object into the basket, facilitating subsequent deployment. This is an optional subtask that is not required for task completion. We include it in our task implementation to make the final policy behavior safe and more predictable after completing the full task. When evaluating the real world performance (e.g. Section VIII-D), we do not include this auxiliary subtask into the full task metric. We also exclude it whenever we discuss task-relevant decompositions since it is not a task related subtask.
Formally, we instantiate the task compositional operator as
$\mathcal{T}=\mathcal{C}(\{\tau^{k}\}_{k=1}^{K})\triangleq\tau^{1}\rightarrow\tau^{2}\rightarrow\cdots\rightarrow\tau^{K}$ (5)
where $\tau^{i}\rightarrow\tau^{j}$ denotes that the two subtasks $\tau^{j}$ and $\tau^{j}$ are two adjacent subtasks and $\tau^{j}$ follows subtask $\tau^{i}$ in a sequential manner. In our case, $K=7$ and $\{\tau^{1},\tau^{2},\cdots,\tau^{7}\}$ corresponds to {Search, MoveTo, Grasp, SearchWObj, MoveToWObj, MoveGripperToWObj, DropInto}.
Because of the decomposition, we can implement reward functions for each subtask, making it modular and easier for implementation. From some tasks, we simply use the sparse subtask success reward (e.g. Search). We incorporated delta-distance based shaping reward for tasks that requires to move the robot (or gripper) close to a target region (e.g. MoveTo, MoveToWObj, MoveGripperToWObj, Grasp, DropInto), in addition to the sparse subtask reward. For SearchWObj and MoveToWObj, we also incorporated an arm retract reward, which is another instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (c.f. Figure A.13). Intuitively, the arm retract reward will encourage the robot’s arm to raise up after grasping the object and move close to its neutral position. Its impact will be investigated in Section VIII-B (No Arm Retract baseline).
An additional benefit of task decomposition is that we have the option to incorporate behavior priors in a flexible way into the policy. More details on the subtasks, rewards and behavior priors are provided in Appendix A-A.
VII Sim-to-Real Gap Reduction Techniques
When training a policy in simulation, we would like to ensure that the policy also performs well in the real world. In other words, we must minimize the “sim2real gaps” as much as possible. Generally, we can divide the sim2real gaps for image-conditioned policies into two distinct categories: the dynamics and visual gaps. Here, we outline essential techniques that proved invaluable in addressing both gaps. We stress that the incorporation of such techniques is crucial to real-world performance. For brevity, we discuss a subset of the techniques employed. In-depth details for these techniques can be found in Appendix C.
VII-A Dynamics Gap Reduction
The dynamics gap is essentially caused by misaligned transition functions $P(\mathbf{s}^{t}|\mathbf{s}^{t-1},\mathbf{a}^{t-1})$ between the simulated and real worlds. This difference in dynamics will result in the sim-trained policy experiencing covariate shift when deployed in the real world, which if large enough, can destroy policy performance. In fact, even minor covariate shifts can result in task failure if the task requires fine-grained high precision coordination, e.g., grasping a small object. Therefore, to minimize the dynamics gap, we must minimize the impact of the covariate shift when transferring to real. In particular, we use the following key techniques for addressing the issue: 1) arm joint tracking error minimization, 2) domain randomization, and 3) object perturbations.
Minimizing Arm Tracking Error via PID Control.
We used PID controller for the robot arm. As opposed to PD controllers which will always have gravity-induced tracking errors, the integral term of PID controllers allow for more robust pose-agnostic tracking given that the joint delta targets are sufficiently small for particular operating frequency. This allows us to greatly reduce the dynamics sim-to-real gap while avoiding system identification. More details can be found in Appendix C-A.
Arm Control Perturbation.
Random structured noise is added to the arm control signal to make the learned policy to be robust w.r.t. different sources of control noises arising when transferring to real.
Arm Mount Perturbation.
We randomly perturb the arm mount position and yaw to make the learned policy robust to the sim-to-realgaps that could arise due to the actual mounting accuracy, torso height variations etc..
Object Perturbations.
The final strategy we employ are object perturbations. Due to the relatively deterministic nature of simulations (in contrast to the high-entropy real world), we observed that policies often converge to deterministic strategies. For example, for picking up an object, a policy may learn to always position itself so that the object has the same relative position to its base. This deterministic “memorizing” behavior can result in task failure if the conditions for success are even slightly off. Therefore, we enforce learning robust, reactive policies by perturbing objects of interest during various stages of task execution. Empirically, this helps with both the exploration (as the probability of stuck in a local state is reduced) and the robust ness of the policy (as the state coverage is enlarged).
VII-B Visual Gap Reduction
The second source of sim2real gap for image-conditioned policies is the visual gap. This gap is caused by an RGB distribution mismatch between simulated and real world pixels. Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios. Accordingly, we apply various visual sim2real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
Image Domain Randomization.
Inspired by SECANT [10], we apply pixel-level perturbations to every RGB image to increase the robustness of the perception model and the chance of its successful sim2real transfer. Specifically, we will randomly select a transformation from the list (Gaussian Noise, SaltPepper Noise, Speckle Noise, Random Brightness, Gaussian Blur) and apply it to each RGB image. We choose not to introduce higher-level image randomization/generation [53] because it might impact image semantics.
Random Textures.
We spawn the robot on an empty ground and randomize the ground texture. The texture can be arbitrary and generally have no specific semantics, as long as they provide enough visual distraction to the student. This distraction will force the student to learn to focus on the most important visual features that are related to the task at hand: objects with reasonable sizes and target colors. When deployed in real, we hope that this visual focus will get transferred and help the robot locate objects.
Random Spatial Augmentation.
To make the visual representation module robust, we incorporate spatial augmentations in the visual representation learning [50, 51].
Random Background Objects.
To further increase the visual complexity of the surrounding environment, we randomly sample and spawn a set of background objects in the room around the robot’s initial position. There are two categories of background objects: 1) primitive shapes (cuboid, cylinder, and sphere), and 2) Google scanned objects [9]. We make sure that these objects only serve the purpose of distraction and the student will not confuse them with task objects. It could collide with an background object during training with some penalty reward. When spawning an object, we also randomize its size, orientation, and texture (only for primitive shapes).
Refer to caption Refer to caption
Figure 6: An example scene populated with random background objects. Left: close-up view; Right: birds-eye view.
VIII Physical System, Experiments and Results
In this section, we briefly go over the robot system and the task used to evaluate SLIM.
VIII-A Physical Robotic System
For our robotic system, we use a Unitree Go1 with a top-mounted WidowX-250S manipulator. An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount, as shown in Fig. 7, serving as the robot’s sole RGB vision feed. The depth stream is not used in this work. Additionally, custom 3D-printed, elongated parallel fingers are fitted on the WidowX gripper to extend its reach. The inner surface of the fingers are padded with a thin layer of foam, which act to increase grasping friction. Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around 12K USD. With the non-edu version, total cost is even lower (around 6K USD) and reflects more closely the cost of mass produced parts..
Figure 7: Qualitative Overview of the Robot System.
[…]
VIII-B Baselines
We compare SLIM with a number of baselines as detailed below:
  • No Arm Retract: removing the arm retract (raise) reward in RL training, which is used to encourage the robot’s arm to raise up after grasping the object (introduced in Section VI);
  • No Perturb: removing arm control randomization, arm mount perturbation and object perturbation (introduced in Section VII-A)
  • No Visual Aug: no background objects and no random spatial augmentation based visual augmentation (introduced in Section VII-B)
  • Distillation Only: agent training without using the RL loss (introduced in Section V), keeping only the distillation and representation loss.
We also tried another baseline No Distillation, which trains the visuomotor policy directly via RL without distillation from the teacher policy (i.e. RL for visuomotor policy learning from scratch). This baseline cannot learn to solve the task at all in simulation, because of the compounded difficulties of representation learning, behavior learning and long-horizon exploration. We have therefore excluded it from the subsequent real-world experiments.
VIII-C Evaluation Protocol
We use the Standard object spatial layout as shown in Figure 8 in a real-world scene Lobby. By doing so, it is easier to maintain repeatability across all methods and seeds. Later in subsequent experiments, we will evaluate the robot in other scenes or with other layouts. For task objects, we use cubes and baskets of different colors. More details on these physical objects are detailed in Appendix E-A.
For each method of each random seed, we roll out the robot’s policy for 20 episodes with varied initial scene states (relative spatial positioning and object colors etc.) according to the evaluation protocol. For the detailed protocol, please refer to Appendix E-B.
TABLE I: Real-world task success rates (mean±$\pm$±stddev), averaged across 3 random seeds, with 20 episodes per seed.
Method Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
[…]
No Visual Aug 98.3±2.498.3.2.498.3$\pm$ 2.498.3±2.4 88.3±4.788.3.4.788.3$\pm$ 4.788.3±4.7 63.3±26.663.3.26.663.3$\pm$ 26.663.3±26.6 56.6±22.556.6.22.556.6$\pm$ 22.556.6±22.5 52.9±18.752.9.18.752.9$\pm$ 18.752.9±18.7
Distillation Only 93.3±4.793.3.4.793.3$\pm$ 4.793.3±4.7 71.1±33.271.1.33.271.1$\pm$ 33.271.1±33.2 60.0±38.960.0.38.960.0$\pm$ 38.960.0±38.9 50.0±35.550.0.35.550.0$\pm$ 35.550.0±35.5 61.2±25.561.2.25.561.2$\pm$ 25.561.2±25.5
[…]
VIII-D Results and Analysis
[…]
Figure 8: Scene Spatial Layouts. Standard: we put the robot in the center of a square space and objects including objects to be grasped and containers to drop the graspable objects into on the four corners. The instructions and object placements vary across episodes. Cluttered: objects are placed close to each other in front and behind the robot, leading to OOD spatial layouts. The instructions and object placements vary across episodes.
We conduct experiments on the baseline and the proposed methods, and measure their success rates in the real world. The results are summarized in Table I.
We run this experiment with 3 different random seeds (for low-level, expert and agent training) and measure the performance in real, using the following two metrics to characterize both accuracy and efficiency:
  1. 1.
    Cumulative Subtask Success Rate: it is defined as the success rate upto the specified subtask starting from the beginning of the full task.
  2. 2.
    Episode Time: it is defined as the time completing the full task from the beginning. Since we applied a time limit of $t_{\rm max}\!=\!90$s in evaluation, for failed episodes, we set their episode time as $t_{\rm max}$.
Then we calculate the mean and standard deviation across the seeds as the final aggregated metric values and the results are summarized in Table I.
It can be observed from the Table I that No Arm Retract baseline performs well for the initial two subtasks, but for later subtasks, it does not perform well. Note that even though it can move to the basket with the object with a success rate of $28.3\%$, it can only drop the object into the basket with a success rate of $5.0\%$. The major reason is that without arm retract reward, the arm is typically at a pose that is not friendly for learning dropping into. No Perturb baseline has the lowest grasping success rate. Actually, without perturbations (object, arm mount and control), the policy is not robust both during learning and can be stuck during exploration. This is reflected in the low and full task success rate, indicating the importance of perturbations for learning. No Visual Aug’s performance on the initial Search subtask is lower than the that of above baselines. The main reason is that without visual augmentation, it can be distracted by other task-irrelevant background objects and can sometimes have a false positive in detecting the target. This implies that the visual augmentation is important for improving the robustness of the robot’s vision. Distillation Only baseline has an overall success rate that is higher than No Arm Retract and No Perturb, reconfirming the importance of arm-retract and perturbations. On the other hand, its performance gap with distillation-guided RL (SLIM which uses both distillation loss and RL loss) shows that distillation loss is not enough for achieving a consistently high performance.
The proposed method SLIM achieves the highest subtask and full task success rate among all the compared methods. Moreover, SLIM is also the most efficient, taking only 43.8s to accomplish the full task on average.
VIII-E Common Failures for SLIM
We also recorded common failures for the full SLIM method. When grasping, in one case, Go1 stopped too late, and kicked the cube far away. In the other case, the low level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach. Drop into often fails because of either early dropping right outside of the basket, or Go1 kicking the basket back when backing toward the basket, right before dropping. We believe most of these can be fixed, in fact, during development, we have seen some of our best policies to be more robust in cases similar to these failures, and achieve around 95% success.
[…]
Figure 9: Real-world Performance of SLIM across Scene Variations. Left: images showing the differences for evaluation. Right: the full task success rate across difference scenes.
VIII-F Analysis of Simulation and Real Performance
We also summarize the performance of different baseline methods as well as SLIM in Figure 10 (averaged over three seeds). The Sim result for each method is obtained by evaluating the corresponding policy in simulation under a setting that is the same as its training setting (in distribution). The Real result denotes the full task success rate when evaluating the policy in the real-world, following the Standard spatial layout as shown in Figure 8. It should be noted that different methods have different abilities and behaviors after their policies are transferred to real. Methods performs well in sim (e.g. the No Visual Aug baseline) does not necessarily translate to a high-performing policy in real. Without arm retract (No Arm Retract), the exploration because more difficult because the camera view is usually restrictive after grasping an object. In addition, it also causes a large sim-to-real gap, as moving with an outreaching arm will cause arm shakiness due to factors such as backlash and arm’s elasticity, which are hard to model in simulation. While one might expect perturbations to complicate the learning process, their inclusion result in much higher performance even in simulation. In the absence of perturbations (No Perturb), the policy is susceptible to converging to suboptimal solutions due to a lack of exposure to a sufficient range of task parameter combinations, effectively lacking a proper curriculum. The proposed SLIM performs well in simulation and maintains the highest success rate when transferred to real.
[…]
Figure 10: Full Task Success Rates of Different Methods in Simulation and Real World.
VIII-G Real-world Performance Across Different Scenes
In addition to the scene used in Table I, here we further run real-world evaluations of SLIM for the following scene variations (c.f. Figure 9 left): Outdoor, Carpet-Room, Kitchen and Lobby. For the first three scenes, we use the standard spatial layout (c.f. Figure 8). For Lobby, we use the Cluttered spatial layout, which is out of the training setting in simulation (Standard layout based). The results are summarized in Figure 9.
VIII-H Policy Behavior Visualization
The behavior of the policy is shown in Figure 11. To make the visualization clear we did not add background objects in the simulation for this episode. The robot start with a randomized position and yaw within a region (Start phase). A red cube and a green cube as well as a red basket and and blue basket are used. Their positions are randomized around the Standard layout and their yaw angles are also randomized. Given an instruction (“Drop the yellow cube into the red basket.”), the robot will first search for the target graspable object (the yellow cube). Note that during the Search phase, the robot actively moves both its body and the arm in searching of the target object (c.f. Figure 11-Search). Once the target object is in view, the robot will move towards it (Figure 11-MoveTo). Once the robot moves close to the target object, it will start to grasp (Figure 11-Grasp). During this stage, the robot body remains stationary and the arm will move towards the target cube and grasp it. After picking up the target cube, the robot will search for the target basket while keeping the cube in gripper. As shown in Figure 11-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory). At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular ee-trajectory in Figure 11-SearchWithObj. Once the target basket is in view, the robot will move towards it (Figure 11-MoveToWithObj) and drop the cube in gripper into that basket (Figure 11-DropInto).
[…]
Figure 11: Policy Trajectory Visualization. For visualization purpose, we removed background objects. We also visualize the ee trajectories. Trajectories corresponding to different subtasks are rendered with different colors. The ego-centric camera view is shown on the top-right of each image. The red colored arrows denote the direction of movement.
IX Limitations
Although not the main focus of this work, the current form of low-level policy introduces several limitations. One limitation is the applicable type of terrains. Currently the low-level policy is only trained on flat terrain in simulation. Although it shows the potential to work across different types of grounds (e.g. concrete, tile floor, carpet, mulch etc.), it does not push the low-level policy learning to its full potential. We can potentially leverage the recent advances in locomotion policy training [61, 6] to further expand the ability and applicability of SLIM across more diverse types terrains, e.g. stairs, ramp etc.
Another low-level-related limitation is that the current form of low-level policy is not a wholebody policy. Although the current form of low-level policy is sufficient for the task considered in this work, by generalizing it to a wholebody policy [14], we can potentially further enlarge its potential workspaces.
As a pure sim-to-real approach that is trained from scratch, we did not leverage any pre-trained foundation models. As a result, we did not address large amount of visual concepts and language variations and instead focused on the right amount of variations that are both practical and representative enough.
There are also a few limitations on the robot system side. Currently for the visual perception, we only used the RGB image from the wrist-mounted camera, which is the minimum setting that is sufficient for the current work. To further improve the robot’s perception and therefore the overall ability in terms of navigation and obstacle avoidance, it might be necessary to incorporate additional sensors for perception such as depth, stereo and lidar etc. In additiona, this can also provide valuable information for the robot to improve its longer-term memories via SLAM-like mechanisms [56, 5].
X Conclusion, Discussion and Future Work
We present a complete legged mobile manipulation robot system for solving language-instructed long horizon task in this work. The policy is trained via the sim-to-real paradigm, together with a progressive policy expansion-based teacher policy for solving the long-horizon task, followed with a distillation-guided RL approach for student visuomotor policy learning. We further identified and reduced a number of crucial sim-to-real gaps. Experimental results compared with a number of baseline methods verified the effectiveness of the proposed SLIM method. Moreover, real-world testing of SLIM across different scenes and spatial layouts shows that the proposed method can generalize well, further demonstrating the effectiveness of the proposed whole pipeline.
For future work, we would like to address the limitations discussed in Section IX, including further expanding the applicability of SLIM across more diverse scenarios by improving locomotion policy training [61, 6, 14]. Extending the visual and language diversity is another future work. Equipping the robot with additional sensors and enable it with SLAM-like ability [56, 5] is also an exciting future direction.
Appendix A RL Formulation
A-A Subtask Boundaries, Rewards and Behavior Priors
Subtask Boundaries.
Since we decompose the full long task into a sequence of subtasks, in practice, we need to decide the task boundaries in order to properly divide the full task. In the sequential decompositional case as described in Eqn.(5), the task boundaries $\tau^{i}|\tau^{i+1}$ between two adjacent subtasks $\tau^{i}\rightarrow\tau^{i+1}$ can be defined by the success condition $\mathcal{S}(\tau^{i})$ of the subtask $\tau^{i}$ (i.e. $\tau^{i}|\tau^{i+1}\equiv\mathcal{S}(\tau^{i})$), since the success of a subtask $\tau^{i}$ naturally leads to the subsequent subtask $\tau^{i+1}$.
We provide the success conditions of each subtask (therefore the subtask boundaries) in the sequel:
  1. 1.
    $\mathcal{S}$(Search): succeeds when the target object is in the view of the gripper-mounted RGB camera;
  2. 2.
    $\mathcal{S}$(MoveTo): succeeds when the target object enters a workspace of the robot within arm’s reach;
  3. 3.
    $\mathcal{S}$(Grasp): succeeds when the target object is in gripper and being lifted above the ground higher than 5CM and lasts at least two timesteps;
  4. 4.
    $\mathcal{S}$(SearchWObj): succeeds when the target object is in gripper and the target basket enters the view of the gripper-mounted RGB camera;
  5. 5.
    $\mathcal{S}$(MoveToWObj): succeeds when the target object is in gripper and the target basket enters a workspace of the robot within arm’s reach;
  6. 6.
    $\mathcal{S}$(MoveGripperToWObj): succeeds when the target object is in gripper and the grippers moved to the top of the target basket;
  7. 7.
    $\mathcal{S}$(DropInto): succeeds when the target object enters the target basket.
Rewards.
We summarize the description of the subtask reward in Table II. The task reward represents the sparse task success ($+1$) or failure ($-1$) reward. For timeout, we used a reward of 00. Apart from the sparse task reward, we also incorporated some shaping rewards as shown in Table II. Most of the shaping rewards are delta-distance based shaping reward, encouraging two position of interest to getting closer to each other. Arm retract reward is also an instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (c.f. Figure A.13). Intuitively, the arm retract reward will encourage the robot’s arm to raise up after grasping the object and move close to its neutral position. Alignment reward is also an instance is delta-distance based shaping reward, where the distance is defined based on the cosine between the left and right finger-tip-to-object vectors. Keep-grasping reward is a contact based reward issuing a penalty of $-0.1$ is the contact between gripper fingers and the object is lost.
TABLE II: Rewards and Behavior Prior for Subtasks
Subtask Reward Behavior Prior
Search task rotational
MoveTo task + distance
Grasp task + distance + alignment stationary
SearchWObj task + arm-retract rotational
MoveToWObj task + distance + arm-retract + keep-grasping
MoveGripperToWObj task + distance + keep-grasping stationary
DropInto task + distance stationary
Behavior Prior.
Another benefit of task decomposition is that we can easily incorporate different behavior priors to the policy. We used two types behavior priors in this work:
  • rotational: the forward velocity command ($\mathbf{c}[0]=0$) to the quadruped is kept as zero and therefore enable the quadruped to move in a rotational fashion. This is used as the behavior prior for search related subtasks (Search, SearchWObj).
  • stationary: both the forward and angular velocity of the command to the quadruped are kept as zero ($\mathbf{c}=0$). This encourages the robot to be stationary. It is used in manipulation related tasks (Grasp, MoveGripperToWObj, DropInto).
In practice, both priors are implemented in a similar way, by initializing the teacher’s actor network for those subtasks with these priors to be of zero mean and small std for the corresponding action dimensions (e.g. the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior). The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution.
Hierarchical RL.
As typical in standard hierarchical RL stuctures, the high-level policy and low-level policies are running at different frequencies. We refer one high-level inference step as a re-planning step. The high-level re-planning period is $5$, meaning the high-level policy (teacher or student) run once every $5$ low-level steps, and the same high-level output is used across this $5$ low-level steps re-planning period for the low-level policy.
Following this structure and given a frozen low-level policy, we are interested in training the high-level (teacher/student) policies. Because of this, the RL is essentially running at the high-level. Therefore, the per-(low-level-)step reward during each re-planning period is accumulated as the step reward for one high-level step, i.e.
$r_{\rm high\_level\_step}=\sum_{i=1}^{5}r_{\rm low\_level\_step}^{i},$ (6)
and this accumulated reward $r_{\rm high\_level\_step}$ is used as the per-step reward signal for high-level RL training.
A-B Low-level Task and Reward
Low-level Tasks.
Details of low-level training tasks/commands and corresponding rewards are summarized in Table III.
TABLE III: Low-level Training Tasks and Rewards
Task/Command Name Sample Range Reward
forward velocity $(-0.5,1.2)$ m/s $e^{-\frac{|v_{\text{forward}}-v_{\text{target}}|}{\max(v_{\text{target}},0.5)}}$
angular velocity $(-1.0,1.0)$ rad/s $e^{-|v_{\text{yaw}}-v_{\text{target}}|}$
stay stationary with probability $0.1$ $\begin{cases}-1&\text{if any foot lifted}\\ 0&\text{otherwise}\end{cases}$
balance and survival
tilt thresh: $0.6$
height thresh: $0.26$
$-(\mathbbm{1}_{\text{over-tilt}}+\mathbbm{1}_{\text{falling}})$
$-0.3\cdot\mathbbm{1}_{\text{killed}}$
gait following (optional)
Low-Level Domain Randomization.
The domain randomization setting for low-level training is summarized in Table IV.
TABLE IV: Environment Randomization Parameters
Env Params Sample Range
base payload $(-0.5,3)$
torso mass center $(-0.15,0.15)$ for each dimension
feet friction $(0.25,1.75)$
feet softness (solimp[2]) [^3] Only the third entry of the solimp parameter of Mujoco is randomized. It is roughly proportional to the size of the deformation of its feet when the robot stands still on the ground. $(0.005,0.02)$
griper payload $(0,0.1)$
Kp of leg joints $(0.7,1.3)$
actuator gain of arm joints $(0.7,1.3)$
bumpiness of the ground $(0,0.15)$ for $90\%$, 00 for $10\%$
delay of sensors $(u-0.002,u+0.002),\ u\in(0.003,0.013)$
Appendix B Model Details
B-A Teacher High Level
The teacher network $\Pi\!=\!\{\pi^{k}\}_{k=1}^{K}$ is a set of sub-policies (referred to as policy set), as shown in Figure 3. The effective size of the policy set is $K\!=\!8$ (including the Idle subtask).
Each policy net $\pi^{k}$ (for subtask $k$) is implemented as as a 3 layer MLP, with the layer size as $512$ ($[512,512,512]$). The critic net for each subtask uses the same MLP structure and with the size of ($[512,512,512]$). Frame stacking is used with a stacking size of $5$.
The re-planning period is $5$, meaning the high-level policy run once every $5$ low-level steps.
We used four 3090 GPUs for training, with a batch size of $2000$, learning rate of $1{\rm e}^{-4}$, gradient clipping of $0.1$, number of parallel environment of $120$, and training iterations of $4$M.
B-B Student High Level
Maps Prediction U-Net.
An illustration of the U-Net [41] for predicting segmentation and depth maps is shown in FigureA.14. We will talk about the architecture details by input modality below.
  1. 1.
    Language. Due to the simplicity of our language instructions, we directly encode each instruction into a sequence of byte tokens. We set the max sequence length to be 100, padding shorter sequences with zeros while truncating longer sequences beyond this limit. Then each byte token is converted into an embedding vector of 128 dimensions by looking up a learnable embedding table. The sequence of language embeddings, with learnable positional encodings, is fed to an instruction encoder transformer [47] with self-attention. The transformer has 2 layers, each layer with 8 attention heads. Each attention head has a key dimensionality of 64 and a value dimensionality of 64. The MLP for generating the residual output has 128 hidden units. Finally, the output embeddings are average pooled to generate a compact instruction-level embedding of 128 dimensions
  2. 2.
    Proprioceptive state. We use a temporal stack of $N=5$. Different proprioceptive observations are simply flattened and concatenated together to produce a 1D robot state vector.
  3. 3.
    Vision. We use a temporal stack of $N=5$ RGB images for vision. The images are stacked along the channel dimension and fed to a U-Net encoder containing 6 convolutional blocks. Each block has a convolutional layer $(filters,kernel\_size,strides,padding)\triangleq(c,k,s,p)$, an instance norm layer [46], a ReLU activation, and a 2D max pooling layer (only starting from the second block). The convolutional layers are configured as $c=[32,64,64,128,128,128]$, and $k=3\times 3,s=1,p=1$ for all blocks. We use FiLM [38] conditioned on the instruction embedding to modulate each intermediate convolutional output, before it is fed to the next convolutional layer. After getting the output from the encoder, we concatenate it with the robot state vector and the instruction embedding, and feed it to the non-skip MLP. The MLP has two hidden layers of 512 units. Finally, the output from MLP is projected back to the output space of the encoder, and is fed to the U-Net decoder to produce the predicted map. The map has $M+1$ channels, where $M=4$ is the predefined maximal number of segmentation classes, and the remaining one channel is for depth prediction.
The maps prediction U-Net is trained by groundtruth segmentation and depth in simulation.
Policy Inputs Encoder.
An illustration of the policy input encoder network is shown in Figure A.15. The network is responsible to digest predicted segmentation&depth maps and proprioceptive observations, and produce a compact representational embedding for downstream RL networks. Again, we use a temporal stack of $N=5$ for both segmentation and depth map inputs. After converting each segmentation map to a one-hot representation, we stack all maps along the channel dimension, resulting in an input of $5\times(4+1)=25$ channels. The input is fed to a ResNet [20] of 5 residual blocks, each block with 64 filters, a kernel size of $3\times 3$, and a stride of 2. The output from the ResNet is spatially flattened, resulting in a sequence of vision tokens. On the other hand, we use an MLP with one hidden layer of 256 units to project the concatenated proprioceptive observation into a token embedding of 256 dimensions. This proprioceptive token is appended to the vision token sequence which is fed to a representation transformer with self-attention. The transformer has 5 layers, each layer with 8 attention heads. Each attention head has a key dimensionality of 128 and a value dimensionality of 128. The MLP for generating the residual output has 256 hidden units. Finally, we take the transformed output corresponding to a special learnable CLS token as the final output of the entire policy input encoder.
Critic and Actor Networks.
Both the critic and the actor have three hidden layers of $(512,512,512)$. The actor outputs a squashed Gaussian distribution as in SAC [18]. Note that the RL and distillation gradients can be propagated all the way back to the policy input encoder and thus affect the representation learning.
Appendix C Sim2Real Techniques
C-A Dynamics Gap Techniques
Minimizing Tracking Error via PID Control.
An arm joint tracking error $\mathbf{e}_{\textrm{arm}}$ is the positional difference between the arm command of the previous timestep and the joint positions of the next timestep,
$\mathbf{e}^{t}_{\textrm{arm}}=\mathbf{s}^{t}_{\textrm{arm}}-\mathbf{a}^{t-1}_{\textrm{arm}}.$
Intuitively, we know that if the tracking error $\mathbf{e}_{\textrm{arm}}$ is equivalent between sim and real for all encounterable states, then there is essentially no dynamics gap. With this, one solution would be to match the tracking errors as closely as possible through accurate embodiment modeling with low-level PD control tuning. In practice, achieving this alignment for the entire state distribution proves to be nontrivial given both the nonlinearity of manipulator dynamics in conjunction with numerous unmodeled effects such as wear-and-tear, motor backlash, etc.
Rather than matching non-zero tracking errors, another approach is to simply achieve near-zero tracking errors in both sim and real. In other words, whatever arm command is outputted, the resulting arm position should be roughly achieved $\mathbf{s}^{t}_{\textrm{arm}}\approx\mathbf{a}^{t-1}_{\textrm{arm}}$. In fact, prior studies have shown that tracking error minimization is one of the most reliable strategies for enabling successful sim2real transfer [3]. To do so, we use a PID joint position controller for our manipulator in both simulation and in real, with a small joint position change limit $z$ (Section III-A) to ensure minimal tracking error for a wide variety of pose configurations.
Though the leg joints also possess a dynamics mismatch, we note that policy’s task success is minimally impacted by sim-to-real locomotion differences. Therefore, we opt to simply use PD position control for the legs.
C-B Visual Gap Techniques
No Camera Depth Data.
For our setting, language-guided visual recognition in the RGB space is the key to finish the task. Without pretrained vision models, this sim-to-real transfer in RGB has to be supported by our method. Given this, we choose not to use camera depth data because this will introduce an extra sim-to-real gap, which has to be reduced with denoising or hole-filling techniques [22]. We observe that depth can be reasonably estimated from stacked RGB images, even with monocular vision.
Color Modeling.
Visual sim-to-real, especially for colors, is notoriously difficult [40]. We simplify the problem by only using four distinct colors in our experiments, namely, red, green (more like cyan), blue and yellow. These colors are sampled from real objects in different lighting conditions and are randomized in the HSV space during simulation training. Color samples are shown in Figure A.12.
Figure A.12: Color samples. Each color had 1000 randomly sampled pixels used to estimate the distribution in simulation.
Appendix D Robot Implementation Details
D-A Control Latency Management
Figure A.13: Asynchronous System Architecture. Note that “Onboard*” refers to the low-level controller frequency being unknown due to the use of closed-source software. Furthermore, when using the WidowX’s PID controller, we also take advantage of Dynamixel’s onboard profiling for smooth control.
We measured the latency along the full control cycle, including motor sensor reading ($\sim$3ms), camera observation (30$\sim$40ms), network connection ($\sim$0.2ms with cable and $\sim$5ms on Wi-Fi), and model inference (20$\sim$40ms for the high level  [^4] Later with TensorRT, we were able to reduce this latency to around 10ms. However, the experiments in this paper were conducted without this option back then., and $\sim$5ms for the low level). We implement the following methods to address such delays as much as possible, while properly modeling them in our simulator.
  1. 1.
    Asynchronous low-high inference: most of the model latency comes from the high-level inference, which takes around 20 ms, but can sometimes peak to 40 ms. This could delay the low-level policy (50HZ) considerably and the robot to falter, if the two inferences are performed in sequence. Alternatively, in the asynchronous mode, the low-level policy will reuse the previous locomotion command if the new one has not arrived yet, and thus still maintain its 50HZ frequency. This leads to a more stable gait of the quadruped.
  2. 2.
    Ethernet cable connection to the robot: Wi-Fi connection is more convenient than cable. However, we find the Wi-Fi connection between Go1 and the laptop to be unstable, sometimes spiking to a latency of 100ms, which is unacceptable for low-level motor control. It is possible to have a Wi-Fi solution with more powerful and stable hardware, or to run low-level inference on Go1 instead of over Wi-Fi, but we simply rely on cable connection in this work.
  3. 3.
    Sensor delay randomizations: motor sensor reading delay is sampled from $(3,13)$ms per episode with $\pm$2ms variation uniformly per control step. Camera observation delay is sampled from $(30,50)$ms uniformly per step with observations taken at 60HZ. Low-level control delay is sampled from $(5,7)$ms uniformly per step. High-level control delay should have been sampled from $(20,40)$ms, but was accidentally set to none during implementation. We did not repeat the full evaluation after discovering this bug, due to limited time. We did test introducing an additional 40ms control delay during physical deployment to the existing high-level policies. It did not impact performance in any significant way. We do plan to train models with the correct delay randomization in the future.
Figure A.14: The architecture of the student’s U-Net for predicting segmentation and depth maps.
Figure A.15: The architecture of the student’s policy input encoder that generates a compact observation embedding to be used by RL.
D-B Safety Measures for the Physical Robot
To reduce damage to the robot itself and the surroundings, we employ many strategies.
  1. 1.
    Bridging the sim-to-real gap is one important part of it, so we can have confidence after observing the robot’s behavior in simulation.
  2. 2.
    PID arm control, with reasonable action clipping, results in much smoother and more reliable arm motion, and reduces damage to the arm motors significantly.
  3. 3.
    The randomization of motor strength, payload, center of mass, arm mounting position, etc. allows the policy to adapt to the real world more easily, resulting in very accurate grasping execution, so the arm does not push into the ground or the target objects, which can lead to motor damage.
  4. 4.
    Control delay caused by policy inference time and communication delay leads to poor low level control. Go1 can topple over, and damage components mounted on top. For example, high level inference can sometimes take over 40ms on the laptop. We use asynchronous inference of high and low level to make sure high level inference does not prevent the robot to receive low level commands. We find that the latency of the Go1 Wi-Fi module can spike to 100ms at times, and decide to rely on cable connection instead. We also allow terminating deployment with a button press. When latency still spikes, e.g. when the deployment machine suddenly freezes or network cable disconnects, our driver onboard Go1 automatically stops the quadruped’s motion. This stopping is done using a scripted policy, and is not always safe, but it prevents damage to the hardware in most cases.
  5. 5.
    Additionally, to prevent the robot from wandering off and bumping into the surroundings, during training, we incorporate a rotational search behavior as explained in Section A-A during policy training, for all baseline methods and the proposed SLIM method.
Appendix E Real-World Experiment Details
E-A Containers and Graspable Objects
We use cubes of different colors as graspable objects, and baskets of different colors as the target container for dropping into, as show in Figure A.16. The size of the basket is about $16{\rm cm}\!\times\!16{\rm cm}\!\times\!16{\rm cm}$. The size of the cube is about $2.5{\rm cm}\!\times\!2.5{\rm cm}\!\times\!2.5{\rm cm}$ and the weight of the cube is about $9{\rm g}$.
[…]
Figure A.16: Containers and Graspable Objects.
E-B Real-world Evaluation Protocols
We detail the protocols used for real-world quantitative evaluation in the main paper (Table I).
TABLE V: Scene Setups for Real-World Quantitative Evaluations. Legends: $\uparrow$ denotes the position robot with its orientation aligned with the direction of $\uparrow$. Uppercase letters denote baskets and lower case letters denote the objects to be grasped. The color of an object is denoted both by the background color and the character (e.g. g denotes the green cube and R denotes the red basket.)
Scene setup Instruction: ”Drop x into Y” (x $\mapsto$ Y)
[…]
g $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
[…]
g $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
[…]
g $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
[…]
g $\mapsto$ B
b $\mapsto$ R
b $\mapsto$ B
[…]
y $\mapsto$ B
r $\mapsto$ R
r $\mapsto$ B
  1. 1.
    For each method, we run 20 episodes with different initial scene states;
  2. 2.
    We apply a time limit of 90 seconds for each evaluation episode.
  3. 3.
    To ensure the repeatability of the scene setup, we use the Standard spatial layout as shown in Figure 8, as it is easy to ensure accuracy of the scene setup in terms of spatial positions. In practice, we mark the positions for the robot at the center of the square. Similarly, we mark the 4 corner positions on the lobby floor for placing the objects and baskets. The size of the square is about $3\times 3$ meters.
  4. 4.
    We follow a scene initialization procedure as detailed in Table V across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.

### v2

@@
 Abstract
-We present a low-cost quadruped manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning (RL) purely in simulation.
-The system comprises 1) a hierarchical design of a high-level policy for visual-mobile manipulation following instructions, and a low-level policy for quadruped movement and limb-control, 2) a progressive policy expansion approach for solving the long-horizon task together with a teacher-student framework for efficient high-level training of the high-level visuomotor policy, and 3) a suite of techniques for minimizing sim-to-real gaps.
-   With budget-friendly but limited reliability and performance hardware, and just one wrist-mounted RGB camera, the entire system fully trained in simulation achieves high success rates for long horizon tasks involving search, move, grasp, and drop-into, with fluid sim-to-real transfer in a wide variety of indoor and outdoor scenes and lighting conditions.Extensive real-world evaluations show that on the long horizon mobile manipulation tasks, our system achieves good performance when transferred to real both in terms of task success rate and execution efficiency.
-Finally, we discuss the necessity of our sim-to-real techniques for legged mobile manipulation, and show their ablation performance.
+We present a low-cost legged mobile manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning purely in simulation.
+This system is made possible by 1) a hierarchical design of a high-level policy for visual-mobile manipulation following instructions and a low-level policy for quadruped movement and limb control, 2) a progressive exploration and learning approach that leverages privileged task decomposition information to train the teacher policy for long-horizon tasks, which will guide an imitation-based student policy for efficient training of the high-level visuomotor policy, and 3) a suite of techniques for minimizing sim-to-real gaps.
+   In contrast to previous approaches that use high-end equipment, our system demonstrates effective performance with more accessible hardware – specifically, a Unitree Go1 quadruped, a WidowX250S arm, and a single wrist-mounted RGB camera – despite the increased challenges of sim-to-real transfer.
+When fully trained in simulation, a single policy autonomously solves long-horizon tasks such as search, move, grasp, and drop-into, achieving nearly 80% success.
+This performance is comparable to that of expert human teleoperation on the same tasks but operates in a more efficient way, at 1.5 times the speed of human expert.
+The sim-to-real transfer is fluid across diverse indoor and outdoor scenes under varying lighting conditions.
+Finally, we discuss the key techniques that enable the entire pipeline, including efficient RL training and sim-to-real, to work effectively for legged mobile manipulation, and present their ablation results.
 Figure 1: SLIM in Real.
-Snapshots of the same SLIM policy running in different scenes in the real world, with large scene variations, including ground properties, backgrounds, and distractors etc.
+Snapshots of the same SLIM policy deployed in diverse real-world scenes, featuring significant variations in ground types, backgrounds, distractors, and other environmental factors.
 These scenes are not available in simulation training.
-The annotations at the bottom are added for understanding task progress stages and are not part of the input to the system.
+The annotations at the bottom are added for understanding task progress and are not part of the input to the system.
 I Introduction
 Until recently, robots have primarily been limited to static, highly repetitive tasks, often in industrial manufacturing.
-As the demand for autonomy grows, the need emerges for robots to navigate and interact with their environment – commonly referred to as mobile manipulators.
-Such systems typically handle primitive tasks, such as searching, grasping, and transporting objects, to achieve overarching long-horizon goals.
-Several promising directions exist in training autonomous mobile manipulators: imitation learning from human demonstrations, reinforcement learning (RL) in the real world, and RL in simulation with sim-to-real transfer.
-Each of the directions has its pros and cons.
-Imitation learning benefits from learning directly from human experts teleoperating the hardware, thus, reasonable behavior and some successes can be achieved even with a small amount of training.
-However, it can be expensive to collect demonstrations to cover all the combinatorial variations of the scene, the task setup, and the hardware.
-Reinforcement learning (RL) can automate policy optimization so minimal human effort is needed to adapt to the environment and task variations.
-Furthermore, the resulting policy is often more efficient than human teleoperation.
-However, mastering long-horizon tasks in the real world still poses a significant challenge.
-Directly training an RL policy in the real world can be challenging due to hardware constraints and safety issues.
-Having to rely on real robots, training efficiency can be low, and scalability is often limited.
-Additionally, a model trained in one physical setup may not generalize to another with for example different background or floor texture.
-RL in simulation solves the scalability and adaptability issues by running massively parallel simulations with randomized setups.
-The difficulty lies in the sim-to-real gap.
-Shortcomings of the design often only shows up in real world testing, at the very end of the development cycle.
-Making things worse, several iterations of modeling, design tweaking, policy training, and real testing are often needed to close up a particular sim-to-real gap.
-Given such a long and often expensive pipeline of tuning steps, RL in simulation can be slow to arrive at a working prototype, and can be expensive in terms of engineering costs and server time for simulation and training.
-Therefore, the hope is to develop general techniques to bridge the sim-to-real gaps, so that efforts spent on developing the training pipeline and the robotic system for some tasks can be reused for training future tasks.
-This work focuses on training an end-to-end RL policy in simulation, and zero-shot deploying in the real world on a relatively low-cost robotic system to solve long-horizon, intricate quadruped manipulation tasks.
-Weak and often faulty hardware, the complexity of legged manipulation, and the long horizon are some of the factors that would compound and significantly expand the sim-to-real gap, making this a daunting task.
-To the best of our knowledge, this is the first working solution of a legged manipulator fully trained with RL in simulation, for long horizon tasks, with low cost hardware.
-We use a quadruped instead of a wheel base, for it can adapt to different terrains, but manipulation is also more challenging on legs, due to the unstable legged base.
-This results in a large 19 DOF (12 leg joints, 6 arm joints, and 1 gripper) system, exacerbating the already difficult long-horizon RL exploration issue.
-To mitigate this, we utilize a hierarchical design consisting of a high-level mobile manipulation policy and low-level quadruped controller, and also rely on a teacher-student training framework similar to prior work [29, 59].
+As the demand for autonomy grows, the need emerges for robots to navigate and interact with their environments - commonly referred to as mobile manipulators.
+Such systems can typically handle primitive tasks, such as searching, grasping, and transporting objects, to achieve overarching long-horizon goals.
+There are several promising methodologies for training autonomous mobile manipulators: imitation learning from human demonstrations, reinforcement learning (RL) in the real world, and RL in simulation with sim-to-real transfer.
+Each of them has its pros and cons.
+Imitation learning (IL) benefits from learning directly from human experts teleoperating the hardware, and thus reasonable behaviors and some success in a restricted domain can be achieved even with a handful of trajectories.
+However, diverse behavior data is required to improve the policy’s robustness even for a single task.
+This makes IL costly, especially if we further want to cover all the combinatorial variations of scenes and tasks for robot ability scaling.
+Reinforcement learning (RL) can potentially generate an infinite amount of training data by allowing the robot to explore and interact with the environment.
+Furthermore, the resulting policy is often more optimal than a human teleoperating policy.
+However, directly training an RL policy in the real world can be challenging due to hardware maintenance, reward specification, and safety issues.
+Data scale is also largely limited by the difficulty of deploying a robot fleet for collecting data in parallel.
+As a result, mastering long-horizon tasks in the real world still poses a significant challenge.
+RL in simulation could address the scaling, generalization, and safety issues by running massively parallel simulation with randomized scenes and tasks.
+But to successfully deploy the trained policy on a real robot, the challenge lies in sim-to-real gaps.
+Therefore, we attempt to develop effective and general techniques to bridge the gaps.
+This work focuses on training an end-to-end RL policy in simulation, and zero-shot deploying in the real world on a relatively low-cost quadruped based robot to solve long-horizon, intricate manipulation tasks.
+The complexity of accurate manipulation on legs, the variations of the physical environment, the scene and task setups, and the long task horizon are some of the factors that compound and increase the difficulty of RL training.
+On top of these factors, limited hardware reliability further exacerbates the sim-to-real gap, making this a daunting task.
+We train an end-to-end model that is able to achieve long-horizon tasks consisting of searching, grasping, transporting and dropping primitive tasks (Figure 1).
+For robot mobility, we choose a quadruped instead of a wheeled base, as it has the potential to adapt to many different types of terrains.
+The choice of a quadruped base results in a large 19 DOF (12 leg joints, 6 arm joints, and 1 gripper) system, exacerbating the already difficult long-horizon RL exploration issue.
+To mitigate this, we utilize a hierarchical teacher-student design, consisting of a high-level mobile manipulation policy and low-level quadruped controller, and also a teacher-student training framework [6] for the high-level policy.
+To the best of our knowledge, this is the first demonstration of a low-cost real-world quadruped manipulator fully trained with RL in simulation, for solving long-horizon tasks.
 The teacher policy has access to privileged information (e.g., structured and accurate object features) that is unavailable in the real world.
@@
 The student policy is then trained to distill the teacher policy while maximizing task rewards, but with access to only standard sensor data such as camera feed and proprioceptive state.
-We also condition the student visuomotor policy on language instructions, allowing us to naturally command the robot to achieve different tasks.
+We also condition the student visuomotor policy on language instructions, providing a natural way to the user for specifying the task to be executed by the robot.
 We name our robotic system SLIM (Sim-to-Real Legged Instructive Manipulation).
 We reiterate that SLIM is a complete robotic system with fluid sim-to-real transfer.
-Unlike prior work [29, 52], SLIM learns all modules in simulation from scratch, without the need for any pretrained model or real-world finetuning.
+Unlike prior work [30, 53, 59], SLIM learns all modules in simulation from scratch, without the need for any pretrained model or real-world finetuning.
 SLIM relies on pure RL, and thus it does not require a single demonstration trajectory.
@@
   4. 4.
-    We conduct an extensive set of simulated and real-world experiments, achieving higher success rate and faster task completion than baselines.
-In the real world, we demonstrate robust task success across a wide range of indoor and outdoor environments.
+    We conduct an extensive set of experiments with real-world evaluations, achieving higher success rate and faster task completion than baselines.
+We also demonstrate robust task success across a wide range of indoor and outdoor environments in real-world.
 II Related Work
@@
   • $\mathbf{s}_{\textrm{r}}$
-    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of all joint position readings.
+    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of proprioceptive reading of all joints.
+$\mathbf{s}_{\textrm{r}}\!=\![\mathbf{s}_{\textrm{leg}},\mathbf{s}_{\textrm{arm}}]$, where $\mathbf{s}_{\textrm{leg}}$ denotes the leg-related proprioceptive states and $s_{\textrm{arm}}$ the arm-related proprioceptive states.
+When used as the input to the low-level locomotion policy, we overload $\mathbf{s}_{\textrm{leg}}$ to denote the leg-related proprioceptive states stacked according to the low-level frame-stacking size.
 Privileged Observations (only available in simulation).
@@
 $(\mathbf{c},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a_{\textrm{leg}}}\in\mathbb{R}^{12}.$
-Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$, and has a much shorter time horizon.
+Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$, thus does not need to handle high dimensional image and language inputs.
 Low-Level.
@@
 […]
-Figure 3: Teacher Network Structure.
+Figure 3: Teacher Policy Network Structure.
 On the left, each individual teacher policy $\pi^{k}$ takes a set of privileged and proprioceptive information as input, flattening and concatenating them before passing it through an MLP.
@@
 Then for policy improvement, we replace the entropy term with the distillation loss, and assign a fixed weight $\alpha$ to it:
-$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{o},\mathbf{s}_{r},\mathbf{s}_{p},l,k)\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)}Q\big{(}(\mathbf{o},\mathbf{s}_{r},l),\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha KL[\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{r},\mathbf{s}_{p})||\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)]\Big{]}.\\ \end{array}$
-To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform a policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion (e.g., std.
+$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{o},\mathbf{s}_{r},\mathbf{s}_{p},l,k)\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)}Q\big{(}(\mathbf{o},\mathbf{s}_{r},l),\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha KL[\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{r},\mathbf{s}_{p})||\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)]\Big{]}.\\ \end{array}$ (5)
+To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform a policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion $\sigma$ (e.g., std.
 for Gaussian).
@@
 For deployment, the three inputs are replaced with real-world observations.
-When designing the student’s representation model, we specifically take visual sim2real gap reduction into consideration, because once after the student is trained, it will be directly deployed in real without any finetuning.
-While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck to reduce the visual sim2real gap (Fig.
+When designing the student’s representation model, we specifically take vision sim2real gap reduction into consideration, because once after the student is trained, it will be directly deployed in real without any finetuning.
+While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck to reduce the vision sim2real gap (Fig.
 4).
@@
 First, we believe that they contain the minimal visual information required to achieve our task, while without incurring much task information loss.
-Second, our simulator already provides groundtruth segmentation and depth maps, as high-quality supervision for learning such a visual bottleneck.
+Second, our simulator already provides groundtruth segmentation and depth maps, as high-quality supervision for learning such a vision bottleneck.
 Thus when training the student, we add a supervision loss for predicting the segmentation mask and depth map from its inputs.
@@
 VII Sim-to-Real Gap Reduction Techniques
-When training a policy in simulation, we would like to ensure that the policy also performs well in the real world.
-In other words, we must minimize the “sim2real gaps” as much as possible.
-Generally, we can divide the sim2real gaps for image-conditioned policies into two distinct categories: the dynamics and visual gaps.
-Here, we outline essential techniques that proved invaluable in addressing both gaps.
-We stress that the incorporation of such techniques is crucial to real-world performance.
-For brevity, we discuss a subset of the techniques employed.
-In-depth details for these techniques can be found in Appendix C.
+sim-to-real is notoriously difficult.
+This work handles two kinds of sim-to-real gaps: the physical dynamics and the visual gaps.
+For brevity, we discuss a subset of the techniques which are crucial in addressing sim-to-real and for successful real-world deployment.
+More techniques can be found in Appendix D.
 VII-A Dynamics Gap Reduction
@@
 This difference in dynamics will result in the sim-trained policy experiencing covariate shift when deployed in the real world, which if large enough, can destroy policy performance.
-In fact, even minor covariate shifts can result in task failure if the task requires fine-grained high precision coordination, e.g., grasping a small object.
+In fact, even minor covariate shifts can result in task failure if the task requires fine motor control, e.g., grasping a small object.
 Therefore, to minimize the dynamics gap, we must minimize the impact of the covariate shift when transferring to real.
@@
 This allows us to greatly reduce the dynamics sim-to-real gap while avoiding system identification.
-More details can be found in Appendix C-A.
+More details can be found in Appendix D-A.
 Arm Control Perturbation.
@@
 Empirically, this helps with both the exploration (as the probability of stuck in a local state is reduced) and the robust ness of the policy (as the state coverage is enlarged).
+Stationary Manipulations.
+We make the quadruped stand as still as possible when manipulating objects, i.e.
+during grasp or drop into.
+This avoids the large sim-to-real gap of a walking and shaking dog when accurate manipulation is needed.
+To achieve this, when training low level policies to follow locomotion speed commands, we also make the quadruped enter a stationary mode with a balanced stance when the speed command is close to zero.
+Without stationary manipulation, not only will success rates suffer, but also it becomes unsafe to deploy, because the gripper fingers can hit the ground, damaging fingers and arm motors.
 VII-B Visual Gap Reduction
-The second source of sim2real gap for image-conditioned policies is the visual gap.
+The second source of sim-to-real gap for image-conditioned policies is the vision gap.
 This gap is caused by an RGB distribution mismatch between simulated and real world pixels.
 Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios.
-Accordingly, we apply various visual sim2real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
+Accordingly, we apply various vision sim-to-real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
 Image Domain Randomization.
-Inspired by SECANT [10], we apply pixel-level perturbations to every RGB image to increase the robustness of the perception model and the chance of its successful sim2real transfer.
+Inspired by SECANT [11], we apply pixel-level perturbations to every RGB image to increase the robustness of the perception model and the chance of its successful sim-to-real transfer.
 Specifically, we will randomly select a transformation from the list (Gaussian Noise, SaltPepper Noise, Speckle Noise, Random Brightness, Gaussian Blur) and apply it to each RGB image.
@@
 VIII-A Physical Robotic System
-For our robotic system, we use a Unitree Go1 with a top-mounted WidowX-250S manipulator.
+For our robotic system, we use a Unitree Go1 with a top-mounted WidowX250S manipulator.
 An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount, as shown in Fig.
@@
 The inner surface of the fingers are padded with a thin layer of foam, which act to increase grasping friction.
-Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around 12K USD.
-With the non-edu version, total cost is even lower (around 6K USD) and reflects more closely the cost of mass produced parts..
+Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around \$12K USD.
+With the non-edu version, total cost is even lower ($\sim$6K USD) and reflects a bit more closely the cost of mass produced parts..
 Figure 7: Qualitative Overview of the Robot System.
@@
 VIII-B Baselines
-We compare SLIM with a number of baselines as detailed below:
+We compare SLIM with a number of baselines below:
   • No Arm Retract: removing the arm retract (raise) reward in RL training, which is used to encourage the robot’s arm to raise up after grasping the object (introduced in Section VI);
@@
   • Distillation Only: agent training without using the RL loss (introduced in Section V), keeping only the distillation and representation loss.
+  • Human Teleop: an expert human trained for half an hour in teleoperating the robot with locomotion and arm end-effector commands observing only the view of the wrist camera.
+While not a baseline per se, human teleoperation provides a valuable reference point.
+See Appendix B-D for more details.
 We also tried another baseline No Distillation, which trains the visuomotor policy directly via RL without distillation from the teacher policy (i.e.
@@
 This baseline cannot learn to solve the task at all in simulation, because of the compounded difficulties of representation learning, behavior learning and long-horizon exploration.
-We have therefore excluded it from the subsequent real-world experiments.
+It is excluded from the tables.
 VIII-C Evaluation Protocol
@@
 By doing so, it is easier to maintain repeatability across all methods and seeds.
-Later in subsequent experiments, we will evaluate the robot in other scenes or with other layouts.
+Besides the Standard in lobby, we also evaluate the robot in several other scenes with other layout.
 For task objects, we use cubes and baskets of different colors.
-More details on these physical objects are detailed in Appendix E-A.
-For each method of each random seed, we roll out the robot’s policy for 20 episodes with varied initial scene states (relative spatial positioning and object colors etc.) according to the evaluation protocol.
-For the detailed protocol, please refer to Appendix E-B.
-TABLE I: Real-world task success rates (mean±$\pm$±stddev), averaged across 3 random seeds, with 20 episodes per seed.
-Method Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
-Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
+More details on these physical objects are detailed in Appendix F-A.
+For each method of each random seed, we roll out the robot’s policy for 20 episodes with varied initial scene setup (i.e.
+different positioning and colors of objects) according to the evaluation protocol.
+For the detailed protocol, please refer to Appendix F-B.
+TABLE I: Real-world task success rates and episode time (mean±$\pm$±stddev over 3 random seeds), with 20 episodes per seed.
+Method Fully Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
+Autonomous Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
 […]
@@
 Distillation Only 93.3±4.793.3.4.793.3$\pm$ 4.793.3±4.7 71.1±33.271.1.33.271.1$\pm$ 33.271.1±33.2 60.0±38.960.0.38.960.0$\pm$ 38.960.0±38.9 50.0±35.550.0.35.550.0$\pm$ 35.550.0±35.5 61.2±25.561.2.25.561.2$\pm$ 25.561.2±25.5
+[…]
 […]
@@
 The instructions and object placements vary across episodes.
-We conduct experiments on the baseline and the proposed methods, and measure their success rates in the real world. The results are summarized in Table I.
-We run this experiment with 3 different random seeds (for low-level, expert and agent training) and measure the performance in real, using the following two metrics to characterize both accuracy and efficiency:
+We conduct experiments on the baselines and the proposed method, and measure their success rates in the real world. The results are summarized in Table I.
+We run the whole training pipeline with three different random seeds to train low-level, the teacher and the student, and measure the performance of the final student policy in real, in terms of accuracy and efficiency:
   1. 1.
-    Cumulative Subtask Success Rate: it is defined as the success rate upto the specified subtask starting from the beginning of the full task.
+    Cumulative Subtask Success Rate: the success rate up to the specified subtask from the beginning of the episode.
   2. 2.
-    Episode Time: it is defined as the time completing the full task from the beginning.
-Since we applied a time limit of $t_{\rm max}\!=\!90$s in evaluation, for failed episodes, we set their episode time as $t_{\rm max}$.
-Then we calculate the mean and standard deviation across the seeds as the final aggregated metric values and the results are summarized in Table I.
-It can be observed from the Table I that No Arm Retract baseline performs well for the initial two subtasks, but for later subtasks, it does not perform well.
-Note that even though it can move to the basket with the object with a success rate of $28.3\%$, it can only drop the object into the basket with a success rate of $5.0\%$.
-The major reason is that without arm retract reward, the arm is typically at a pose that is not friendly for learning dropping into.
+    Episode Time: the time spent completing the full task from the beginning. We apply a time limit of $t_{\rm max}\!=\!90$s. For failed episodes, we use $t_{\rm max}$ as episode time.
+Finally, we report the mean and standard deviation across the seeds as the final aggregated metric value in Table I.
+Table I shows how important each of the interventions used contribute to the final successes.
+No Arm Retract baseline performs well for the initial two subtasks of Search and MoveTo, but not for the later subtasks.
+Even though half of the times the robot can move to the basket carrying the cube (a cumulative success rate of $28.3\%$), it only successfully drops the object into the basket less than one fifth of the cases (a cumulative success rate of $5.0\%$).
+The main reason is that without arm retract reward, the arm could be at an over-extended pose, getting closer to the physical torque limits of the shoulder motors responsible for raising the arm.
+This leads to more arm shake, cube slipping out of the fingers, motor failures, and makes it harder for the arm to raise up to search for the basket and to hover above the basket to complete DropInto.
 No Perturb baseline has the lowest grasping success rate.
-Actually, without perturbations (object, arm mount and control), the policy is not robust both during learning and can be stuck during exploration.
-This is reflected in the low and full task success rate, indicating the importance of perturbations for learning.
-No Visual Aug’s performance on the initial Search subtask is lower than the that of above baselines.
-The main reason is that without visual augmentation, it can be distracted by other task-irrelevant background objects and can sometimes have a false positive in detecting the target.
-This implies that the visual augmentation is important for improving the robustness of the robot’s vision.
-Distillation Only baseline has an overall success rate that is higher than No Arm Retract and No Perturb, reconfirming the importance of arm-retract and perturbations.
-On the other hand, its performance gap with distillation-guided RL (SLIM which uses both distillation loss and RL loss) shows that distillation loss is not enough for achieving a consistently high performance.
-The proposed method SLIM achieves the highest subtask and full task success rate among all the compared methods.
-Moreover, SLIM is also the most efficient, taking only 43.8s to accomplish the full task on average.
-VIII-E Common Failures for SLIM
-We also recorded common failures for the full SLIM method.
-When grasping, in one case, Go1 stopped too late, and kicked the cube far away.
-In the other case, the low level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach.
-Drop into often fails because of either early dropping right outside of the basket, or Go1 kicking the basket back when backing toward the basket, right before dropping.
-We believe most of these can be fixed, in fact, during development, we have seen some of our best policies to be more robust in cases similar to these failures, and achieve around 95% success.
+Without perturbations (object, arm mount and control), the policy is not robust enough to deal with the variations in the location of the object relative to the robot.
+This is likely because during simulation training, the policy can get very efficient at approaching the target object and stopping just at the success boundary of the MoveTo subtask, so that the Grasp subtask usually starts with very similar locations of the target object relative to the robot, while in the real world such clear cut positioning is rarely seen.
+No Visual Aug’s performance on the initial Search subtask is lower than the that of the above baselines.
+The main reason is that without visual augmentations, it can be distracted by background objects irrelevant to the task, i.e.
+false positive detection of the target.
+This shows that visual augmentations are important for improving the robustness of the robot’s vision module.
+Distillation Only baseline simply imitates from the highly successful teacher policy, and can indeed achieve a relatively high success rate.
+However, its performance is not consistent across random seeds.
+Our full method (SLIM) uses distillation-guided RL, and results in consistently high success rates across seeds.
+It shows that distillation loss alone is not enough for robust policy learning.
+Additionally, task completion is also slower than the full method which additionally uses RL loss to optimize the policy.
+Our full method SLIM achieves the highest subtask and full task success rate among all the methods, and is also the most efficient, taking only 43.8 seconds to complete the full task on average.
 […]
 Figure 9: Real-world Performance of SLIM across Scene Variations. Left: images showing the differences for evaluation. Right: the full task success rate across difference scenes.
-VIII-F Analysis of Simulation and Real Performance
-We also summarize the performance of different baseline methods as well as SLIM in Figure 10 (averaged over three seeds).
-The Sim result for each method is obtained by evaluating the corresponding policy in simulation under a setting that is the same as its training setting (in distribution).
-The Real result denotes the full task success rate when evaluating the policy in the real-world, following the Standard spatial layout as shown in Figure 8.
-It should be noted that different methods have different abilities and behaviors after their policies are transferred to real.
-Methods performs well in sim (e.g.
-the No Visual Aug baseline) does not necessarily translate to a high-performing policy in real.
-Without arm retract (No Arm Retract), the exploration because more difficult because the camera view is usually restrictive after grasping an object.
-In addition, it also causes a large sim-to-real gap, as moving with an outreaching arm will cause arm shakiness due to factors such as backlash and arm’s elasticity, which are hard to model in simulation.
-While one might expect perturbations to complicate the learning process, their inclusion result in much higher performance even in simulation.
-In the absence of perturbations (No Perturb), the policy is susceptible to converging to suboptimal solutions due to a lack of exposure to a sufficient range of task parameter combinations, effectively lacking a proper curriculum.
-The proposed SLIM performs well in simulation and maintains the highest success rate when transferred to real.
-[…]
-Figure 10: Full Task Success Rates of Different Methods in Simulation and Real World.
-VIII-G Real-world Performance Across Different Scenes
-In addition to the scene used in Table I, here we further run real-world evaluations of SLIM for the following scene variations (c.f.
+VIII-E Generalization to Different Real-world Scenes
+In addition to the standard scene used in Table I, we further run real-world evaluations of SLIM under more scene variations (c.f.
 Figure 9 left): Outdoor, Carpet-Room, Kitchen and Lobby.
@@
 Figure 8).
-For Lobby, we use the Cluttered spatial layout, which is out of the training setting in simulation (Standard layout based).
-The results are summarized in Figure 9.
-VIII-H Policy Behavior Visualization
+For Lobby, we use the Cluttered spatial layout, which is different from the training (Standard) layout.
+Results as summarized in Figure 9 are close to the Standard layout or a bit lower due to being out of distribution.
+This shows that the model trained fully in simulation can zero-shot adapt to a wide range of real world settings, with vastly different lighting conditions, background, and floor materials and textures.
+Furthermore, our qualitative evaluation showed that the robot also generalizes well to scenarios where:
+  1. 1.
+    task objects are randomly scattered on the ground (e.g., last two rows in Figure 1),
+  2. 2.
+    distractor objects are present (e.g., last row in Figure 1),
+  3. 3.
+    human interrupts the task progress, or
+  4. 4.
+    objects to grasp have slightly different shapes.
+We refer the reader to our demo videos for more emergent generalization behaviors. In general, the robot is quite robust to the real-world scene complexity.
+VIII-F Comparison with Simulation Performance
+Different methods or even the same method with different seeds can exhibit very different behaviors, mostly due to the variability in the RL training of the teacher policies.
+Additionally, because of the sim-to-real gap, real deployment can differ from simulation as well.
+Figure 10 shows simulation performance with the real performance for our methods, so we know more clearly how each method helps the training process.
+The Sim success rate for each method denotes the average full task success rate over 200 episodes run in simulation, averaged across the three random seeds.
+The simulation environment follows the same randomization setting as training, with the policy rollout being fully greedy.
+The Real success rate denotes the full task success rate of the policy in the real-world, following the Standard spatial layout as shown in Figure 8.
+Without arm retract (No Arm Retract), searching for the basket becomes more difficult in simulation as well as in real, because after grasp, the camera still points down to the ground, but searching for the basket requires many steps of the arm raising up, creating an exploration bottleneck for RL training.
+Real world performance suffers more, because moving with an out-stretched arm causes large sim-to-real gaps: The tip of the arm tends to shake more due to motor backlash, hard to model in simulation, and the shoulder motors operate close to their physical torque limits, leading to more motor overload.
+For No Perturb, while one might expect perturbations to complicate RL learning, their inclusion results in a much higher performance even in simulation.
+Without perturbations, the policy seems to converge to suboptimal solutions maybe due to a lack of exposure to a sufficiently large range of setups.
+Without visual augmentations (No Visual Aug), simulation task is easier to solve, thus simulation success is artificially high, but real world performance suffers when compared to the full method.
+RL training loss in the full SLIM method increases success in simulation (over the imitation only Distillation Only) and achieves the highest success rate in real.
+It’s interesting to note that the RL loss seems to help the No Visual Aug baseline to achieve slightly better real world success than Distillation Only.
+[…]
+Figure 10: Full Task Success Rates of Different Methods (averaged across three seeds) in Simulation and Real.
+[…]
+Figure 11: Policy Trajectory Visualization.
+For visualization purpose, we removed background objects.
+We also visualize the ee trajectories.
+Trajectories corresponding to different subtasks are rendered with different colors.
+The ego-centric camera view is shown on the top-right of each image.
+The red colored arrows denote the direction of movement.
+VIII-G Policy Behavior Visualization
 The behavior of the policy is shown in Figure 11.
@@
 As shown in Figure 11-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory).
-At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular ee-trajectory in Figure 11-SearchWithObj.
+At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular EE trajectory in Figure 11-SearchWithObj.
 Once the target basket is in view, the robot will move towards it (Figure 11-MoveToWithObj) and drop the cube in gripper into that basket (Figure 11-DropInto).
-[…]
-Figure 11: Policy Trajectory Visualization.
-For visualization purpose, we removed background objects.
-We also visualize the ee trajectories.
-Trajectories corresponding to different subtasks are rendered with different colors.
-The ego-centric camera view is shown on the top-right of each image.
-The red colored arrows denote the direction of movement.
+VIII-H Failure Modes
+We also recorded common failure cases for SLIM.
+In one case, when grasping the quadruped stopped too late and kicked the cube far away.
+In the other case, the low-level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach.
+DropInto sometimes fails because of either early dropping right outside of the basket, or the quadruped kicking the basket away right before dropping when backing towards the basket.
+In rare cases where the gripper holds only the upper half of a cube, the robot tends to get stuck at hovering over the basket while holding the cube, because the cube is completely blocked by the gripper fingers in the wrist camera view.
+See our failure demo videos for more details.
 IX Limitations
@@
   1. 1.
-    $\mathcal{S}$(Search): succeeds when the target object is in the view of the gripper-mounted RGB camera;
+    $\mathcal{S}$(Search): succeeds when the target object is in the view of the wrist-mounted RGB camera;
   2. 2.
@@
   4. 4.
-    $\mathcal{S}$(SearchWObj): succeeds when the target object is in gripper and the target basket enters the view of the gripper-mounted RGB camera;
+    $\mathcal{S}$(SearchWObj): succeeds when the target object is in gripper and the target basket enters the view of the wrist-mounted RGB camera;
   5. 5.
@@
 DropInto task + distance stationary
-Behavior Prior.
+Behavior Priors.
 Another benefit of task decomposition is that we can easily incorporate different behavior priors to the policy. We used two types behavior priors in this work:
@@
 This is used as the behavior prior for search related subtasks (Search, SearchWObj).
-  • stationary: both the forward and angular velocity of the command to the quadruped are kept as zero ($\mathbf{c}=0$).
+  • stationary: both the forward and angular velocity of the command to the quadruped are hardcoded as zero ($\mathbf{c}=0$).
 This encourages the robot to be stationary.
@@
 the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior).
-The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution.
+The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution and being consistent with subtasks without these priors.
 Hierarchical RL.
@@
 The domain randomization setting for low-level training is summarized in Table IV.
-TABLE IV: Environment Randomization Parameters
+TABLE IV: Low-level Domain Randomization Parameters
 Env Params Sample Range
@@
 Each policy net $\pi^{k}$ (for subtask $k$) is implemented as as a 3 layer MLP, with the layer size as $512$ ($[512,512,512]$).
-The critic net for each subtask uses the same MLP structure and with the size of ($[512,512,512]$).
-Frame stacking is used with a stacking size of $5$.
-The re-planning period is $5$, meaning the high-level policy run once every $5$ low-level steps.
-We used four 3090 GPUs for training, with a batch size of $2000$, learning rate of $1{\rm e}^{-4}$, gradient clipping of $0.1$, number of parallel environment of $120$, and training iterations of $4$M.
+A squashed Normal projection layer (a layer for generating Normal distributions) is used as the output layer for each policy net following the standard practice [19], generating the action distribution conditioned on the current input.
+The critic net for each subtask uses the same MLP structure and with the size of ($[512,512,512]$). Frame stacking is used with a stacking size of $5$.
+The re-planning period is $5$, meaning that the high-level policy predicts once every $5$ low-level steps.
 B-B Student High Level
@@
 The actor outputs a squashed Gaussian distribution as in SAC [19].
-Note that the RL and distillation gradients can be propagated all the way back to the policy input encoder and thus affect the representation learning.
-Appendix C Sim2Real Techniques
-C-A Dynamics Gap Techniques
+Note that the RL and distillation gradients will be propagated all the way back to the policy input encoder for learning the representation.
+Stationary Bit.
+The teacher policy incorporates a stationary behavior prior into the locomotion command $\mathbf{c}$ for several subtasks (Table II).
+With this prior, the corresponding locomotion command distribution always has a zero mean and a small std.
+Taking argmax of the distribution, the teacher policy generates exactly zero locomotion command and thus the quadruped can remain stationary.
+However, if we directly distill this distribution into the student’s counterpart, it is almost impossible to guarantee a zero mode, due to the continuous nature of the distribution.
+Thus we employed a sationary bit trick to solve this issue.
+Specifically, we create an extra binary action distribution for the student for predicting whether the quadruped should be stationary or not, which is determined by checking if the locomotion command of the teacher is exactly zero.
+We then add a new prediction loss for learning this stationary bit, while masking out the KLD loss (only for the two locomotion command dimensions) if the stationary bit is true.
+For student inference, whenever the stationary bit distribution outputs 1, the locomotion command $\mathbf{c}$ will be overwritten to be zero.
+B-C Low Level
+For the low-level training, we used PPO [44] with a regularized online adaptation module Fu et al. [15].
+Policy Input and Locomotion Command.
+As shown in Figure 2, the low-level locomotion policy takes two types of inputs: 1) the standard proprioceptive state $\mathbf{s}_{\textrm{leg}}$ obtained from the environment, 2) the locomotion command $\mathbf{c}$ from an external command issuer (e.g..
+The second type of input is crucial to make the low-level policy useful in downstream tasks where locomotion is involved as a basic skill for exploration and learning.
+During training, the command is re-sampled periodically within each episode and the low-level policy is trained to follow the command $\mathbf{c}$ via the command-following rewards detailed in Table III.
+The arm pose is also adjusted periodically within each episode, to make the low-level policy’s command following robust against arm movement.
+For the standard proprioceptive input to the low-level policy, while it is also possible to include those from the arm.
+Empirically, we discover that it might trigger some shakiness behavior in certain cases and therefore we only used $\mathbf{s}_{\textrm{leg}}$.
+After training, the low-level policy can be used to follow commands issued by an external command issuer, e.g.
+either a higher level policy (as in SLIM and all the other baseline except Human Teleop), or human tele-operator (as in Human Teleop) in this work.
+Actor and Value Networks.
+Both the actor and value networks have three hidden layers of size $(256,256,256)$.
+The actor network outputs a Beta distribution instead of the typical Gaussian.
+For numerical stability, we set the minimum value of the shape parameters of the Beta distribution to $1$.
+Online Adaptation Module for Sim-to-Real Transfer
+.
+Similar as in [15, Figure 2], our online adaptation module consists of two encoders, the privileged encoder and the adaptation encoder.
+The privileged encoder takes privileged information as input and predicts a latent representation of environment extrinsics, while the adaptation encoder predicts the same latent representation depending only on robot sensory observations.
+Both encoders are trained to match each other and jointly with the RL objective.
+The privileged encoder network has two hidden layers of $(128,64)$ and the adaptation encoder has hidden layers of $(256,128)$.
+B-D Human Teleoperation
+To show that we’ve made reasonable efforts to make Human Teleop work, we list its failure modes.
+1) The most common failure mode is task timeout.
+This is sometimes due to not fully grasping the cube and the cube slipping out of the fingers during movement, leaving no time to search and re-grasp.
+Failing to grasp fully can be due to either stopping a bit far from the cube, or incorrectly assessing the distance to the cube under monocular vision.
+2) The arm can enter a twisted pose due to poor IK solutions, despite having already tuned IK as we can.
+Human teleoperation is in general less accurate and slower than RL policy, due to Low level policy shaking, and human operator fatigue.
+Besides being slower, human teleoperation seems to be less energy efficient than our RL policies.
+When RL learned high-level policies can run smoothly on a particular low level policy, human teleoperation on the same low level policy seems to trigger over-current protection of Go1.
+This often happens when trying to raise an over-extended arm, causing Go1 to lose control of its leg motors.
+In evaluations, we ignore these Go1 hardware failures, and let the human operator retry for free.
+B-E Model Parameter Count
+We summarize in Table V the rough number of model parameters for different components of our system, including low level, teacher high level and student high level.
+TABLE V: Number of Trainable Parameters.
+Component Approx. Param. Count
+Low Level $3$M
+Teacher High Level (all subtasks) $10$M
+Student High Level $16$M
+Our deployed policy (“Low Level” + “Student High Level”) has a total parameter count of less than 19M.
+Appendix C Important Training Hyperparameters
+Below we list some important training hyperparameters of SLIM.
+Teacher High Level.
+We used four 3090 GPUs for training, with a batch size of $2000$, a learning rate of $1\times 10^{-4}$, gradient clipping of $0.1$ with the Adam optimizer, $120$ parallel environments, and training iterations of $4$M (corresponding to about $60$M environment steps).
+The total training wall-clock time was about 1 week.
+Student High Level.
+We used 60 parallel environments to collect rollout data for training the student, where the data was stored in to a replay buffer with a capacity of $600$k environment steps.
+The full policy was trained on four 3090 GPUs for about 2 weeks, for a total of $1.3$M iterations.
+This amounts to roughly $40$M environment steps in total.
+We used a batch size of $64$, a learning rate of $1\times 10^{-4}$, gradient clipping of $100$, and the Adam optimizer.
+For the mixed rollout schedule, we kept the probability $\beta$ of sampling from the student policy as $0.5$ until the $750$K-th iteration, and linearly increased $\beta$ so that it reaches $1.0$ at the $1.25$M-th iteration.
+The distillation loss weight $\alpha$ in Equation 5 was set to $0.01$.
+The fixed modal dispersion $\sigma$ for distilling the teacher policy was $0.05$.
+Low level.
+We used 5000 parallel environments to roll out and train the low-level policy.
+The policy was trained on one 3090 GPU for about three days for a total of 4B environment steps.
+We used a batch size of 5000.
+The Adam optimizer had an initial learning rate of $2\times 10^{-4}$ and a step decay to its $1/5$ after $70\%$ of the total training progress.
+Appendix D More Sim2Real Techniques
+D-A More Dynamics Gap Reduction Techniques
 Minimizing Tracking Error via PID Control.
@@
 Therefore, we opt to simply use PD position control for the legs.
-C-B Visual Gap Techniques
+D-B More Vision Gap Reduction Techniques
 No Camera Depth Data.
@@
 Color Modeling.
-Visual sim-to-real, especially for colors, is notoriously difficult [40].
+Vision sim-to-real, especially for colors, is notoriously difficult [41].
 We simplify the problem by only using four distinct colors in our experiments, namely, red, green (more like cyan), blue and yellow.
@@
 Figure A.12: Color samples. Each color had 1000 randomly sampled pixels used to estimate the distribution in simulation.
-Appendix D Robot Implementation Details
-D-A Control Latency Management
+Appendix E Robot Implementation Details
+E-A Control Latency Management
 Figure A.13: Asynchronous System Architecture.
@@
     Asynchronous low-high inference: most of the model latency comes from the high-level inference, which takes around 20 ms, but can sometimes peak to 40 ms.
-This could delay the low-level policy (50HZ) considerably and the robot to falter, if the two inferences are performed in sequence.
+This could delay the low-level policy (50HZ) considerably and make the robot falter, if the two levels are executed in sequence.
 Alternatively, in the asynchronous mode, the low-level policy will reuse the previous locomotion command if the new one has not arrived yet, and thus still maintain its 50HZ frequency.
@@
 Low-level control delay is sampled from $(5,7)$ms uniformly per step.
-High-level control delay should have been sampled from $(20,40)$ms, but was accidentally set to none during implementation.
-We did not repeat the full evaluation after discovering this bug, due to limited time.
-We did test introducing an additional 40ms control delay during physical deployment to the existing high-level policies.
-It did not impact performance in any significant way.
-We do plan to train models with the correct delay randomization in the future.
+About the high-level inference delay, we evaluated the policies trained without high-level inference delay and compared the performances with and without test time high-level inference delays.
+We did not find clear performance differences in terms of the success rate.
+Therefore, we did not use high-level control delay in policy training in this work.
+Meanwhile, it is an option that is provided in our implementation and can be turned on whenever necessary.
 Figure A.14: The architecture of the student’s U-Net for predicting segmentation and depth maps.
 Figure A.15: The architecture of the student’s policy input encoder that generates a compact observation embedding to be used by RL.
-D-B Safety Measures for the Physical Robot
+E-B Safety Measures for the Physical Robot
 To reduce damage to the robot itself and the surroundings, we employ many strategies.
@@
     Additionally, to prevent the robot from wandering off and bumping into the surroundings, during training, we incorporate a rotational search behavior as explained in Section A-A during policy training, for all baseline methods and the proposed SLIM method.
-Appendix E Real-World Experiment Details
-E-A Containers and Graspable Objects
+Appendix F Real-World Experiment Details
+F-A Containers and Graspable Objects
 We use cubes of different colors as graspable objects, and baskets of different colors as the target container for dropping into, as show in Figure A.16.
@@
 Figure A.16: Containers and Graspable Objects.
-E-B Real-world Evaluation Protocols
+F-B Real-world Evaluation Protocols
 We detail the protocols used for real-world quantitative evaluation in the main paper (Table I).
-TABLE V: Scene Setups for Real-World Quantitative Evaluations.
+TABLE VI: Scene Setups for Real-World Quantitative Evaluations.
 Legends: $\uparrow$ denotes the position robot with its orientation aligned with the direction of $\uparrow$.
@@
 Similarly, we mark the 4 corner positions on the lobby floor for placing the objects and baskets.
-The size of the square is about $3\times 3$ meters.
+The size of the square is about $2\times 2$ meters.
   4. 4.
-    We follow a scene initialization procedure as detailed in Table V across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.
+    We follow a scene initialization procedure as detailed in Table VI across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.
+F-C Visualization Predicted Segmentation and Depth Maps
+[…]
+Figure A.17: Maps prediction results for some key frames of a deployment session.
+Time flows from top to bottom, and each row corresponds to one time step.
+From left to right, the columns represent the RGB frame, the predicted depth map, the predicted segmentation mask, and our added note for better interpretation of different stages.
+In Figure A.17, we show the robot’s prediction results for depth and object segmentation, during an arbitrary outdoor deployment session.
+We emphasize again that the robot was deployed in this zero-shot scenario never modeled by our simulator, and no model finetuning was performed.
+Despite this, we can see that the prediction results are high-quality, given only a very short history ($0.5$ seconds) of a single wrist-mounted RGB camera and the robot proprioceptive state.
+While searching for the target cube, the perception model had to locate it with just a handful of pixels (first row).
+Right after the gripper holding and lifting the cube above the ground, the model correctly predicted that the target had switched from the cube to a basket (cube’s green color turned red; third row).
+Moreover, while searching for the target basket in the outdoor environment, the cluttered background did not confuse the depth and segmentation prediction (fourth row).
+The perception model was also able to correctly predict the segmentation mask for the robot itself (gripper fingers and quadruped in orange; fifth row).
+Overall, we find that in practice, this high-quality visual information bottleneck not only greatly improves the generalization ability of our RL policy, but also facilitates our debugging of the entire robotic system as it separates the errors of the perception model from those of the policy.

### v3

@@
 We present a low-cost legged mobile manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning purely in simulation.
-This system is made possible by 1) a hierarchical design of a high-level policy for visual-mobile manipulation following instructions and a low-level policy for quadruped movement and limb control, 2) a progressive exploration and learning approach that leverages privileged task decomposition information to train the teacher policy for long-horizon tasks, which will guide an imitation-based student policy for efficient training of the high-level visuomotor policy, and 3) a suite of techniques for minimizing sim-to-real gaps.
-   In contrast to previous approaches that use high-end equipment, our system demonstrates effective performance with more accessible hardware – specifically, a Unitree Go1 quadruped, a WidowX250S arm, and a single wrist-mounted RGB camera – despite the increased challenges of sim-to-real transfer.
-When fully trained in simulation, a single policy autonomously solves long-horizon tasks such as search, move, grasp, and drop-into, achieving nearly 80% success.
-This performance is comparable to that of expert human teleoperation on the same tasks but operates in a more efficient way, at 1.5 times the speed of human expert.
-The sim-to-real transfer is fluid across diverse indoor and outdoor scenes under varying lighting conditions.
-Finally, we discuss the key techniques that enable the entire pipeline, including efficient RL training and sim-to-real, to work effectively for legged mobile manipulation, and present their ablation results.
+This system is made possible by 1) a hierarchical design of a high-level policy for visual-mobile manipulation following task instructions, and a low-level quadruped locomotion policy, 2) a teacher and student training pipeline for the high level, which trains a teacher to tackle long-horizon tasks using privileged task decomposition and target object information, and further train a student for visual-mobile manipulation via RL guided by the teacher’s behavior, and 3) a suite of techniques for minimizing the sim-to-real gap.
+   In contrast to previous works that use high-end equipment, our system demonstrates effective performance with more accessible hardware – specifically, a Unitree Go1 quadruped, a WidowX-250S arm, and a single wrist-mounted RGB camera – despite the increased challenges of sim-to-real transfer.
+Trained fully in simulation, a single policy autonomously solves long-horizon tasks involving search, move to, grasp, transport, and drop into, achieving nearly 80% real world success.
+This performance is comparable to that of expert human teleoperation on the same tasks but significantly more efficient, operating at about 1.5$\times$ the speed.
+Finally, we perform extensive ablations on key techniques for efficient RL training and effective sim-to-real transfer, and demonstrate effective deployment across diverse indoor and outdoor scenes under various lighting conditions.
 Figure 1: SLIM in Real.
-Snapshots of the same SLIM policy deployed in diverse real-world scenes, featuring significant variations in ground types, backgrounds, distractors, and other environmental factors.
+Snapshots of the same SLIM policy deployed in diverse real-world scenes, featuring significant variations in terrain, background, distractors, and other environmental factors.
 These scenes are not available in simulation training.
-The annotations at the bottom are added for understanding task progress and are not part of the input to the system.
+The subtask annotations at the bottom are added for understanding task progress and are not part of the input to the system.
 I Introduction
-Until recently, robots have primarily been limited to static, highly repetitive tasks, often in industrial manufacturing.
-As the demand for autonomy grows, the need emerges for robots to navigate and interact with their environments - commonly referred to as mobile manipulators.
-Such systems can typically handle primitive tasks, such as searching, grasping, and transporting objects, to achieve overarching long-horizon goals.
-There are several promising methodologies for training autonomous mobile manipulators: imitation learning from human demonstrations, reinforcement learning (RL) in the real world, and RL in simulation with sim-to-real transfer.
-Each of them has its pros and cons.
-Imitation learning (IL) benefits from learning directly from human experts teleoperating the hardware, and thus reasonable behaviors and some success in a restricted domain can be achieved even with a handful of trajectories.
-However, diverse behavior data is required to improve the policy’s robustness even for a single task.
-This makes IL costly, especially if we further want to cover all the combinatorial variations of scenes and tasks for robot ability scaling.
-Reinforcement learning (RL) can potentially generate an infinite amount of training data by allowing the robot to explore and interact with the environment.
-Furthermore, the resulting policy is often more optimal than a human teleoperating policy.
-However, directly training an RL policy in the real world can be challenging due to hardware maintenance, reward specification, and safety issues.
-Data scale is also largely limited by the difficulty of deploying a robot fleet for collecting data in parallel.
-As a result, mastering long-horizon tasks in the real world still poses a significant challenge.
-RL in simulation could address the scaling, generalization, and safety issues by running massively parallel simulation with randomized scenes and tasks.
-But to successfully deploy the trained policy on a real robot, the challenge lies in sim-to-real gaps.
-Therefore, we attempt to develop effective and general techniques to bridge the gaps.
-This work focuses on training an end-to-end RL policy in simulation, and zero-shot deploying in the real world on a relatively low-cost quadruped based robot to solve long-horizon, intricate manipulation tasks.
-The complexity of accurate manipulation on legs, the variations of the physical environment, the scene and task setups, and the long task horizon are some of the factors that compound and increase the difficulty of RL training.
-On top of these factors, limited hardware reliability further exacerbates the sim-to-real gap, making this a daunting task.
-We train an end-to-end model that is able to achieve long-horizon tasks consisting of searching, grasping, transporting and dropping primitive tasks (Figure 1).
-For robot mobility, we choose a quadruped instead of a wheeled base, as it has the potential to adapt to many different types of terrains.
-The choice of a quadruped base results in a large 19 DOF (12 leg joints, 6 arm joints, and 1 gripper) system, exacerbating the already difficult long-horizon RL exploration issue.
-To mitigate this, we utilize a hierarchical teacher-student design, consisting of a high-level mobile manipulation policy and low-level quadruped controller, and also a teacher-student training framework [6] for the high-level policy.
-To the best of our knowledge, this is the first demonstration of a low-cost real-world quadruped manipulator fully trained with RL in simulation, for solving long-horizon tasks.
-The teacher policy has access to privileged information (e.g., structured and accurate object features) that is unavailable in the real world.
-To tackle the difficulties of long-horizon task learning, we also incorporate task decomposition as a form of privileged information, where boundaries between subtasks are defined.
-The student policy is then trained to distill the teacher policy while maximizing task rewards, but with access to only standard sensor data such as camera feed and proprioceptive state.
-We also condition the student visuomotor policy on language instructions, providing a natural way to the user for specifying the task to be executed by the robot.
-We name our robotic system SLIM (Sim-to-Real Legged Instructive Manipulation).
-We reiterate that SLIM is a complete robotic system with fluid sim-to-real transfer.
-Unlike prior work [30, 53, 59], SLIM learns all modules in simulation from scratch, without the need for any pretrained model or real-world finetuning.
-SLIM relies on pure RL, and thus it does not require a single demonstration trajectory.
-Our contributions are as follows:
+Legged mobile manipulation combines a robotic manipulator with a legged mobile platform, enabling robots to perform a wide variety of complex tasks in diverse environments [16, 59, 41, 33, 43, 65].
+Unlike stationary or wheeled manipulators, legged systems can adapt to uneven terrains, significantly expanding the scope of potential applications.
+Recent advancements in hardware and algorithms [16, 46, 39] have also made legged manipulators more accessible and cost-effective, encouraging their adoption in research and industry.
+Despite its promise, legged mobile manipulation presents unique challenges.
+The expanded scope of the tasks, the long task horizon, the diverse scenes and terrains, and unstable legged base all compound the difficulty of achieving reliable performance.
+One popular approach to tackle mobile manipulation is Imitation Learning (IL) from expert demonstrations [5, 18, 54, 4].
+It is straightforward, leveraging human expertise to produce reasonable behaviors quickly with supervised training.
+However, IL relies heavily on large and comprehensive datasets to ensure generalization and robustness, making data collection resource-intensive, both in terms of human labor and hardware wear-and-tear.
+Moreover, its performance is upper-bounded by the quality of human demonstrations.
+These limitations become even more pronounced when addressing the complexity of legged manipulation.
+Another paradigm in robotics learning is sim-to-real Reinforcement Learning (RL).
+This method bypasses the need for extensive real-world demonstration data by leveraging virtually infinite data available from simulation, ensuring robust policy training through diverse and comprehensive data [50, 26, 28, 46, 33, 11].
+Furthermore, simulation allows safe exploration without physical wear-and-tear and is not constrained by the performance of demonstrations, unlike IL.
+However, the success of this paradigm hinges on tackling two key challenges: (1) effectively training RL policies in simulation over long horizons, and (2) addressing sim-to-real gaps, including those both in dynamics and vision.
+These challenges are further exacerbated by the additional complexities introduced by legged manipulators.
+Based on sim-to-real RL, we present SLIM (Sim-to-Real Legged Instructive Manipulation), a standalone pipeline for training robotic policies entirely in simulation and deploying them zero-shot in the real world.
+SLIM employs a hierarchical policy structure that decouples task-dependent high-level visuomotor commands from low-level control (Section III-B).
+In high-level policy training, we utilize a teacher-student learning framework to enhance training efficiency, where a teacher policy leverages privileged information to guide the student.
+We train the teacher policy to solve long-horizon tasks via task decomposition and progressive learning (Section IV).
+We then train a student policy, conditioned on language and visual inputs, in RL guided by the teacher’s behavior (Section V).
+Together with a suite of carefully designed sim-to-real techniques (Section VII), including visual and dynamics randomization and low-level controller tuning, SLIM produces visuomotor policies that can maintain sim performance when transferred to real.
+In this work, we implement SLIM on a Unitree Go1 quadruped with a top-mounted WidowX-250S manipulator and a wrist-mounted Intel RealSense D435 camera (only using the RGB stream), resulting in a large 19 degree-of-freedom (DOF) system.
+We focus on the long-horizon task consisting of multi-stage search, move to, grasp, transport, and drop into, achieving nearly 80% real-world success (Section VIII).
+Our key contributions are as follows:
   1. 1.
-    We develop a low-cost quadruped manipulation system SLIM. To our best knowledge, SLIM is the first end-to-end system for solving long-horizon real-world mobile manipulation tasks from RL alone.
+    We develop a low-cost legged manipulation system SLIM.
+To the best of our knowledge, SLIM is the first end-to-end robotic system for solving long-horizon real-world legged manipulation tasks from sim-to-real RL alone.
   2. 2.
-    To address the long-horizon RL exploration challenge, we propose a hierarchical framework for training SLIM’s high-level mobile manipulation policy and low-level quadruped controller, and adopt a teacher-student setup for high-level training efficiency.
+    We conduct an extensive set of real-world experiments with 400 episodes achieving human expert-level success ($\sim$80%) with $\sim$1.5$\times$ the speed of human expert teleoperation.
+We also demonstrate robust task success across a wide range of indoor and outdoor scenarios.
   3. 3.
-    We identify several techniques that are crucial for SLIM to achieve successful sim-to-real transfer with inexpensive hardware, as shown in real-world ablation studies.
-  4. 4.
-    We conduct an extensive set of experiments with real-world evaluations, achieving higher success rate and faster task completion than baselines.
-We also demonstrate robust task success across a wide range of indoor and outdoor environments in real-world.
+    We identify crucial techniques for achieving successful sim-to-real transfer with inexpensive hardware, validated through extensive real-world ablations.
+We plan to open source the code to facilitate future efforts on related research.
 II Related Work
@@
 More recently, there has been explosive progress in using learning-based approaches for achieving quadruped locomotion.
-In particular, sim-to-real reinforcement learning (RL) has arisen as a robust solution showcasing impressive feats such as parkour [61, 8, 4] and speeds of up to 3.9m/s [35].
+In particular, sim-to-real reinforcement learning (RL) has arisen as a robust solution showcasing impressive feats such as parkour [67, 8, 6] and high running speeds [38].
 Other works have also focused on exploiting the strengths of legged locomotion over their wheeled counterparts through utilizing the strong inherent coupling of proprioception with egocentric vision [17, 1] and active estimation of the environment [40, 31, 34].
-Furthermore, minimal tuning is necessary compared to classical approaches [14] with massively parallelized simulations having been shown capable of learning gaits in just minutes [43].
+Furthermore, necessary tuning can be minimal compared to classical approaches [15], for instance, massively parallelized simulation has been shown capable of learning gaits in just minutes [46].
 Finally, data-driven legged locomotion has shown great promise in its ability to zero-shot generalize to new morphologies [14], surfaces [39, 31, 34] and agile skills [56, 30].
@@
 Significant advancements in robust locomotion have further enabled researchers to push the boundaries of legged mobile manipulation.
-Building upon data-driven approaches, quadrupeds have been demonstrated pushing objects with their body [26], dribble around balls [27], and manipulate objects using egocentric [49] or calve mounted grippers [29].
+Building upon data-driven approaches, quadrupeds have been demonstrated pushing objects with their body [29], dribbling around balls [30], and manipulating objects using egocentric [55] or calve mounted grippers [32].
 Still, the most popular setup has been the traditional top-mounted 6DOF manipulator design [16, 20, 33, 65, 59, 37, 41, 48].
@@
 Classical approaches to solving top-mounted legged mobile manipulation consist of designing wholebody controllers [48] or even combining locomotion learning with model-based manipulation control [37].
-Others simplify the need for low-level controllers by using high-level API calls ($\dot{x},\dot{\theta}$) provided by their respective quadruped robots to achieve graspability-aware policies [59], navigational pick-and-place [53], and language-conditioned mobile manipulation [40].
+Others simplify the need for low-level controllers by using high-level API calls provided by their respective quadruped robots to achieve graspability-aware policies [65], navigational pick-and-place [59], and language-conditioned mobile manipulation [43].
 Complex tasks have also been achieved by leveraging teacher-student training setups in simulation for grasping [33] and door opening [66].
 Finally, to expand the robot’s workspace even further, wholebody loco-manipulation has arisen where leg joints are actuated in a way to assist manipulation [16, 33, 41].
-Most similar to our work, Liu et al.
-[30] introduced VBC, which trains a sim-to-real visuomotor whole-body loco-manipulation policy using a high-low hierarchical model and teacher-student training setup for the object pickup task.
-They achieve impressive results across a diverse range of objects by leveraging whole-body manipulation, though this approach comes with several key caveats.
-First, VBC requires the user to manually click on the target object to perform an initial segmentation using a third-party general-purpose model.
-This segmentation is then used to track the object through another third-party model.
-Consequently, to ensure consistent segmentation, the object must always remain within the frame.
-This requirement inherently limits the feasible starting state distribution of VBC, as the target object must be placed within a short distance in front of the robot and in view.
-Finally, due to the close proximity and visibility of the object, the VBC task horizon is relatively short, ending as soon as the object is lifted.
-In comparison,
+Most closely related to our work, Liu et al.
+[33] introduces a visual whole-body control approach (VBC), which trains a sim-to-real visuomotor whole-body loco-manipulation policy using a high-low hierarchical model and teacher-student training setup for the object pickup task, similar to our design.
+The system can grasp an impressive set of diverse objects by leveraging whole-body manipulation with capable hardware.
+VBC takes a few shortcuts to simplify the task.
+VBC requires the user to manually click on the target object to perform initial segmentation using a third-party vision model.
+This segmentation is then used to track the object through a third-party tracking model, which requires the object to stay within the camera view the whole time.
+Additionally, VBC uses a scripted policy to put the object into the basket on the back of the quadruped, once the object is lifted, making the task scope more limited and the task horizon much shorter.
+In comparison, SLIM offers the follow advantages:
   1. 1.
     SLIM is a complete and self-contained system.
-All modules–vision, high-level visual loco-manipulation, low-level motor control–are trained within a single framework, resulting in both minimal latency and full autonomy.
+All modules–vision, high-level visual policy, low-level motor control–are trained in our simulation environment integrated within a single framework without any third party modules, resulting in a lean 19 million parameter neural network system, allowing both minimal latency and full autonomy.
   2. 2.
-    SLIM is fluid and intuitive. Conditioned on language, SLIM is able to rapidly accomplish tasks in a wide variety of scenes, requiring just a user language command.
+    SLIM is fluid and intuitive to use. Given a language instruction, SLIM is able to rapidly accomplish tasks in a wide variety of scenes.
   3. 3.
-    Finally, SLIM is practical. With SLIM, we demonstrate impressive real-world, long-horizon task completion for a multi-stage mobile search and manipulation task.
+    Finally, SLIM is practical.
+SLIM can accomplish complex long-horizon real-world tasks autonomously from the beginning to end, as apposed to some of the existing methods that requires human involvement (e.g.
+manually clicking for segmentation) and can only automate part of the task with restrictions on the feasible starting state distribution [33]
 III Architecture Overview
@@
 Figure 2: Hierarchical Framework and Visuomotor Policy Pipeline (Left).
-Given a language instruction and various sensor information, the policy generates two control signals: 1) arm joint position targets and 2) a locomotion command.
-The first control signal is directly passed to the arm driver (passthrough), while the second signal is passed into a low-level locomotion policy to generate leg joint position targets.
-Low-Level and High-Level Teacher-Student Policy Training (Right).
-SLIM’s training pipeline can be decomposed into the following three sequential stages of policy training.
-First, a low-level locomotion policy is trained to track a specified locomotion command via RL.
-Next, a high-level teacher policy is trained with privileged information via RL to solve the long-horizon task.
-Finally, the student visuomotor policy is trained by distilling the teacher policy while maximizing task rewards, conditioned on visual inputs.
-Both the teacher and student policies use the same frozen locomotion policy from the first stage to execute their locomotion commands.
+Given language instruction and sensor inputs, the high-level policy generates a set of two control signals: 1) the arm and gripper control signals, and 2) the locomotion command.
+The arm control signals are directly passed to the arm driver, and the locomotion command is passed to the low-level policy to control the leg joints of the quadruped.
+Low-Level Training (Top Right), and High-Level Teacher and Student Training (Middle and Bottom Right).
+SLIM training is divided into three sequential stages.
+First, the low-level locomotion policy is trained via RL to follow a sampled linear and angular velocity command.
+Second, the high-level teacher policy is trained via RL with privileged low-dimensional state input to solve the long-horizon task.
+Finally, the student visuomotor policy is trained by distilling the teacher behavior while maximizing task rewards, using visual, sensory, and language instruction as input.
+Both the teacher and student policies command the same frozen low-level policy produced by the first stage.
+The teacher is only run in simulation and the student can be deployed in real.
+Fig.
+2 illustrates the hierarchical (high-low policy) framework (left) and the teacher-student training pipeline (right) used in SLIM.
+Given language instruction and sensor inputs, the high-level policy generates intermediate control signals that will be used as input to the low-level policy.
+A low-level locomotion policy is trained to track a specified locomotion command via RL and is frozen after training.
+Then a high-level teacher policy is trained with privileged information via RL to solve the long-horizon task, and a student visuomotor policy is trained by distilling the teacher policy while maximizing task rewards, conditioned on visual inputs and language.
 III-A Observations and Actions
@@
     A temporal stack of RGB images from a single wrist-mounted camera with a shape of $(N,H,W,3)$, where $N$ is the stack size, and $H$ and $W$ are the image height and width, respectively.
-The environment is always partially observed by the robot.
+The environment is always partially observed by the robot, and most of the workspace will not be contained in this camera view during the task.
   • $\mathbf{s}_{\textrm{r}}$
-    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of proprioceptive reading of all joints.
-$\mathbf{s}_{\textrm{r}}\!=\![\mathbf{s}_{\textrm{leg}},\mathbf{s}_{\textrm{arm}}]$, where $\mathbf{s}_{\textrm{leg}}$ denotes the leg-related proprioceptive states and $s_{\textrm{arm}}$ the arm-related proprioceptive states.
-When used as the input to the low-level locomotion policy, we overload $\mathbf{s}_{\textrm{leg}}$ to denote the leg-related proprioceptive states stacked according to the low-level frame-stacking size.
+    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of proprioceptive readings of all joints.
+$\mathbf{s}_{\textrm{r}}\!=\![\mathbf{s}_{\textrm{leg}},\mathbf{s}_{\textrm{arm}}]$, where $\mathbf{s}_{\textrm{leg}}$ denotes the leg-related proprioceptive state (leg joint positions) and $s_{\textrm{arm}}$ the arm proprioceptive state (arm joint positions).
 Privileged Observations (only available in simulation).
@@
   • $\mathbf{s}_{\textrm{p}}$
-    A group of temporal object feature stacks, where each stack has a shape of $(N,Q_{m})$ and $0\leq m<M$ denotes the feature index.
+    A group of temporal stacks of privileged object features (e.g., positions, scales, categories, etc.), where each stack has a shape of $(N,Q_{m})$ and $0\leq m<M$ denotes the object index.
 $M$ is the maximal number of task related objects in the scene.
 Note that object features are also partially observed and their visibility is always determined by the robot’s camera field of view.
+The features are set to zero if they are not in the view.
+This is to ensure that the object visibility is consistent between the guidance policy and the robot policy, and avoid any information gap for policy distillation.
 The privileged information can be extracted or computed from the underlying simulator state.
@@
   • $\mathbf{a}_{\textrm{arm}}$
-    A vector in $[-z,z]^{6}$, indicating the delta changes to the arm joint positions. $z$ is the maximally allowed joint position change for a control interval.
+    A vector in $[-z,z]^{I}$, indicating the delta changes to the arm joint positions.
+$z$ is the maximally allowed joint position change for a control interval.
+$I$ is the number of actuated arm joints, excluding the gripper joint.
   • $g$
@@
   • $\mathbf{a}_{\textrm{leg}}$
-    The target quadruped joint positions in $\mathbb{R}^{12}$.
-In total, a complete action $\mathbf{a}$ has 19 dimensions.
+    The target quadruped joint positions in $\mathbb{R}^{J}$. $J$ is the number of actuated leg joints.
+In total, a complete action $\mathbf{a}$ has $(I+1+J)$ dimensions.
 III-B Hierarchical Policy Structure
-The system as a whole takes visual and instructional input and controls all joints of the quadruped and the arm, which is a daunting training task with high-dimensional input and output.
-We use a two-level hierarchical policy to divide the complexity of visual mobile manipulation training from legged locomotion training, similar to [30], illustrated in the left half of Fig.
-2.
-The high level takes visual, instructional, and proprioceptive inputs, and outputs a simple locomotion speed command and arm control command, while the low level takes proprioceptive input and outputs leg joint commands to track the locomotion command.
-The assumption is that quadruped locomotion control is largely independent of the high-level task semantics.
-Intermediary Action.
-To enable this two-level policy structure, we introduce an intermediary action as follows,
+The overall system takes visual and instructional input and controls all joints of the quadruped and the arm, which involves a non-trivial training task with high-dimensional input and output.
+We use a two-level hierarchical policy to divide the complexity of visual mobile manipulation training from legged locomotion training, illustrated in the left part of Fig.
+2, similar to [33].
+The high level policy takes visual, instructional, and proprioceptive inputs, and outputs a locomotion speed command and arm control command.
+The low level policy tracks the locomotion command from the high level, using proprioceptive inputs to command the quadruped’s leg joints.
+The assumption is that given the locomotion command, quadruped locomotion control is largely independent of the high-level task semantics.
+Intermediate Action.
+The following intermediate action separates the high and low-level policies,
   • $\mathbf{c}$
@@
 With this, a high-level policy is defined as
-$(\mathbf{l},\mathbf{o},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a}_{\text{hi}}\triangleq(\mathbf{a_{\textrm{arm}}},g,\mathbf{c})\in\mathbb{R}^{9},$
+$(\mathbf{l},\mathbf{o},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a}_{\text{hi}}\triangleq(\underbrace{[\mathbf{a_{\textrm{arm}}},g]}_{\textrm{arm and gripper control}},\underbrace{{\mathbf{{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}[}c{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}]}}}}_{\textrm{locomotion command}})\in\mathbb{R}^{I+3},$
 while the low-level policy is defined as
-$(\mathbf{c},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a_{\textrm{leg}}}\in\mathbb{R}^{12}.$
-Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$, thus does not need to handle high dimensional image and language inputs.
-Low-Level.
-The low-level policy is a quadruped controller that generates joint position targets for PD control to follow a task-agnostic 2D command $\mathbf{c}$ [36].
-The low-level policy is trained in simulation using PPO [44] and transferred to real.
-An illustration of the training can be found in the first block in the right column of Fig.
+$(\mathbf{c},\mathbf{s_{\textrm{leg}}})\rightarrow\mathbf{a_{\textrm{leg}}}\in\mathbb{R}^{J}.$
+Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$ and thus, does not need to process high-dimensional image and language inputs.
+Low Level.
+The low-level policy is a quadruped controller that generates joint position targets for PD control to follow a task-agnostic 2D linear and angular-velocity command $\mathbf{c}$ [39] and is trained in simulation using PPO [47].
+An illustration of the training can be found in the top row of the right column of Fig.
 2.
-It is worth noting that we randomly sample both command $\mathbf{c}$ and arm joint actions $\mathbf{a_{\textrm{arm}}}$ during low-level training.
-Since the arm mounted on the quadruped can have varying poses during the task, the low-level command following ability must generalize to various arm configurations.
+It is worth noting that during low-level training, the same embodiment as shown in Fig.
+2 (arm mounted on top of the quadruped) is used in low-level training.
+We randomly sample both command $\mathbf{c}$ and arm joint actions $\mathbf{a_{\textrm{arm}}}$ during training.
+Since the arm mounted on the quadruped can have varying poses during the task, this enforces the low-level command following ability to generalize to various arm configurations.
 That is, the quadruped base has to learn to keep balance while achieving locomotion commands regardless of the arm’s current joint positions and movements.
-Similar to most sim2real approaches, we also randomize the simulation environment during training.
-Besides widely-used domain randomization parameters as in [15], we further randomize the simulated delay of each sensor and foot softness of the robot, to better adapt to variations in real deployment.
-Rewards design and full details of the training setup can be found in Appendix A-B.
-High-Level.
-The high-level policy is responsible for outputting the command $\mathbf{c}$ and the arm control $(\mathbf{a_{\textrm{arm}}},g)$, conditioned on language $\mathbf{l}$ and the current stack of RGB images $\mathbf{o}$.
-As demonstrated in Fig.
-2, the command $\mathbf{c}$ is forwarded to the low-level policy which then follows the received command for a number of time steps, before the high-level policy outputs the next command.
-For the manipulator, we opt to operate in joint space (as opposed to task space) and thus no inverse kinematics is needed.
-Once the low-level policy is trained, it is frozen and used as a base controller by the high-level policy.
-In the remainder of this paper, we will only talk about the high-level policy unless otherwise stated.
-III-C Teacher-Student Training
-Given the challenges of an enormous exploration space implied by long-horizon tasks, we adopt a teacher-student learning framework  [11, 30, 60] for better high-level training efficiency.
-The overall structure of the framework is shown in the second and third blocks of the right side of Fig.
+We intentionally remove arm joint state and joint command from low level input, see details in Appendix B-C.
+Similar to most sim-to-real approaches, we also randomize the simulation environment during training.
+Besides widely-used domain randomization parameters as in [16], we further randomize the simulated delay of each sensor and foot softness of the robot to better adapt to variations in real deployment.
+Full details of the training setup and reward design can be found in Appendix A-C.
+High Level.
+The high-level policy is responsible for outputting the locomotion command $\mathbf{c}$ and the arm controls $(\mathbf{a_{\textrm{arm}}},g)$, given language instruction $\mathbf{l}$, proprioceptive state $\mathbf{s}_{\textrm{r}}$, and the current stack of RGB observations $\mathbf{o}$.
+As shown in Fig.
+2, the command $\mathbf{c}$ is forwarded to the low-level policy, which then follows the received command for a number of time steps, before the high-level policy outputs the next command.
+For the manipulator, we opt to operate in joint space, as opposed to task space or end-effector pose, so we do not have to worry about inverse kinematics computation.
+Once the low-level policy is trained, it is frozen and used as a base controller by the high-level (teacher or student) policies.
+In the remainder of main text, we only talk about the high-level policies unless otherwise stated.
+III-C Teacher-Student For High-Level Training
+With the low level policy taking over the responsibilities of legged locomotion control, high level can focus on task dependent decision making.
+However, the high level policy needs to process high dimensional visual and language input, and complete the long-horizon task.
+This presents an enormous space for the policy to explore, and a challenge for efficient RL training.
+We adopt a teacher-student learning framework  [12, 33, 66] for efficient high-level training.
+The overall structure of the framework is shown in the middle and bottom right sections of Fig.
 2.
-As shown, the teacher is trained purely with RL from privileged, structured, and low-dimensional environment data $(k,\mathbf{s_{\textrm{p}}},\mathbf{s_{\textrm{r}}})$ that can only be obtained from the simulator.
-After a high performing teacher policy is learned, it is frozen and used to guide the student’s policy learning via a distillation loss.
-As the student will eventually be deployed in the real world, its inputs are consistent with what our robot system has.
-Besides policy distillation, we also use an RL objective to allow the student’s behavior being shaped by the same set of task rewards used by the teacher.
-Later, we will show improvements of this distillation-guided RL setting, compared to pure policy distillation (Section VIII-B Distillation Only baseline.).
-In the next two sections, we will provide more technical details of the teacher and student.
-IV The Teacher: Progressive Policy Expansion for Long-Horizon Task Learning
-IV-A Long-Horizon Task Learning Challenges
-Since we are targeting at solving long-horizon tasks via RL, we need an effective approach to train a policy to solve long-horizon tasks successfully.
-There are are few key challenges that must be addressed for solving the long horizon tasks:
+As shown, the teacher is trained purely with RL from privileged, structured, and low-dimensional inputs $(k,\mathbf{s_{\textrm{p}}},\mathbf{s_{\textrm{r}}})$ some of which can be obtained from the simulator but not easily in the real world.
+When a teacher policy is successfully learned, it is frozen and used to guide the student via behavior distillation.
+As the student will eventually be deployed in the real world, its inputs no longer contain privileged information.
+Besides policy distillation, we also use an RL objective to allow the student’s behavior to be shaped by the same set of task rewards used by the teacher, for a potentially more effective policy than using distillation alone (details in Section VIII-B, Distillation Only baseline).
+In the two sections below, we explain the teacher and the student policies in more detail.
+IV The Teacher: Long-Horizon RL with Task Decomposition and Policy Expansion
+The Teacher works with privileged object state information, thus, offloading the complexities of visual representation learning to the student, and only needs to focus on the following challenges of long-horizon task learning,
   1. 1.
-    Continual exploration: When learning a long-horizon task via RL, there may be a number of intermediate milestones that must be sequentially achieved.
-Therefore, the teacher must be able to keep exploring at its frontier despite solving an intermediate milestone stage, in pursuit of solving the full task.
-Without the ability of continual exploration, it is more likely to converge to a sub-optimal solution that solves the task only up to some intermediate milestones, therefore cannot solve the full task completely.
+    Continual exploration: For long-horizon tasks, there may be a number of intermediate milestones that must be sequentially achieved.
+Therefore, even after reaching an intermediate milestone, the teacher must continue exploring new frontiers in order to ultimately solve the entire task.
+Without carefully encouraging continual exploration, the agent can stop exploration early and settle on a suboptimal solution.
   2. 2.
-    Loss of plasticity and catastrophic forgetting: As the teacher progresses into a task, it needs to properly cope with both the loss of capacity/plasticity [32, 33] of the network due to the evolving nature of the problem as well as the data imbalance during long-horizon RL task exploration and training, without impacting skills that have already been acquired (catastrophic forgetting).
-IV-B Progressive Policy Expansion
-Given the challenges listed above, we build upon the idea of Policy Expansion (PEX) [58] originally used in the context of offline-to-online RL.
-Intuitively, PEX divides the full learning process into stages (e.g.
-offline and online stages in [58]) and uses dedicated networks for learning within each stage.
-This way, the trainer allocates new network capacity for continual exploration and learning of new skills without affecting any existing skill (catastrophic forgetting).
-Here we generalize this idea for addressing the challenges in long-horizon task learning as mentioned in Section IV-A, by extending it to multiple stages.
-More concretely, we partition the full long horizon task into a sequence of stages (subtasks), and then perform a policy expansion operation for each new encountered subtask.
-By doing this, it can effectively combat the challenges in long-horizon task learning, keep exploring beyond the currently solved subtasks to solve the subsequent subtasks, circumvent the loss of plasticity by leveraging the newly allocated policy network  [^1] We use the term policy network to represent all the networks that are required for learning a policy.
-For example, in the context of Actor-Critic formulation, it encompasses both the actor and critic networks.
-for learning, and avoid catastrophic forgetting by retaining the policy networks for previous subtasks.
-Given these progressive multi-stage expansions, we dub this strategy Progressive Policy Expansion.
-Formally, we represent the full long-horizon task $\mathcal{T}$ as a composition of a set of subtasks
-$\mathcal{T}=\mathcal{C}(\{\tau^{k}\}_{k=1}^{K})$ (1)
-where $\mathcal{C}$ is a composition operator, assembling the set of subtasks $\{\tau^{k}\}$ into the full task $\mathcal{T}$.
-$K$ denotes the maximum number of subtasks.
-We also define a task decomposional operator as follows, mapping an input state to the corresponding subtask index:
-$k=\mathcal{D}(\mathbf{s})\quad k\in\{1,2,\cdots,K\},$ (2)
-where $\mathbf{s}=[\mathbf{s}_{\textrm{r}},\mathbf{s}_{\textrm{p}}]$.
-We use $\pi_{\rm tea}^{k}(\mathbf{a}|\mathbf{s})$ to denote the teacher sub-network that is responsible for solving its corresponding subtask $\tau^{k}$.
-When there is no confusion, we omit the subscript of the teacher network $\pi_{\rm tea}^{k}$ as $\pi^{k}$.
-Figure 3 provides graphical illustrations of the teacher sub-network structures.
-Each sub-network (Figure 3 left) takes low-dimensional proprioceptive and privileged observation as input and is responsible for learning to solve the corresponding subtask.
+    Loss of capacity and catastrophic forgetting: As the teacher progresses in training, it needs to cope with both the loss of capacity issue [35, 36], which hinders the network from continual learning, and the catastrophic forgetting issue, which could destroy skills that have already been acquired during the learning of new tasks.
+Addressing these challenges are key for successful teacher policy learning. We propose to use task decomposition and further integrate with policy expansion (PEX) [64] to address these two challenges.
+We first decompose the long horizon task $\mathcal{T}$ into $K$ subtasks $\{\tau^{k}\}_{k=1}^{K}$ with shorter-horizons.
+To incorporate the task decompositional structure in learning, we leverage the fact that privileged information is accessible to the Teacher and thus include the subtask index $k$ as part of the privileged observation.
+Subtask successes keep pushing the agent to learn new subtasks one after another, addressing the exploration issue.
+Secondly, we solve the loss of plasticity and forgetting in long horizon task learning via a progressive application of the PEX scheme [64].
+Originally designed for two-stage offline-to-online training, PEX uses dedicated networks for learning within each stage.
+This way, the trainer creates new network instances for continual exploration and learning of new skills without affecting any skills acquired in the previous stage.
+Here we apply this idea progressively to the decomposed long-horizon task with multiple subtasks, where the stages corresponds to the subtasks.
+We dub this strategy as Progressive PEX.
+Because of policy expansion at each stage transition in Progressive PEX, it addresses the two challenges brought by the long-horizon task learning naturally: 1) it can achieve multi-stage continual exploration and the full ability of exploration is always ensured when entering a new stage; 2) since the policy for solving one subtask is encapsuled in a dedicated network, it mitigates the catastrophic forgetting issue for the already learned subtask policies.
+For a newly encountered subtask, the dedicated network that will be newly allocated addresses the issue of lost plasticity.
 […]
 Figure 3: Teacher Policy Network Structure.
+The teacher network is a set of identical submodules $\{\pi^{k}\}_{k=1}^{K}$ gated by the subtask id $k$ (right side of figure).
+Each submodule $\pi^{k}$ is only trained with the experience when the agent is working on subtask $k$, taking privileged and proprioceptive input, flattening and concatenating them before passing it through an MLP (left side of figure).
 On the left, each individual teacher policy $\pi^{k}$ takes a set of privileged and proprioceptive information as input, flattening and concatenating them before passing it through an MLP.
-In the case of actor-critic learning, both critic and actor networks have the same architecture, with the difference in that the actor network outputs an action distribution while the critic outputs a scalar value.
 On the right, the full teacher network $\Pi$ is composed of multiple structurally identical networks $\{\pi^{k}\}_{k=1}^{K}$.
-Given an input id $k$, the $k$-th component network is activated in the computation, i.e.
-$\Pi[k]\equiv\pi^{k}$.
-We can then solve the long horizon task $\mathcal{T}$ via Progressive PEX as follows.
-We initiate the exploration and learning with a single policy network $\Pi=\{\pi^{1}\}$.
-This policy network is responsible for learning to solve the initial subtask.
-Whenever a new subtask is encountered, a policy expansion operation is conducted, adding a new policy into the policy set $\Pi=\{\pi^{1},\pi^{2}\}$.
-By doing this progressively for the full task, we get
+Given an input id $k$, the $k$-th component network is activated in the computation, i.e., $\Pi[k]\equiv\pi^{k}$.
+We use $\pi_{\rm tea}^{k}(\mathbf{a}_{\text{hi}}|\mathbf{s})$ to denote the teacher sub-network that is responsible for solving its corresponding subtask $\tau^{k}$.
+When there is no confusion, we omit the subscript of the teacher network $\pi_{\rm tea}^{k}$ as $\pi^{k}$.
+The left side of Fig.
+3 provides a graphical illustration of the teacher sub-network structures.
+Each sub-network takes low-dimensional proprioceptive and privileged observation as input and is responsible for learning to solve the corresponding subtask.
+Progressive PEX works as follows.
+We initiate the exploration and learning with a single policy network [^1] We use the term policy network to represent all the networks that are required for learning a policy.
+For example, in the context of Actor-Critic formulation, it encompasses both the actor and critic networks.
+$\Pi=\{\pi^{1}\}$ that is responsible for learning to solve the initial subtask.
+Whenever a new subtask is encountered, a policy expansion operation is conducted, adding a new policy into the policy set, i.e., $\Pi=\{\pi^{1},\pi^{2}\}$.
+By doing this progressively for the all the $K$ subtasks, we get
 $\Pi$$\triangleq\{\pi^{k}\}_{k=1}^{K},$ (1)
-as illustrated in Figure 3 right.
-A forward pass through $\Pi$ is carried out via the proper sub-policy by indexing the policy set with the task decompositional operator $\mathcal{D}$ conditioned on the current state
-$\Pi(\mathbf{a}_{\text{hi}}|\mathcal{D},\mathbf{s})$ $=\pi^{\mathcal{D}(\mathbf{s})}(\mathbf{a}_{\text{hi}}|\mathbf{s}).$ (4)
-There are many possible ways to implement the task decompositional operator $\mathcal{D}$.
-In our case, since it is used in teacher policy learning, we can leverage the fact that privileged observation can be used by the teacher policy and implement it as part of the privileged observation provided by the simulator, e.g., providing the subtask index $k$ directly in the privileged observation.
-Based this implementation, the training of $\Pi$ is closely related to multi-task RL [56, 46], where we sample a batch of multi-task-transitions (transitions containing the task index $k$) from the replay buffer and use each transition in the batch for training the sub-network associated with its corresponding task index.
-In this work, we train the teacher policy using a multi-task variant of SAC algorithm [19, 56, 46],
+as illustrated in the right side of Fig. 3.
+Based on this implementation, the training of $\Pi$ is closely related to multi-task RL [62, 49].
+For each training iteration, we get a batch of samples containing the subtask indices from the replay buffer and use each sample in the batch for training the sub-network associated with the subtask index of that sample.
+In this work, we train the teacher policy using a multi-task variant of SAC algorithm [21, 62, 49].
+The performance comparison between the method using a standard network structure and the Progressive PEX approach on solving long-horizon tasks is provided in Appendix J-B, which clearly shows Progressive PEX has a much stronger ability in long-horizon task learning.
 V The Student: Policy Distillation Guided RL
-The student needs to perceive the surrounding environment through RGB images.
-There is no input regarding which subtask the student is currently solving, and it has to infer this information from the inputs.
-Because of this, the student learns only a single policy across the entire long-horizon task (no sub-task dedicated policy).
-To tackle these challenges, we train the student by distilling the teacher’s multiple sub-task policies into a single task policy $\pi_{\rm stu}$.
-However, distillation will only give us a policy that works fine but not quite well.
-It could lead to the state drifting or out-of-distribution (OOD) issues when the policy is rolled out for a long horizon.
-There is also always an observation gap between the teacher and the student, which sometimes creates a difficulty in minimizing the distillation loss.
-Moreover, due to the differences in model architectures and inductive biases, the student might need a different way of mastering certain skills.
-Finally, distillation makes the student’s performance upper bounded by the teacher’s.
-Given these considerations, we choose to boost the student’s policy with RL under the same set of task rewards used by the teacher (Fig.
+The student needs to perceive the surrounding environment using RGB, motor sensors and language instruction.
+Because the student needs to be deployed in the real world, it cannot use any privileged information e.g.
+which subtask it needs to solve.
+Therefore, the student has to learn a single policy over the entire long-horizon task.
+To tackle these challenges, we train the student by distilling the teacher’s behavior into a single task policy $\pi_{\rm stu}$.
+However, distillation on its own is likely to produce a policy that works reasonably well but it could lead to state drifting / out-of-distribution (OOD) issues when the policy is rolled out over a long horizon.
+There is also always an observation gap between the teacher and student, which may create difficulty in minimizing the distillation loss.
+Moreover, due to differences in model architectures and inductive biases, the student may need some adaptation of the teacher’s skills.
+Given these considerations, we choose to boost the student’s policy with RL under the same set of task rewards used to train the teacher (Fig.
 4).
@@
 First, we use a mixed rollout strategy to generate replay data.
-When rolling out a new episode, with a probability of $\beta$ we will sample from the student policy, or otherwise from the teacher policy.
-On one hand, we wish to use the teacher to efficiently generate high-performing transitions along the long horizon.
+When rolling out a new episode, with a probability of $\beta$, we will sample actions from the student policy, otherwise from the teacher policy.
+On the one hand, we wish to use the teacher to efficiently generate high-performing transitions along the long horizon.
 On the other hand, we want to keep exploring with the student’s own policy.
-Second, following SACLite [55] we remove the entropy reward from policy evaluation.
-Then for policy improvement, we replace the entropy term with the distillation loss, and assign a fixed weight $\alpha$ to it:
-$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{o},\mathbf{s}_{r},\mathbf{s}_{p},l,k)\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)}Q\big{(}(\mathbf{o},\mathbf{s}_{r},l),\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha KL[\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{r},\mathbf{s}_{p})||\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)]\Big{]}.\\ \end{array}$ (5)
-To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform a policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion $\sigma$ (e.g., std.
+Second, following SACLite [61], we remove the entropy reward from policy evaluation.
+Then for policy improvement, we replace the entropy term with the distillation loss and assign a fixed weight $\alpha$ to it:
+$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{s}_{\textrm{stu}},\mathbf{s}_{\textrm{p}},k)\sim\mathcal{D}_{\text{replay}}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{s}_{\textrm{stu}})}Q\big{(}\mathbf{s}_{\textrm{stu}},\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha{\rm KL}\!\left(\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{\textrm{r}},\mathbf{s}_{\textrm{p}})\,||\,\pi_{\rm stu}(\cdot|\mathbf{s}_{\textrm{stu}})\right)\Big{]},\\ \end{array}$ (2)
+where $\mathbf{s}_{\textrm{stu}}=[\mathbf{o},\mathbf{s}_{\textrm{r}},\mathbf{l}]$.
+To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion $\sigma$ (e.g., std.
 for Gaussian).
-This makes our distillation loss the best of two worlds: imitation learning and entropy regularization.
+This results in a distillation loss that elegantly achieves two goals: imitation learning and entropy regularization.
 […]
-Figure 4: A close look at the student’s architecture.
-Red dashed lines represent data flows that only exist for training in simulation.
-Solid lines denote the inference process of the student policy.
-For deployment, the three inputs are replaced with real-world observations.
-When designing the student’s representation model, we specifically take vision sim2real gap reduction into consideration, because once after the student is trained, it will be directly deployed in real without any finetuning.
-While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck to reduce the vision sim2real gap (Fig.
-4).
+Figure 4: Student Architecture Overview. Red dashed lines represent data flows that only exist in simulation training. Solid lines denote the inference process of the student policy.
+When designing the student’s representation model, we specifically take visual sim-to-real gap reduction into consideration, because once the student is trained, it will be directly deployed in real without any finetuning.
+While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck (Fig.
+4) to reduce the visual sim-to-real gap while achieving better interpretability.
 The idea is to let the policy depend on as minimal visual information as possible, namely, to limit the complexity of its input vision space.
-This potentially leads to a better generalization ability.
+This can leads to potentially better generalization.
 While there might be many choices for the vision bottleneck, from discrete latent codes [22] to canonical images [28], we choose to use a pair of segmentation and depth maps.
-There are mainly two reasons for this choice.
-First, we believe that they contain the minimal visual information required to achieve our task, while without incurring much task information loss.
-Second, our simulator already provides groundtruth segmentation and depth maps, as high-quality supervision for learning such a vision bottleneck.
+There are mainly three reasons for this choice.
+First, we believe that they contain the minimal visual information required to achieve our task without incurring much task-relevant information loss.
+Second, ground-truth segmentation and depth maps can be efficiently obtained from the simulator for high-quality supervised learning of such a vision bottleneck.
 Thus when training the student, we add a supervision loss for predicting the segmentation mask and depth map from its inputs.
@@
 Using a segmentation mask for selecting interesting objects is similarly adopted by prior works [33, 59].
+Finally, this design leads to interpretable visual representations, which is helpful for us to diagnose the visual sim-to-real gap.
 VI Task Implementation In Simulation
-Figure 5: Task Decomposition as Privileged Information.
+Figure 5: Task Decomposition as Privileged Information (for teacher only).
 Given a task, we decompose it into a sequential series of primitive subtasks.
-As shown above, the task of finding an object A, grasping it, and then dropping into an object B can be decomposed into a series of subtasks.
+As shown above, the task of finding an object A, grasping it, and then dropping into an object B can be decomposed into a sequence of subtasks.
 In simulation, the subtask boundaries and transitions are well-defined, serving as a form of privileged information to the teacher.
-The teacher can leverage this information, and use a separate policy to learn each subtask, thereby conquering the long horizon task.
+The teacher can leverage this information, and use a separate policy to learn each subtask.
 This expertise is then distilled into the student, which is conditioned on both visual camera feed and language commands.
-We use a long-horizon pick-and-place task as a concrete example in this work (c.f.
-Figure 2 left for an illustration of the task), and we will use this task as a concrete example for illustrating the task decomposition as well.
-Given a language instruction (e.g., “Drop the blue cube into the green basket.”), the robot needs to locate the target object, picking it up and then drop it into the target basket.
-As mentioned previously in Section IV, SLIM employs a task decompositional approach paired with progressive PEX for training a teacher policy to solve the long horizon task.
-More concretely, for the long-horizon task under consideration, it can be naturally broken down into the following sequential steps (c.f.
-Figure 5):
+As mentioned previously in Section IV, SLIM employs a task decompositional approach paired with progressive PEX for training a teacher policy to solve the long-horizon task.
+In this work, we use a multi-stage search and pick-and-place task as our exemplar long-horizon task (Fig.
+A.16).
+This task will also serve as a concrete example for illustrating the task decomposition.
+Given a language instruction (e.g.,Drop the blue cube into the green basket.”), the robot needs to locate an object A (the blue cube), pick it up, and then drop it into a container B (the green basket).
+This task can be naturally broken down into the following sequential steps, where object A and B are specified by the language command itself:
   1. 1.
-    Search: the robot should first search for the target object (the green cube) for picking up (a subtask), based on the RGB images from the gripper mounted ego-centric camera;
+    Search: The robot must first search for object A using the RGB images from the wrist-mounted ego-centric camera.
   2. 2.
-    MoveTo: after finding the target object, the robot moves to the object until it is within the reach of the arm;
+    MoveTo: After locating object A, the robot should approach it until it is within the reach of the arm.
   3. 3.
-    Grasp: then the robot moves the gripper towards the target object for picking it up.
-During the grasping process, the robot base should ideally move less compared to the MoveTo subtask to facilitate accurate grasping;
+    Grasp: The robot attempts to grasp and pick up object A.
   4. 4.
-    SearchWObj: after picking up the target object, the robot should keep it in the gripper and then search for the target basket; within this stage, there should be some coordination between arm and locomotion movement to facilitate search while avoid the cube to be dropped.
+    SearchWithObj: After picking up object A, the robot should then search for container B.
+Within this stage and those after, there should be some coordination between the arm and locomotion movement to facilitate subtask success while avoiding dropping the grasped object.
   5. 5.
-    MoveToWObj: after finding the target basket, the robot should move towards it while keeping the object in gripper, until the target basket is within a range that is appropriate for the robot to reach for dropping the object in its gripper;
+    MoveToWithObj: After locating container B, the robot should approach it until it is within the reach of the arm.
   6. 6.
-    MoveGripperToWObj: the robot can now move the gripper towards the target basket until reaching an appropriate position on top of the basket for dropping;
+    MoveGripperToWithObj: The robot then moves its gripper towards container B until reaching an appropriate position over the top of the container.
   7. 7.
-    DropInto: the robot releases its gripper and drops the cube into the basket.
-In addition, we can further append an auxiliary Idle subtask, which encourage the robot to adjust its arm back to neutral position after dropping the object into the basket, facilitating subsequent deployment.
-This is an optional subtask that is not required for task completion.
-We include it in our task implementation to make the final policy behavior safe and more predictable after completing the full task.
-When evaluating the real world performance (e.g.
-Section VIII-D), we do not include this auxiliary subtask into the full task metric.
-We also exclude it whenever we discuss task-relevant decompositions since it is not a task related subtask.
-Formally, we instantiate the task compositional operator as
-$\mathcal{T}=\mathcal{C}(\{\tau^{k}\}_{k=1}^{K})\triangleq\tau^{1}\rightarrow\tau^{2}\rightarrow\cdots\rightarrow\tau^{K}$ (6)
-where $\tau^{i}\!\rightarrow\!\tau^{j}$ denotes that the two subtasks $\tau^{j}$ and $\tau^{j}$ are two adjacent subtasks and $\tau^{j}$ follows subtask $\tau^{i}$ in a sequential manner.
-In our case, $K\!=\!7$ and $\{\tau^{1},\tau^{2},\cdots,\tau^{7}\}$ corresponds to {Search, MoveTo, Grasp, SearchWObj, MoveToWObj, MoveGripperToWObj, DropInto}.
-Because of the decomposition, we can implement reward functions for each subtask, making it modular and easier for implementation.
-From some tasks, we simply use the sparse subtask success reward (e.g.
-Search).
-We incorporated delta-distance based shaping reward for tasks that requires to move the robot (or gripper) close to a target region (e.g.
-MoveTo, MoveToWObj, MoveGripperToWObj, Grasp, DropInto), in addition to the sparse subtask reward.
-For SearchWObj and MoveToWObj, we also incorporated an arm retract reward, which is another instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (c.f.
-Figure A.13).
-Intuitively, the arm retract reward will encourage the robot’s arm to raise up after grasping the object and move close to its neutral position.
+    DropInto: Finally, the robot releases its gripper and drops object A into container B.
+An illustration of the full task decomposition can be observed in Fig.
+5.
+Note that in this figure, we append an auxiliary Idle subtask, which encourages the robot to adjust its arm back to its starting position after finishing DropInto.
+This serves as an optional subtask and is used purely to facilitate subsequent deployments.
+Herein after, we do not include Idle in evaluations or discussions as it is unrelated to the task.
+Task decomposition offers many benefits.
+For one, we can implement reward functions for each subtask, aiding modularity and ease of implementation.
+For all subtasks, we incorporate a sparse subtask success reward.
+Additionally, for those requiring moving the robot (or gripper) to a target region (e.g., MoveTo, Grasp, etc.), we incorporate an additional dense delta distance-based shaping reward.
+For SearchWithObj and MoveToWithObj, we also incorporate an arm retract reward (another delta distance-based shaping reward), where the arm is encouraged to stay close to a predefined target pose, shown in Fig.
+6.
+This serves to promote good vantage and discourage suboptimal behaviors after grasping such as looking at the ground when searching.
 Its impact will be investigated in Section VIII-B (No Arm Retract baseline).
-An additional benefit of task decomposition is that we have the option to incorporate behavior priors in a flexible way into the policy.
-More details on the subtasks, rewards and behavior priors are provided in Appendix A-A.
+Finally, task decomposition offers the extra advantage of allowing behavior priors to be flexibly integrated into the policy.
+Such behavior priors can significantly increase sim-to-real task success as we will show later in Section VIII-D.
+More details on the subtasks, rewards, and behavior priors are provided in Appendix A-A.
+[…]
+Category Technique Description Reference Importance
+Rating (1-3)
+[…]
+a deterministic behavior that fails in the real world
+[…]
+[…]
+mismatch in the absence of system identification
+[…]
+[…]
+[…]
+for better generalization in real
+[…]
+[…]
+[…]
+[…]
+[…]
 VII Sim-to-Real Gap Reduction Techniques
-sim-to-real is notoriously difficult.
-This work handles two kinds of sim-to-real gaps: the physical dynamics and the visual gaps.
-For brevity, we discuss a subset of the techniques which are crucial in addressing sim-to-real and for successful real-world deployment.
-More techniques can be found in Appendix D.
-VII-A Dynamics Gap Reduction
-The dynamics gap is essentially caused by misaligned transition functions $P(\mathbf{s}^{t}|\mathbf{s}^{t-1},\mathbf{a}^{t-1})$ between the simulated and real worlds.
-This difference in dynamics will result in the sim-trained policy experiencing covariate shift when deployed in the real world, which if large enough, can destroy policy performance.
-In fact, even minor covariate shifts can result in task failure if the task requires fine motor control, e.g., grasping a small object.
-Therefore, to minimize the dynamics gap, we must minimize the impact of the covariate shift when transferring to real.
-In particular, we use the following key techniques for addressing the issue: 1) arm joint tracking error minimization, 2) domain randomization, and 3) object perturbations.
-Minimizing Arm Tracking Error via PID Control.
-We used PID controller for the robot arm.
-As opposed to PD controllers which will always have gravity-induced tracking errors, the integral term of PID controllers allow for more robust pose-agnostic tracking given that the joint delta targets are sufficiently small for particular operating frequency.
-This allows us to greatly reduce the dynamics sim-to-real gap while avoiding system identification.
-More details can be found in Appendix D-A.
-Arm Control Perturbation.
-Random structured noise is added to the arm control signal to make the learned policy to be robust w.r.t. different sources of control noises arising when transferring to real.
-Arm Mount Perturbation.
-We randomly perturb the arm mount position and yaw to make the learned policy robust to the sim-to-realgaps that could arise due to the actual mounting accuracy, torso height variations etc..
-Object Perturbations.
-The final strategy we employ are object perturbations.
-Due to the relatively deterministic nature of simulations (in contrast to the high-entropy real world), we observed that policies often converge to deterministic strategies.
-For example, for picking up an object, a policy may learn to always position itself so that the object has the same relative position to its base.
-This deterministic “memorizing” behavior can result in task failure if the conditions for success are even slightly off.
-Therefore, we enforce learning robust, reactive policies by perturbing objects of interest during various stages of task execution.
-Empirically, this helps with both the exploration (as the probability of stuck in a local state is reduced) and the robust ness of the policy (as the state coverage is enlarged).
-Stationary Manipulations.
-We make the quadruped stand as still as possible when manipulating objects, i.e.
-during grasp or drop into.
-This avoids the large sim-to-real gap of a walking and shaking dog when accurate manipulation is needed.
-To achieve this, when training low level policies to follow locomotion speed commands, we also make the quadruped enter a stationary mode with a balanced stance when the speed command is close to zero.
-Without stationary manipulation, not only will success rates suffer, but also it becomes unsafe to deploy, because the gripper fingers can hit the ground, damaging fingers and arm motors.
-VII-B Visual Gap Reduction
+In this section, we discuss crucial techniques for addressing sim-to-real gaps and achieving successful real-world deployment.
+The dynamics sim-to-real gap is the mismatch between the transition function of the simulation versus that of the real world.
+It can be due to the mismatch in the physical properties of objects and motors, or by things that are not properly simulated, such as friction and backlash.
+This difference in dynamics can cause policies trained purely in simulation to fail when deployed in the real world, especially if the task requires accurate motor control, such as grasping a small object.
 The second source of sim-to-real gap for image-conditioned policies is the vision gap.
-This gap is caused by an RGB distribution mismatch between simulated and real world pixels.
+This gap is caused by an RGB distribution mismatch between simulated and real-world pixels.
 Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios.
-Accordingly, we apply various vision sim-to-real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
-Image Domain Randomization.
-Inspired by SECANT [11], we apply pixel-level perturbations to every RGB image to increase the robustness of the perception model and the chance of its successful sim-to-real transfer.
-Specifically, we will randomly select a transformation from the list (Gaussian Noise, SaltPepper Noise, Speckle Noise, Random Brightness, Gaussian Blur) and apply it to each RGB image.
-We choose not to introduce higher-level image randomization/generation [54] because it might impact image semantics.
-Random Textures.
-We spawn the robot on an empty ground and randomize the ground texture.
-The texture can be arbitrary and generally have no specific semantics, as long as they provide enough visual distraction to the student.
-This distraction will force the student to learn to focus on the most important visual features that are related to the task at hand: objects with reasonable sizes and target colors.
-When deployed in real, we hope that this visual focus will get transferred and help the robot locate objects.
-Random Spatial Augmentation.
-To make the visual representation module robust, we incorporate spatial augmentations in the visual representation learning [51, 52].
-Random Background Objects.
-To further increase the visual complexity of the surrounding environment, we randomly sample and spawn a set of background objects in the room around the robot’s initial position.
-There are two categories of background objects: 1) primitive shapes (cuboid, cylinder, and sphere), and 2) Google scanned objects [10].
-We make sure that these objects only serve the purpose of distraction and the student will not confuse them with task objects.
-It could collide with an background object during training with some penalty reward.
-When spawning an object, we also randomize its size, orientation, and texture (only for primitive shapes).
-Refer to caption Refer to caption
-Figure 6: An example scene populated with random background objects. Left: close-up view; Right: birds-eye view.
+For brevity, we list techniques we deemed most relevant in Table I. Comprehensive details can be found in Appendix E.
 VIII Physical System, Experiments and Results
-In this section, we briefly go over the robot system and the task used to evaluate SLIM.
 VIII-A Physical Robotic System
-For our robotic system, we use a Unitree Go1 with a top-mounted WidowX250S manipulator.
-An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount, as shown in Fig.
-7, serving as the robot’s sole RGB vision feed.
-The depth stream is not used in this work.
+For our robotic system, we use a Unitree Go1 with a top-mounted WidowX-250S manipulator.
+An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount.
+The RGB stream of D435 is used as the robot’s sole vision feed (Fig.
+6).
 Additionally, custom 3D-printed, elongated parallel fingers are fitted on the WidowX gripper to extend its reach.
 The inner surface of the fingers are padded with a thin layer of foam, which act to increase grasping friction.
+Finally, the base of the WidowX is raised a few inches to allow for easy access to the onboard Raspberry PI’s ports.
 Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around \$12K USD.
@@
 Figure 6: Qualitative Overview of the Robot System.
+TABLE II: Dimension values for the observations and actions of our robot. For the meanings of these notations, we refer the reader to Section III.
+$L$ $H$ $W$ $N$ $D$ $K$ $Q_{m}$ $M$ $z$ $I$ $J$
+$100$ $90$ $160$ $5$ $66$ $7$ $19$ $6$ $0.05$ $6$ $12$
 […]
-VIII-B Baselines
-We compare SLIM with a number of baselines below:
-  • No Arm Retract: removing the arm retract (raise) reward in RL training, which is used to encourage the robot’s arm to raise up after grasping the object (introduced in Section VI);
-  • No Perturb: removing arm control randomization, arm mount perturbation and object perturbation (introduced in Section VII-A)
-  • No Visual Aug: no background objects and no random spatial augmentation based visual augmentation (introduced in Section VII-B)
-  • Distillation Only: agent training without using the RL loss (introduced in Section V), keeping only the distillation and representation loss.
-  • Human Teleop: an expert human trained for half an hour in teleoperating the robot with locomotion and arm end-effector commands observing only the view of the wrist camera.
-While not a baseline per se, human teleoperation provides a valuable reference point.
-See Appendix B-D for more details.
-We also tried another baseline No Distillation, which trains the visuomotor policy directly via RL without distillation from the teacher policy (i.e.
-RL for visuomotor policy learning from scratch).
+Given our hardware configuration, the dimension values from Section III are summarized in Table II.
+VIII-B Ablations and Baselines
+We compare SLIM against the following baselines, where the first four are ablations and the last is a human teleoperator as a reference:
+  • No Arm Retract: Remove the arm retract (raise) reward in RL training (Section VI).
+  • No Perturbation: Remove arm control randomization, arm mount perturbation, and object perturbation (Section E-A).
+  • No Visual Aug: Remove random background objects and spatial visual augmentations (Section E-B).
+  • Distillation Only: Train student without the RL loss, keeping only the distillation and representation loss (Section V).
+  • Human Teleop: An expert human teleoperator that provides locomotion commands and delta EE pose to the robot via a joystick.
+Delta EE is handled via IK.
+The teleoperator shares the same observation space as the policy, i.e., observing solely through the wrist camera stream, and is allowed to practice on the system for an hour.
+While not a baseline intrinsically, human teleoperation provides a valuable reference point.
+See Appendix I for more details.
+We also tried another baseline No Distillation, which trains the student policy directly via RL without distillation from the teacher policy (i.e., RL for visuomotor policy learning from scratch).
 This baseline cannot learn to solve the task at all in simulation, because of the compounded difficulties of representation learning, behavior learning and long-horizon exploration.
-It is excluded from the tables.
-VIII-C Evaluation Protocol
-We use the Standard object spatial layout as shown in Figure 8 in a real-world scene Lobby.
-By doing so, it is easier to maintain repeatability across all methods and seeds.
-Besides the Standard in lobby, we also evaluate the robot in several other scenes with other layout.
+As such, it is excluded from the results.
+VIII-C Metrics and Evaluation Protocol
+To accurately evaluate policy performance, we run the entire training pipeline across three different random seeds for all methods, training from scratch the low-level, teacher, and student policies.
+Note that each method for a particular seed shares the same low-level policy.
+We report the following two key metrics when deploying the final student policy:
+  1. 1.
+    Cumulative Subtask Success Rate: The success rate up to each subtask from the beginning of the episode.
+  2. 2.
+    Episode Time: The time spent completing the full task from the beginning. We apply a time limit of $t_{\rm max}\!=\!90$ s. For failed episodes, we use $t_{\rm max}$ as the episode time.
+For the deployment environment, we use the Standard object spatial layout in an indoor Lobby scene, as shown in Fig.
+7, to maintain repeatability across all methods and seeds.
 For task objects, we use cubes and baskets of different colors.
-More details on these physical objects are detailed in Appendix F-A.
-For each method of each random seed, we roll out the robot’s policy for 20 episodes with varied initial scene setup (i.e.
-different positioning and colors of objects) according to the evaluation protocol.
-For the detailed protocol, please refer to Appendix F-B.
-TABLE I: Real-world task success rates and episode time (mean±$\pm$±stddev over 3 random seeds), with 20 episodes per seed.
-Method Fully Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
-Autonomous Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
+For each method of each random seed, we roll out the robot’s policy for 20 episodes with varying object types and positions according to a consistent evaluation protocol, with 360 real-world episodes of experiments.
+More details on the physical objects and protocol can be found in Appendices G-A and G-B, respectively.
+In addition to the Standard spatial setup in Lobby, we also evaluate the robot in several other scene and layout combinations in subsequent experiments, with another 40 real-world episodes, leading to a total of 400 real-world episodes of experiments.
+TABLE III: Real-world task success rates and episode time (mean±$\pm$±stddev over 3 random seeds), with 20 episodes per seed.
+Method Autonomous Cumulative Success Rate [%] $\uparrow$ Episode Time [s] $\downarrow$
+Search+MoveTo Grasp Search+MoveTo(WObj) DropInto (Full Task)
 […]
@@
 […]
-Figure 8: Scene Spatial Layouts.
+Figure 7: Scene Spatial Layouts in Lobby Environment.
 Standard: we put the robot in the center of a square space and objects including objects to be grasped and containers to drop the graspable objects into on the four corners.
 The instructions and object placements vary across episodes.
-Cluttered: objects are placed close to each other in front and behind the robot, leading to OOD spatial layouts.
+Cluttered: objects are placed close to each other in front and behind the robot, leading to an OOD spatial layouts.
 The instructions and object placements vary across episodes.
-We conduct experiments on the baselines and the proposed method, and measure their success rates in the real world. The results are summarized in Table I.
-We run the whole training pipeline with three different random seeds to train low-level, the teacher and the student, and measure the performance of the final student policy in real, in terms of accuracy and efficiency:
-  1. 1.
-    Cumulative Subtask Success Rate: the success rate up to the specified subtask from the beginning of the episode.
-  2. 2.
-    Episode Time: the time spent completing the full task from the beginning. We apply a time limit of $t_{\rm max}\!=\!90$s. For failed episodes, we use $t_{\rm max}$ as episode time.
-Finally, we report the mean and standard deviation across the seeds as the final aggregated metric value in Table I.
-Table I shows how important each of the interventions used contribute to the final successes.
-No Arm Retract baseline performs well for the initial two subtasks of Search and MoveTo, but not for the later subtasks.
-Even though half of the times the robot can move to the basket carrying the cube (a cumulative success rate of $28.3\%$), it only successfully drops the object into the basket less than one fifth of the cases (a cumulative success rate of $5.0\%$).
-The main reason is that without arm retract reward, the arm could be at an over-extended pose, getting closer to the physical torque limits of the shoulder motors responsible for raising the arm.
-This leads to more arm shake, cube slipping out of the fingers, motor failures, and makes it harder for the arm to raise up to search for the basket and to hover above the basket to complete DropInto.
-No Perturb baseline has the lowest grasping success rate.
-Without perturbations (object, arm mount and control), the policy is not robust enough to deal with the variations in the location of the object relative to the robot.
+All success rate and task completion results are summarized in Table III. As shown, we can see the negative impact of ablating away each technique on the final task success.
+In particular, No Arm Retract performs fairly well up until Grasp, but then soon its performance drops off substantially.
+In fact, No Arm Retract possesses the lowest cumulative success rates past grasping across the board, with a sharp dropoff of approximately 30% when attempting to search the container.
+There are several reasons for this dropoff.
+For one, without the arm retract reward acting as a soft constraint on the arm’s position, the arm could be at an over-extended pose, getting closer to the physical torque limits of the shoulder motors responsible for raising the arm.
+This leads to more arm shake, which in turn can result in the grasped cube slipping out of the fingers or even motor failures.
+Another key reason for search-based failure arises from possible low vantage search strategies, e.g., the camera is pointed toward the ground.
+Intuitively, No Perturbation has the lowest grasping success rate.
+This makes sense as without perturbations of any kind (object, arm mount, and control), the policy is not robust enough to deal with the variations in the location of the object relative to the robot.
 This is likely because during simulation training, the policy can get very efficient at approaching the target object and stopping just at the success boundary of the MoveTo subtask, so that the Grasp subtask usually starts with very similar locations of the target object relative to the robot, while in the real world such clear cut positioning is rarely seen.
-No Visual Aug’s performance on the initial Search subtask is lower than the that of the above baselines.
-The main reason is that without visual augmentations, it can be distracted by background objects irrelevant to the task, i.e.
-false positive detection of the target.
-This shows that visual augmentations are important for improving the robustness of the robot’s vision module.
-Distillation Only baseline simply imitates from the highly successful teacher policy, and can indeed achieve a relatively high success rate.
+Performance of No Visual Aug on each of the Search subtasks suffer as the model sometimes gets distracted, having false positive detections of the target in the background.
+There’s a particularly large drop of roughly 25% when searching for the basket.
+This indicates that visual augmentations are important for improving the robustness of the robot’s vision module.
+Finally, Distillation Only simply distills the highly successful teacher policy, and can indeed achieve a relatively high success rate.
 However, its performance is not consistent across random seeds.
-Our full method (SLIM) uses distillation-guided RL, and results in consistently high success rates across seeds.
+In comparison, SLIM uses distillation-guided RL, resulting in consistently high success rates across all seeds indicated by the low standard deviations.
 It shows that distillation loss alone is not enough for robust policy learning.
-Additionally, task completion is also slower than the full method which additionally uses RL loss to optimize the policy.
-Our full method SLIM achieves the highest subtask and full task success rate among all the methods, and is also the most efficient, taking only 43.8 seconds to complete the full task on average.
+Additionally, task completion is also slower as the RL loss allows the policy to further optimize.
+Overall, SLIM achieves the highest subtask and full task success rates among all the methods including Human Teleop, and is also the most efficient, taking on average 43.8 seconds to complete the full task, which is about 1.5$\times$ the speed of a teleoperation expert.
+We provide some visualization of the policy behaviors and prediction outputs in Appendix H.
 […]
-Figure 9: Real-world Performance of SLIM across Scene Variations. Left: images showing the differences for evaluation. Right: the full task success rate across difference scenes.
+Figure 8: The same SLIM Policy across Scene Variations.
+Left: images showing the differences for evaluation.
+Right: the average full task success rate across 10 real-world deployment episodes for each of the scenes shown on the left.
 VIII-E Generalization to Different Real-world Scenes
-In addition to the standard scene used in Table I, we further run real-world evaluations of SLIM under more scene variations (c.f.
-Figure 9 left): Outdoor, Carpet-Room, Kitchen and Lobby.
-For the first three scenes, we use the standard spatial layout (c.f.
-Figure 8).
-For Lobby, we use the Cluttered spatial layout, which is different from the training (Standard) layout.
-Results as summarized in Figure 9 are close to the Standard layout or a bit lower due to being out of distribution.
-This shows that the model trained fully in simulation can zero-shot adapt to a wide range of real world settings, with vastly different lighting conditions, background, and floor materials and textures.
+In addition to the standard scene used in Table III, we further run real-world evaluations of SLIM under more scene variations (left side of Fig.
+8): Outdoor, Carpet Room, Kitchen, and Lobby-Cluttered.
+For the first three scenes, we use the same Standard spatial layout as before.
+For the Lobby-Cluttered scene, we use the Cluttered spatial layout (right side of Fig.
+7), which is different from the Standard layout used during training.
+We conduct ten trials for each environment using the same SLIM policy.
+All results for these four additional scenes are summarized in the right side fo Fig.
+8.
+For the three novel scenes, we observe that success rates are all fairly close to the results from Table III ($\sim$78%).
+For the Lobby-Cluttered scene, we can see that the success rate is quite close to the Standard layout and is just a bit lower due to being out of distribution.
+This shows that the model trained fully in simulation can zero-shot adapt to a wide range of real world settings, with vastly different lighting conditions, backgrounds, and floor materials and textures.
 Furthermore, our qualitative evaluation showed that the robot also generalizes well to scenarios where:
   1. 1.
-    task objects are randomly scattered on the ground (e.g., last two rows in Figure 1),
+    task objects are randomly scattered on the ground (last two rows in Fig. 1),
   2. 2.
-    distractor objects are present (e.g., last row in Figure 1),
+    distractor objects are present (last row in Fig. 1),
   3. 3.
-    human interrupts the task progress, or
+    a human interrupts the task progress (in demo videos), or
   4. 4.
-    objects to grasp have slightly different shapes.
-We refer the reader to our demo videos for more emergent generalization behaviors. In general, the robot is quite robust to the real-world scene complexity.
-VIII-F Comparison with Simulation Performance
-Different methods or even the same method with different seeds can exhibit very different behaviors, mostly due to the variability in the RL training of the teacher policies.
+    objects to grasp have slightly different shapes (in demo videos).
+We refer the reader to our demo videos in the supplementary file for more emergent generalization behaviors.
+VIII-F Sim vs. Real Performance
+Different methods or even the same method with different seeds can exhibit very different behaviors, mostly due to the variance in the RL training of the teacher policies.
 Additionally, because of the sim-to-real gap, real deployment can differ from simulation as well.
-Figure 10 shows simulation performance with the real performance for our methods, so we know more clearly how each method helps the training process.
-The Sim success rate for each method denotes the average full task success rate over 200 episodes run in simulation, averaged across the three random seeds.
+To better understand how each technique impacts sim-to-real gap reduction, we showcase the simulation and real performance difference for all methods in Fig.
+9.
+The sim success rate for each method denotes the average full task success rate over 100 episodes run in simulation, averaged across the three random seeds.
 The simulation environment follows the same randomization setting as training, with the policy rollout being fully greedy.
-The Real success rate denotes the full task success rate of the policy in the real-world, following the Standard spatial layout as shown in Figure 8.
-Without arm retract (No Arm Retract), searching for the basket becomes more difficult in simulation as well as in real, because after grasp, the camera still points down to the ground, but searching for the basket requires many steps of the arm raising up, creating an exploration bottleneck for RL training.
-Real world performance suffers more, because moving with an out-stretched arm causes large sim-to-real gaps: The tip of the arm tends to shake more due to motor backlash, hard to model in simulation, and the shoulder motors operate close to their physical torque limits, leading to more motor overload.
-For No Perturb, while one might expect perturbations to complicate RL learning, their inclusion results in a much higher performance even in simulation.
+The real success rate denotes the full task success rates from Table III.
+Without arm retract (No Arm Retract), searching for the basket becomes more difficult in simulation as well as in real, because after grasping, the camera still points down to the ground, but searching for the basket requires many steps of the arm raising up (without explicit rewards encouraging it), creating an exploration bottleneck for RL.
+Real-world performance suffers more, because moving with an out-stretched arm causes large sim-to-real gaps: the tip of the arm tends to shake more due to motor backlash, which are hard to be modeled precisely in simulation, and the shoulder motors operate close to their physical torque limits, leading to more motor overloading.
+For No Perturbation, while one might expect perturbations to complicate RL learning, their inclusion results in much higher performance even in simulation.
 Without perturbations, the policy seems to converge to suboptimal solutions maybe due to a lack of exposure to a sufficiently large range of setups.
-Without visual augmentations (No Visual Aug), simulation task is easier to solve, thus simulation success is artificially high, but real world performance suffers when compared to the full method.
-RL training loss in the full SLIM method increases success in simulation (over the imitation only Distillation Only) and achieves the highest success rate in real.
-It’s interesting to note that the RL loss seems to help the No Visual Aug baseline to achieve slightly better real world success than Distillation Only.
+Without visual augmentations (No Visual Aug), the simulation task is easier to be solved, and thus simulation success is artificially high, but real-world performance suffers when compared to the full method.
+RL training loss in the full SLIM method increases success in simulation (over Distillation Only) and achieves the highest success rate in real.
 […]
-Figure 10: Full Task Success Rates of Different Methods (averaged across three seeds) in Simulation and Real.
-[…]
-Figure 11: Policy Trajectory Visualization.
-For visualization purpose, we removed background objects.
-We also visualize the ee trajectories.
-Trajectories corresponding to different subtasks are rendered with different colors.
-The ego-centric camera view is shown on the top-right of each image.
-The red colored arrows denote the direction of movement.
-VIII-G Policy Behavior Visualization
-The behavior of the policy is shown in Figure 11.
-To make the visualization clear we did not add background objects in the simulation for this episode.
-The robot start with a randomized position and yaw within a region (Start phase).
-A red cube and a green cube as well as a red basket and and blue basket are used.
-Their positions are randomized around the Standard layout and their yaw angles are also randomized.
-Given an instruction (“Drop the yellow cube into the red basket.”), the robot will first search for the target graspable object (the yellow cube).
-Note that during the Search phase, the robot actively moves both its body and the arm in searching of the target object (c.f.
-Figure 11-Search).
-Once the target object is in view, the robot will move towards it (Figure 11-MoveTo).
-Once the robot moves close to the target object, it will start to grasp (Figure 11-Grasp).
-During this stage, the robot body remains stationary and the arm will move towards the target cube and grasp it.
-After picking up the target cube, the robot will search for the target basket while keeping the cube in gripper.
-As shown in Figure 11-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory).
-At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular EE trajectory in Figure 11-SearchWithObj.
-Once the target basket is in view, the robot will move towards it (Figure 11-MoveToWithObj) and drop the cube in gripper into that basket (Figure 11-DropInto).
-VIII-H Failure Modes
-We also recorded common failure cases for SLIM.
-In one case, when grasping the quadruped stopped too late and kicked the cube far away.
-In the other case, the low-level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach.
-DropInto sometimes fails because of either early dropping right outside of the basket, or the quadruped kicking the basket away right before dropping when backing towards the basket.
-In rare cases where the gripper holds only the upper half of a cube, the robot tends to get stuck at hovering over the basket while holding the cube, because the cube is completely blocked by the gripper fingers in the wrist camera view.
+Figure 9: Full Task Sim and Real Success Rate Comparison. All success rates are the averaged success rate across three random seeds.
+VIII-G Failure Modes
+We also recorded failures for SLIM and report the most common cases here.
+In one case, the quadruped stopped a bit late during grasping and kicked the cube away.
+In another case, the low-level policy due to sim-to-real gap shook and moved backward, causing the cube to be just out of reach.
+DropInto sometimes failed because of either early dropping right outside of the basket, or the quadruped kicking the basket away, right before dropping, when backing towards the basket.
+In rare cases where the gripper held only the top part of the cube, the gripper fingers completely blocked the cube from the camera view, and the robot got stuck hovering over the basket while holding the cube.
 See our failure demo videos for more details.
@@
 Currently the low-level policy is only trained on flat terrain in simulation.
-Although it shows the potential to work across different types of grounds (e.g.
-concrete, tile floor, carpet, mulch etc.), it does not push the low-level policy learning to its full potential.
-We can potentially leverage the recent advances in locomotion policy training [62, 7] to further expand the ability and applicability of SLIM across more diverse types terrains, e.g.
-stairs, ramp etc.
-Another low-level-related limitation is that the current form of low-level policy is not a wholebody policy.
-Although the current form of low-level policy is sufficient for the task considered in this work, by generalizing it to a wholebody policy [15], we can potentially further enlarge its potential workspaces.
+Although it shows the potential to work across different types of grounds (e.g., concrete, tile floor, carpet, mulch, etc.), it does not push the low-level policy learning to its full potential.
+We can potentially leverage the recent advances in locomotion policy training [67, 8] to further expand the ability and applicability of SLIM across more diverse types terrains, e.g., stairs, ramp, etc.
+Another low-level related limitation is that the current form of low-level policy is not a wholebody policy.
+Although the current form of low-level policy is sufficient for the task considered in this work, by generalizing it to a wholebody policy [16], we can further enlarge its potential workspace.
 As a pure sim-to-real approach that is trained from scratch, we did not leverage any pre-trained foundation models.
-As a result, we did not address large amount of visual concepts and language variations and instead focused on the right amount of variations that are both practical and representative enough.
+As a result, we did not address a large amount of visual concepts and language variations and instead focused on the right amount of variations that are both practical and representative enough.
 There are also a few limitations on the robot system side.
 Currently for the visual perception, we only used the RGB image from the wrist-mounted camera, which is the minimum setting that is sufficient for the current work.
-To further improve the robot’s perception and therefore the overall ability in terms of navigation and obstacle avoidance, it might be necessary to incorporate additional sensors for perception such as depth, stereo and lidar etc.
-In additiona, this can also provide valuable information for the robot to improve its longer-term memories via SLAM-like mechanisms [57, 5].
+To further improve the robot’s perception and therefore the overall ability in terms of navigation and obstacle avoidance, it might be necessary to incorporate additional sensors for perception such as depth, stereo, lidar etc.
+In addition, this can also provide valuable information for the robot to improve its longer-term memories via SLAM-like mechanisms [63, 7].
 X Conclusion, Discussion and Future Work
-We present a complete legged mobile manipulation robot system for solving language-instructed long horizon task in this work.
+We present a complete legged mobile manipulation robot system for solving language-instructed long-horizon tasks in this work.
 The policy is trained via the sim-to-real paradigm, together with a progressive policy expansion-based teacher policy for solving the long-horizon task, followed with a distillation-guided RL approach for student visuomotor policy learning.
-We further identified and reduced a number of crucial sim-to-real gaps.
+We further identified and designed a suite of crucial techniques for reducing sim-to-real gap.
 Experimental results compared with a number of baseline methods verified the effectiveness of the proposed SLIM method.
-Moreover, real-world testing of SLIM across different scenes and spatial layouts shows that the proposed method can generalize well, further demonstrating the effectiveness of the proposed whole pipeline.
+Moreover, real-world testing of SLIM across different scenes and spatial layouts shows that the proposed method can generalize well, further demonstrating the effectiveness of the whole system.
 For future work, we would like to address the limitations discussed in Section IX, including further expanding the applicability of SLIM across more diverse scenarios by improving locomotion policy training [67, 8, 16].
@@
 Since we decompose the full long task into a sequence of subtasks, in practice, we need to decide the task boundaries in order to properly divide the full task.
-In the sequential decompositional case as described in Eqn.(6), the task boundaries $\tau^{i}|\tau^{i+1}$ between two adjacent subtasks $\tau^{i}\rightarrow\tau^{i+1}$ can be defined by the success condition $\mathcal{S}(\tau^{i})$ of the subtask $\tau^{i}$ (i.e.
-$\tau^{i}|\tau^{i+1}\equiv\mathcal{S}(\tau^{i})$), since the success of a subtask $\tau^{i}$ naturally leads to the subsequent subtask $\tau^{i+1}$.
+In the sequential decompositional case as described in Section VI, the task boundaries $\tau^{i}|\tau^{i+1}$ between two adjacent subtasks ($\tau^{i}\rightarrow\tau^{i+1}$) can be defined by the success condition $\mathcal{S}(\tau^{i})$ of the subtask $\tau^{i}$ (i.e., $\tau^{i}|\tau^{i+1}\equiv\mathcal{S}(\tau^{i})$), since the success of a subtask $\tau^{i}$ naturally leads to the subsequent subtask $\tau^{i+1}$.
 We provide the success conditions of each subtask (therefore the subtask boundaries) in the sequel:
@@
   4. 4.
-    $\mathcal{S}$(SearchWObj): succeeds when the target object is in gripper and the target basket enters the view of the wrist-mounted RGB camera;
+    $\mathcal{S}$(SearchWithObj): succeeds when the target object is in gripper and the target basket enters the view of the wrist-mounted RGB camera;
   5. 5.
-    $\mathcal{S}$(MoveToWObj): succeeds when the target object is in gripper and the target basket enters a workspace of the robot within arm’s reach;
+    $\mathcal{S}$(MoveToWithObj): succeeds when the target object is in gripper and the target basket enters a workspace of the robot within arm’s reach;
   6. 6.
-    $\mathcal{S}$(MoveGripperToWObj): succeeds when the target object is in gripper and the grippers moved to the top of the target basket;
+    $\mathcal{S}$(MoveGripperToWithObj): succeeds when the target object is in gripper and the grippers moved to the top of the target basket;
   7. 7.
@@
 Rewards.
-We summarize the description of the subtask reward in Table II.
+We summarize the description of the subtask reward in Table IV.
 The task reward represents the sparse task success ($+1$) or failure ($-1$) reward.
 For timeout, we used a reward of 00.
-Apart from the sparse task reward, we also incorporated some shaping rewards as shown in Table II.
+Apart from the sparse task reward, we also incorporated some shaping rewards, as shown in Table IV.
 Most of the shaping rewards are delta-distance based shaping reward, encouraging two position of interest to getting closer to each other.
-Arm retract reward is also an instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (c.f.
-Figure A.13).
-Intuitively, the arm retract reward will encourage the robot’s arm to raise up after grasping the object and move close to its neutral position.
+Arm retract reward is also an instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (please refer to Fig.
+6 for a reference of the neutral arm pose).
+Intuitively, the arm retract reward will encourage the robot’s arm to rise up after grasping the object and move close to its neutral position.
 Alignment reward is also an instance is delta-distance based shaping reward, where the distance is defined based on the cosine between the left and right finger-tip-to-object vectors.
 Keep-grasping reward is a contact based reward issuing a penalty of $-0.1$ is the contact between gripper fingers and the object is lost.
-TABLE II: Rewards and Behavior Prior for Subtasks
+TABLE IV: Rewards and Behavior Prior for Subtasks
 Subtask Reward Behavior Prior
@@
 Grasp task + distance + alignment stationary
-SearchWObj task + arm-retract rotational
-MoveToWObj task + distance + arm-retract + keep-grasping
-MoveGripperToWObj task + distance + keep-grasping stationary
+SearchWithObj task + arm-retract rotational
+MoveToWithObj task + distance + arm-retract + keep-grasping
+MoveGripperToWithObj task + distance + keep-grasping stationary
 DropInto task + distance stationary
@@
   • rotational: the forward velocity command ($\mathbf{c}[0]=0$) to the quadruped is kept as zero and therefore enable the quadruped to move in a rotational fashion.
-This is used as the behavior prior for search related subtasks (Search, SearchWObj).
+This is used as the behavior prior for search related subtasks (Search, SearchWithObj).
   • stationary: both the forward and angular velocity of the command to the quadruped are hardcoded as zero ($\mathbf{c}=0$).
 This encourages the robot to be stationary.
-It is used in manipulation related tasks (Grasp, MoveGripperToWObj, DropInto).
-In practice, both priors are implemented in a similar way, by initializing the teacher’s actor network for those subtasks with these priors to be of zero mean and small std for the corresponding action dimensions (e.g.
-the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior).
+It is used in manipulation related tasks (Grasp, MoveGripperToWithObj, DropInto).
+In practice, both priors are implemented in a similar way, by initializing the teacher’s actor network for those subtasks with these priors to be of zero mean and small std for the corresponding action dimensions (e.g., the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior).
 The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution and being consistent with subtasks without these priors.
-Hierarchical RL.
+A-B Hierarchical RL and High-Level RL Training
 As typical in standard hierarchical RL stuctures, the high-level policy and low-level policies are running at different frequencies.
@@
 Because of this, the RL is essentially running at the high-level.
-Therefore, the per-(low-level-)step reward during each re-planning period is accumulated as the step reward for one high-level step, i.e.
+Therefore, the per-(low-level-)step reward during each re-planning period is accumulated as the step reward for one high-level step, i.e.,
 $r_{\rm high\_level\_step}=\sum_{i=1}^{5}r_{\rm low\_level\_step}^{i},$ (3)
 and this accumulated reward $r_{\rm high\_level\_step}$ is used as the per-step reward signal for high-level RL training.
-A-B Low-level Task and Reward
+A-C Low-level Task and Reward
 Low-level Tasks.
-Details of low-level training tasks/commands and corresponding rewards are summarized in Table III.
-TABLE III: Low-level Training Tasks and Rewards
+Details of low-level training tasks, commands, and corresponding rewards are summarized in Table V.
+Quadraped Gait Control.
+Motivated by [39], we also introduce gait alignment rewards that incentivize the robot to follow predefined gait patterns.
+The gait patterns are defined by a phase offset vector of length three, where each entry represents the phase difference of each of the three legs (excluding the first one) relative to the first leg (fore left).
+The range of each entry is $[-0.5,0.5]$, where 00 indicates a synchronous movement with the first leg and $-0.5$ or $0.5$ indicates a half cycle phase difference.
+The phase of a leg is evaluated independently for each leg, based on the leg’s schedule at current time.
+The schedule is a number in $[0,1]$.
+For gait alignment rewards design, schedule $1$ represents the “swing” phase where smaller movements in the X-Y plane will be penalized; schedule 00 represents the “contact” phase where lower contact force will be penalized; for any schedule between $(0,1)$, both aspects will be penalized with weights modulated by the schedule value.
+For training of the gait following ability, we expand the input command $\mathbf{c}$ with randomly sampled phase offset vector $\mathbf{g}_{\rm offset}$ and gait cycle vector $\mathbf{g}_{\rm cycle}$.
+$\mathbf{g}_{\rm offset}$ is sampled from five predefined gait patterns, as summarized in Table VI.
+$\mathbf{g}_{\rm cycle}$ is a two-dimensional vector, with each dimension representing the frequency and the stance ratio respectively.
+The frequency takes values in $[2.0,4.0]$, indicating the number of cycles that each leg goes through per second.
+The stance ratio takes values in $(0.4,0.6)$, indicating the portion of time within a cycle when each foot is in contact with the floor.
+After low-level training, we fix the gait pattern to be “trot”, i.e., $\mathbf{g}_{\rm offset}\!=\![-0.5,-0.5,0.0]$, and the gait cycle vector as $\mathbf{g}_{\rm cycle}\!=\![2.4,0.5]$.
+We have found that this gait is suitable for the mobile manipulation task tackled in this work in terms of movement efficiency and stability.
+For the stationary task of low-level training, we always set $\mathbf{g}_{\rm cycle}\!=\![0,1]$.
+TABLE V: Low-level Training Tasks and Rewards
 Task/Command Name Sample Range Reward
@@
 $-0.3\cdot\mathbbm{1}_{\text{killed}}$
-gait following (optional)
+TABLE VI: Predefined Gait Patterns
+Gait Names Phase Offsets
+pronk $(0.0,0.0,0.0)$
+walk $(-0.5,0.25,-0.25)$
+trot $(-0.5,-0.5,0.0)$
+bound $(0.0,0.5,0.5)$
+pace $(-0.5,0.0,-0.5)$
 Low-Level Domain Randomization.
-The domain randomization setting for low-level training is summarized in Table IV.
-TABLE IV: Low-level Domain Randomization Parameters
+The domain randomization setting for low-level training is summarized in Table VII.
+TABLE VII: Low-level Domain Randomization Parameters
 Env Params Sample Range
@@
 Appendix B Model Details
-B-A Teacher High Level
-The teacher network $\Pi\!=\!\{\pi^{k}\}_{k=1}^{K}$ is a set of sub-policies (referred to as policy set), as shown in Figure 3.
+B-A Teacher (High Level)
+The teacher network $\Pi\!=\!\{\pi^{k}\}_{k=1}^{K}$ is a set of sub-policies (referred to as policy set), as shown in Fig.
+3.
 The effective size of the policy set is $K\!=\!8$ (including the Idle subtask).
@@
 The re-planning period is $5$, meaning that the high-level policy predicts once every $5$ low-level steps.
-B-B Student High Level
+Figure A.10: The architecture of the student’s U-Net for predicting segmentation and depth maps.
+Figure A.11: The architecture of the student’s policy input encoder that generates a compact observation embedding for RL.
+B-B Student (High Level)
 Maps Prediction U-Net.
-An illustration of the U-Net [42] for predicting segmentation and depth maps is shown in FigureA.14. We will talk about the architecture details by input modality below.
+An illustration of the U-Net [45] for predicting segmentation and depth maps is shown in Fig.A.10. The architecture details will presented according to input modality below.
   1. 1.
@@
 Policy Inputs Encoder.
-An illustration of the policy input encoder network is shown in Figure A.15.
+An illustration of the policy input encoder network is shown in Fig.
+A.11.
 The network is responsible to digest predicted segmentation&depth maps and proprioceptive observations, and produce a compact representational embedding for downstream RL networks.
@@
 Stationary Bit.
-The teacher policy incorporates a stationary behavior prior into the locomotion command $\mathbf{c}$ for several subtasks (Table II).
+The teacher policy incorporates a stationary behavior prior into the locomotion command $\mathbf{c}$ for several subtasks (Table IV).
 With this prior, the corresponding locomotion command distribution always has a zero mean and a small std.
@@
 B-C Low Level
-For the low-level training, we used PPO [44] with a regularized online adaptation module Fu et al. [15].
+For low-level training, we use PPO [47] with a regularized online adaptation module [16].
 Policy Input and Locomotion Command.
-As shown in Figure 2, the low-level locomotion policy takes two types of inputs: 1) the standard proprioceptive state $\mathbf{s}_{\textrm{leg}}$ obtained from the environment, 2) the locomotion command $\mathbf{c}$ from an external command issuer (e.g..
-The second type of input is crucial to make the low-level policy useful in downstream tasks where locomotion is involved as a basic skill for exploration and learning.
-During training, the command is re-sampled periodically within each episode and the low-level policy is trained to follow the command $\mathbf{c}$ via the command-following rewards detailed in Table III.
-The arm pose is also adjusted periodically within each episode, to make the low-level policy’s command following robust against arm movement.
-For the standard proprioceptive input to the low-level policy, while it is also possible to include those from the arm.
-Empirically, we discover that it might trigger some shakiness behavior in certain cases and therefore we only used $\mathbf{s}_{\textrm{leg}}$.
-After training, the low-level policy can be used to follow commands issued by an external command issuer, e.g.
-either a higher level policy (as in SLIM and all the other baseline except Human Teleop), or human tele-operator (as in Human Teleop) in this work.
+As shown in Fig.
+2, the low-level locomotion policy takes two types of inputs: 1) the standard proprioceptive state $\mathbf{s}_{\textrm{leg}}$ obtained from the environment, 2) the locomotion command $\mathbf{c}$ from an external command issuer (e.g., The second type of input is crucial to make the low-level policy useful in downstream tasks where locomotion is involved as a basic skill for exploration and learning.
+During training, the command is re-sampled periodically within each episode and the low-level policy is trained to follow the command $\mathbf{c}$ via the command-following rewards detailed in Table V.
+The arm pose is also adjusted periodically within each episode, to make sure the low-level policy can follow locomotion commands while being robust against arm movements.
+For the standard proprioceptive input to the low-level policy, while it is also possible to include those from the arm, empirically, we discover that it might cause the quadruped to shake more often, and decide to only use $\mathbf{s}_{\textrm{leg}}$.
+We also discover that the joint velocities in the hardware are delayed more than joint position readings from the motor sensors, possibly due to temporal smoothing.
+Since we do not know the exact algorithm that computes the joint velocities, we simply do not use them as input, and rely on feeding the neural networks the previous few frames for the policy to estimate the joint velocities.
+After training, the low-level policy can be used to follow commands issued by an external command issuer.
+This can be either a higher level policy (as in SLIM or all the autonomous baselines), or human teleoperation (Human Teleop) in this work.
 Actor and Value Networks.
@@
 .
-Similar as in [15, Figure 2], our online adaptation module consists of two encoders, the privileged encoder and the adaptation encoder.
+Similar as in [16, Fig.2], our online adaptation module consists of two encoders, the privileged encoder and the adaptation encoder.
 The privileged encoder takes privileged information as input and predicts a latent representation of environment extrinsics, while the adaptation encoder predicts the same latent representation depending only on robot sensory observations.
@@
 The privileged encoder network has two hidden layers of $(128,64)$ and the adaptation encoder has hidden layers of $(256,128)$.
-B-D Human Teleoperation
-To show that we’ve made reasonable efforts to make Human Teleop work, we list its failure modes.
-1) The most common failure mode is task timeout.
-This is sometimes due to not fully grasping the cube and the cube slipping out of the fingers during movement, leaving no time to search and re-grasp.
-Failing to grasp fully can be due to either stopping a bit far from the cube, or incorrectly assessing the distance to the cube under monocular vision.
-2) The arm can enter a twisted pose due to poor IK solutions, despite having already tuned IK as we can.
-Human teleoperation is in general less accurate and slower than RL policy, due to Low level policy shaking, and human operator fatigue.
-Besides being slower, human teleoperation seems to be less energy efficient than our RL policies.
-When RL learned high-level policies can run smoothly on a particular low level policy, human teleoperation on the same low level policy seems to trigger over-current protection of Go1.
-This often happens when trying to raise an over-extended arm, causing Go1 to lose control of its leg motors.
-In evaluations, we ignore these Go1 hardware failures, and let the human operator retry for free.
-B-E Model Parameter Count
-We summarize in Table V the rough number of model parameters for different components of our system, including low level, teacher high level and student high level.
-TABLE V: Number of Trainable Parameters.
+B-D Model Parameter Count
+We summarize in Table VIII the rough number of model parameters for different components of our system, including low level, teacher high level and student high level.
+TABLE VIII: Number of Trainable Parameters.
 Component Approx. Param. Count
@@
 The Adam optimizer had an initial learning rate of $2\times 10^{-4}$ and a step decay to its $1/5$ after $70\%$ of the total training progress.
-Appendix D More Sim2Real Techniques
-D-A More Dynamics Gap Reduction Techniques
-Minimizing Tracking Error via PID Control.
+Appendix D Simulator Choice
+All simulations were done in the MuJoCo simulator [51].
+At the beginning of the project, we evaluated several popular simulators at the time, including MuJoCo [^4] https://mujoco.org/, Isaac Gym [^5] https://developer.nvidia.com/isaac-gym, and PyBullet [^6] https://pybullet.org/wordpress/.
+We settled on MuJoCo due to its reasonably fast simulation (on CPU) and rendering (on GPU) at the same time.
+Besides camera observation rendering for student training, due to partial observability, we also rely on segmentation rendering to check which objects are in the camera view and use it as privileged information for the teacher, during both teacher and student training.
+Thus RGB and segmentation rendering efficiencies are critical to our simulator choice.
+Appendix E Sim-to-Real Techniques
+E-A Dynamics Gap Reduction Techniques
+The dynamics sim-to-real gap is the mismatch between the transition function $P(\mathbf{s}^{t}|\mathbf{s}^{t-1},\mathbf{a}^{t-1})$ of the simulation versus that of the real world.
+It can be due to the mismatch in the physical properties of objects and motors, or by things that are not properly simulated, such as friction and backlash.
+This difference in dynamics can cause policies trained purely in simulation to fail when deployed in the real world, especially if the task requires accurate motor control, such as grasping a small object.
+We use the following key techniques to address the dynamics gap.
+Arm PID Control to Minimize Tracking Errors.
 An arm joint tracking error $\mathbf{e}_{\textrm{arm}}$ is the positional difference between the arm command of the previous timestep and the joint positions of the next timestep,
 $\mathbf{e}^{t}_{\textrm{arm}}=\mathbf{s}^{t}_{\textrm{arm}}-\mathbf{a}^{t-1}_{\textrm{arm}}.$
-Intuitively, we know that if the tracking error $\mathbf{e}_{\textrm{arm}}$ is equivalent between sim and real for all encounterable states, then there is essentially no dynamics gap.
+Prior works [16] have used PD control for the arm joints.
+PD control needs to have large tracking errors, the difference between the target and the current joint positions, to compensate for lifting the arm and any grasped object against gravity.
+Intuitively, we know that if the tracking error $\mathbf{e}_{\textrm{arm}}$ is equivalent between sim and real for all visited states, then there is essentially no dynamics gap.
 With this, one solution would be to match the tracking errors as closely as possible through accurate embodiment modeling with low-level PD control tuning.
-In practice, achieving this alignment for the entire state distribution proves to be nontrivial given both the nonlinearity of manipulator dynamics in conjunction with numerous unmodeled effects such as wear-and-tear, motor backlash, etc.
+In practice, achieving this alignment for the entire state distribution requires nontrivial work, given the nonlinearity and drifting of motor dynamics due to effects such as motor backlash, wear-and-tear, etc.
 Rather than matching non-zero tracking errors, another approach is to simply achieve near-zero tracking errors in both sim and real.
 In other words, whatever arm command is outputted, the resulting arm position should be roughly achieved $\mathbf{s}^{t}_{\textrm{arm}}\approx\mathbf{a}^{t-1}_{\textrm{arm}}$.
-In fact, prior studies have shown that tracking error minimization is one of the most reliable strategies for enabling successful sim2real transfer [3].
+In fact, prior studies have shown that tracking error minimization is one of the most reliable strategies for enabling successful sim-to-real transfer [3].
 To do so, we use a PID joint position controller for our manipulator in both simulation and in real, with a small joint position change limit $z$ (Section III-A) to ensure minimal tracking error for a wide variety of pose configurations.
-Though the leg joints also possess a dynamics mismatch, we note that policy’s task success is minimally impacted by sim-to-real locomotion differences.
-Therefore, we opt to simply use PD position control for the legs.
-D-B More Vision Gap Reduction Techniques
+Our tuned PID controller executes smoother and slower trajectories than those from the PD controller, allowing safe and accurate deployment.
+We stay with PD control for the quadruped motors as the hardware driver does not support PID.
+Arm Control Perturbation.
+We add ±$\pm 0.02$ noise to the arm joint position targets uniformly randomly, sampled every 50 high-level control steps or 5 seconds real time.
+This noise is applied to both teacher and student training.
+The noise level is increased to ±$\pm 0.05$ after 3M iterations for teacher training.
+This is to make the learned policy to be robust w.r.t.
+different sources of control noise arising when transferred to real.
+Arm Mount Perturbation.
+We randomly perturb the arm mount position and yaw at the beginning of each episode to make the learned policy robust to sim-to-real gaps that may arise from the actual mounting accuracy, torso height variations, etc.
+In this work, we sample perturbations for $x$, $y$, $z$ and yaw $\theta$ from the following ranges: $x:\pm 1$cm, $y:\pm 1$cm, $z:\pm 1$cm, $\theta:\pm 2\times 10^{-3}$ rad.
+Object Perturbations.
+Due to the relatively deterministic nature of simulations (in contrast to the high-entropy real world), we observed that policies often converge to deterministic strategies.
+For example, for picking up an object, a policy may learn to always position itself so that the object has the same relative position to its base.
+This deterministic “memorizing” behavior can result in task failure if the conditions for success are even slightly off.
+Therefore, we enforce learning robust, reactive policies by perturbing objects of interest during various stages of task execution.
+Empirically, this helps with both the exploration (as the probability of stuck in a local state is reduced) and the robustness of the policy (as the state coverage is enlarged).
+We perturb the object within circle with a radius of $r_{\rm perturb}$ centered around the current position of the object.
+We randomly sample a position within this circle and check its eligibility.
+If it touches or within the space of robot torso projected on to the ground plane, then this perturbation position is regarded as invalid.
+Otherwise it is regarded as valid.
+If a valid perturbation is found within 10 trails, then we transport the object to the new location as a form of object perturbation.
+Otherwise, we keep the position of the object unchanged.
+In this work, $r_{\rm perturb}\!=\!10$cm is used.
+Stationary Manipulation.
+We make the quadruped stand as still as possible when manipulating objects, i.e., during grasp or drop into.
+This avoids the large sim-to-real gap of a walking and shaking dog when accurate manipulation is needed.
+To achieve this, when training low-level policies to follow locomotion commands, we also encourage the quadruped to enter a stationary mode with a balanced stance when the speed command is close to zero.
+Without stationary manipulation, not only will success rates suffer, but also it becomes unsafe to deploy, because the gripper fingers can hit the ground, damaging fingers, arm motors and the surrounding.
+Reducing Quadruped Shaking
+When transitioning from locomotion to stationary, the quadruped after adding the arm on top can shake, sometimes quite violently, before settling.
+This only happens in real deployment, not in simulation.
+Our experiments seem to show that hiding the arm command and observations from the low-level locomotion policy during training and deployment reduces most of the shaking.
+Shaking when transitioning to stationary is still present in most low-level policies.
+For the experiments in this paper, we ran low level training for ten random seeds, and chose three of the best ones (less shaking and more stable stationary pose) to continue high level policy training and evaluation.
+Dynamics Sim-to-Real Gaps Not Addressed.
+Unlike prior work, we did not explicitly model the motor dynamics, and thus the effects of backlash and motor operating near torque limit boundaries are not modeled.
+For the leg motors of Go1, we follow existing work which uses P control plus damping to mimic the physical PD controller.
+Damping is applied on the absolute joint velocity of the motor, while the D factor in PD control is applied on the difference between the targeted and the actual joint velocities.
+During fast leg joint movements, this difference can cause a large sim-to-real gap.
+E-B Vision Gap Reduction Techniques
+The second source of sim-to-real gap for image-conditioned policies is the vision gap.
+This gap is caused by a visual distribution between simulated and real world pixels.
+Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios.
+Accordingly, we apply various vision sim-to-real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
+Image Domain Randomization.
+Inspired by SECANT [12], we apply pixel-level perturbations to every RGB image to increase the robustness of the perception model and the chance of its successful sim-to-real transfer.
+Specifically, we will randomly select a transformation from the list (Gaussian Noise, SaltPepper Noise, Speckle Noise, Random Brightness, Gaussian Blur) and apply it to each RGB image.
+We choose not to introduce higher-level image randomization/generation [60] because it might impact image semantics.
+Random Textures.
+We spawn the robot on a flat ground plane and randomize the ground texture.
+The texture can be arbitrary and generally has no specific semantics, as long as they provide enough visual distraction to the student.
+This distraction will force the student to learn to focus on the most important visual features that are related to the task at hand: objects with reasonable sizes and target colors.
+When deployed in real, we hope that this visual focus will get transferred and help the robot locate objects.
+Random Spatial Augmentation.
+To make the visual representation module robust, we incorporate spatial augmentations in the visual representation learning [57, 58].
+Random Background Objects.
+To further increase the visual complexity of the surrounding environment, we randomly sample and spawn a set of background objects in the room around the robot’s initial position (Figure A.12).
+There are two categories of background objects: 1) primitive shapes (cuboid, cylinder, and sphere) and 2) Google scanned objects [10].
+We make sure that these objects only serve the purpose of distraction and the student will not confuse them with task objects.
+It could collide with an background object during training with some penalty reward.
+When spawning an object, we also randomize its size, orientation, and texture (only for primitive shapes).
+Refer to caption Refer to caption
+Figure A.12: Example Training Scene with Randomized Objects. Left: isometric view; Right: birds-eye view.
 No Camera Depth Data.
@@
 These colors are sampled from real objects in different lighting conditions and are randomized in the HSV space during simulation training.
-Color samples are shown in Figure A.12.
+To obtain the randomization range, we randomly sample 1000 pixels for each color, as shown in Fig.
+A.13.
+Afterwards, we compute the means and standard deviations of the samples.
+HSV values are then sampled during training from a uniform distribution with a matching mean and $\pm$ bounds that cover the real-world distribution.
 Figure A.13: Color samples. Each color had 1000 randomly sampled pixels used to estimate the distribution in simulation.
-Appendix E Robot Implementation Details
-E-A Control Latency Management
-Figure A.13: Asynchronous System Architecture.
-Note that “Onboard*” refers to the low-level controller frequency being unknown due to the use of closed-source software.
-Furthermore, when using the WidowX’s PID controller, we also take advantage of Dynamixel’s onboard profiling for smooth control.
+Appendix F Robot Implementation Details
+F-A Control Latency Management
 We measured the latency along the full control cycle, including motor sensor reading ($\sim$3 ms), camera observation (30$\sim$40 ms), network connection ($\sim$0.2 ms with cable and $\sim$5 ms on Wi-Fi), and model inference (20$\sim$40 ms for the high level  [^7] Later with TensorRT, we were able to reduce this latency to around 10 ms.
@@
 Low-level control delay is sampled from $(5,7)$ ms uniformly per step.
-About the high-level inference delay, we evaluated the policies trained without high-level inference delay and compared the performances with and without test time high-level inference delays.
+  4. 4.
+    High-level inference delay: we trained policies without high-level inference delay turned on, and compared the performance under different extra test-time high-level inference delays.
 We did not find clear performance differences in terms of the success rate.
@@
 Meanwhile, it is an option that is provided in our implementation and can be turned on whenever necessary.
-Figure A.14: The architecture of the student’s U-Net for predicting segmentation and depth maps.
-Figure A.15: The architecture of the student’s policy input encoder that generates a compact observation embedding to be used by RL.
-E-B Safety Measures for the Physical Robot
-To reduce damage to the robot itself and the surroundings, we employ many strategies.
+Figure A.14: Asynchronous System Architecture.
+Note that “Onboard*” refers to the low-level controller frequency being unknown due to the use of closed-source software.
+Furthermore, when using the WidowX’s PID controller, we also take advantage of Dynamixel’s onboard profiling for smooth control.
+F-B Safety Measures for the Physical Robot
+To reduce damages to surroundings and the robot itself, we employed the following safety measures.
   1. 1.
-    Bridging the sim-to-real gap is one important part of it, so we can have confidence after observing the robot’s behavior in simulation.
+    We first addressed all sim-to-real gaps of the arm alone by training and deploying a tabletop policy.
+Similarly, we trained and deployed a standalone quadruped policy, without the mounted arm.
+Only after the two standalone policies were transferred well, we started training and deploying the whole-body policy.
+This helped us avoid potential catastrophic failures from the first-time deployment of a complex quadruped manipulation system.
   2. 2.
-    PID arm control, with reasonable action clipping, results in much smoother and more reliable arm motion, and reduces damage to the arm motors significantly.
+    We used PID control for the arm (Section E-A), with a reasonably small action clipping, resulting in much smoother and more reliable arm movement.
+This strategy reduced damages to the arm motor gears significantly.
   3. 3.
-    The randomization of motor strength, payload, center of mass, arm mounting position, etc.
-allows the policy to adapt to the real world more easily, resulting in very accurate grasping execution, so the arm does not push into the ground or the target objects, which can lead to motor damage.
+    The randomizations of motor strength, payload, center of mass, arm mounting position, etc.
+(Section E-A), though making simulation training more difficult, actually allowed the policy to adapt to the real world more easily, resulting in very accurate grasping executions.
+Thus the arm did not push into the ground or the target objects due to dynamics mismatch, avoiding motor damages.
   4. 4.
-    Control delay caused by policy inference time and communication delay leads to poor low level control.
-Go1 can topple over, and damage components mounted on top.
-For example, high level inference can sometimes take over 40ms on the laptop.
-We use asynchronous inference of high and low level to make sure high level inference does not prevent the robot to receive low level commands.
-We find that the latency of the Go1 Wi-Fi module can spike to 100ms at times, and decide to rely on cable connection instead.
-We also allow terminating deployment with a button press.
-When latency still spikes, e.g.
-when the deployment machine suddenly freezes or network cable disconnects, our driver onboard Go1 automatically stops the quadruped’s motion.
-This stopping is done using a scripted policy, and is not always safe, but it prevents damage to the hardware in most cases.
+    As high-level inference can sometimes take over 40ms to finish, we adopted asynchronous low-high inference (Fig.
+A.14) to make sure that the high-level inference did not prevent the robot from receiving closed-loop low-level commands.
+Otherwise, the quadruped could topple over and damage components mounted on top.
   5. 5.
-    Additionally, to prevent the robot from wandering off and bumping into the surroundings, during training, we incorporate a rotational search behavior as explained in Section A-A during policy training, for all baseline methods and the proposed SLIM method.
-Appendix F Real-World Experiment Details
-F-A Containers and Graspable Objects
-We use cubes of different colors as graspable objects, and baskets of different colors as the target container for dropping into, as show in Figure A.16.
+    We also enabled deployment termination with a single button press.
+When latency spikes, especially when the deployment machine suddenly freezes or the network disconnects, our driver on Go1 will automatically stop the quadruped.
+This preventative stop was done using a scripted policy and it prevented damages to the hardware in most cases.
+  6. 6.
+    To prevent the robot from wandering off and bumping into surroundings, during training we incorporated a rotational search behavior as explained in Section A-A.
+This was done for all baseline methods and SLIM.
+  7. 7.
+    We attached a leash to the quadruped to prevent it falling down due to any reason. Later when the policy became more reliable in real, we removed the leash.
+  8. 8.
+    Our 3D-printed camera mount (Fig. 6) was designed to break easily, sacrificing itself to protect the camera in the event of a side fall.
+  9. 9.
+    During deployment, we frequently monitored the temperature of Go1’s motors and took a break if necessary. A very hot Go1 motor will result in unpredictable, dangerous locomotion behaviors.
+Appendix G Real-World Experiment Details
+G-A Containers and Graspable Objects
+We use cubes of different colors as graspable objects, and baskets of different colors as the target container for dropping into, as show in Fig.
+A.15.
 The size of the basket is about $16{\rm cm}\!\times\!16{\rm cm}\!\times\!16{\rm cm}$.
@@
 Figure A.15: Containers and Graspable Objects.
-F-B Real-world Evaluation Protocols
-We detail the protocols used for real-world quantitative evaluation in the main paper (Table I).
-TABLE VI: Scene Setups for Real-World Quantitative Evaluations.
+G-B Real-world Evaluation Protocols
+We detail the protocols used for real-world quantitative evaluation in the main paper (Table III).
+TABLE IX: Scene Setups for Real-World Quantitative Evaluations.
 Legends: $\uparrow$ denotes the position robot with its orientation aligned with the direction of $\uparrow$.
 Uppercase letters denote baskets and lower case letters denote the objects to be grasped.
-The color of an object is denoted both by the background color and the character (e.g.
-g denotes the green cube and R denotes the red basket.)
+The color of an object is denoted both by the background color and the character (e.g., g denotes the green cube and R denotes the red basket.)
 Scene setup Instruction: ”Drop x into Y” (x $\mapsto$ Y)
@@
   3. 3.
-    To ensure the repeatability of the scene setup, we use the Standard spatial layout as shown in Figure 8, as it is easy to ensure accuracy of the scene setup in terms of spatial positions.
+    To ensure the repeatability of the scene setup, we use the Standard spatial layout as shown in Fig.
+7, as it is easy to ensure accuracy of the scene setup in terms of spatial positions.
 In practice, we mark the positions for the robot at the center of the square.
@@
   4. 4.
-    We follow a scene initialization procedure as detailed in Table VI across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.
-F-C Visualization Predicted Segmentation and Depth Maps
+    We follow a scene initialization procedure as detailed in Table IX across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.
+Appendix H Visualizations
+[…]
+Figure A.16: Policy Trajectory Visualization.
+For visualization purpose, we removed background objects.
+We also visualize the EE-trajectories.
+Trajectories corresponding to different subtasks are rendered with different colors.
+The ego-centric camera view is shown on the top-right of each image.
+The red colored arrows denote the direction of movement.
 […]
@@
 From left to right, the columns represent the RGB frame, the predicted depth map, the predicted segmentation mask, and our added note for better interpretation of different stages.
-In Figure A.17, we show the robot’s prediction results for depth and object segmentation, during an arbitrary outdoor deployment session.
+H-A Visualization of Policy Behaviors
+The visualization of the policy behavior in real is shown in Fig.
+1.
+We also provide demo videos in the supplementary file.
+Here we further visualize the policy behavior across the complete long-horizon task with the end-effector (EE) trajectory being rendered.
+To do this, we run the policy in the simulator, and render the images from a fixed top-down perspective and overlay the EE-trajectory.
+The EE-trajectory is color-coded in such a way that the segment corresponding to different subtask is rendered with a different color.
+The behavior of the policy is shown in Fig.
+A.16.
+To make the visualization clear we did not add background objects in the simulation for this episode.
+The robot start with a randomized position and yaw within a region (Start phase).
+A red cube and a green cube as well as a red basket and and blue basket are used.
+Their positions are randomized around the Standard layout and their yaw angles are also randomized.
+Given an instruction (“Drop the yellow cube into the red basket.”), the robot will first search for the target graspable object (the yellow cube).
+Note that during the Search phase, the robot actively moves both its body and the arm in searching of the target object (c.f.
+Fig.
+A.16-Search).
+Once the target object is in view, the robot will move towards it (Fig.
+A.16-MoveTo).
+Once the robot moves close to the target object, it will start to grasp (Fig.
+A.16-Grasp).
+During this stage, the robot body remains stationary and the arm will move towards the target cube and grasp it.
+After picking up the target cube, the robot will search for the target basket while keeping the cube in gripper.
+As shown in Fig.
+A.16-SearchWithObj, the robot rotates around its current position in searching of the target basket (as indicated by the circular ee-trajectory).
+At the same time, the arm retracts from the initial out-reaching pose after grasping towards a neutral arm pose, as indicated by the shrinking radius of the circular EE-trajectory in Fig.
+A.16-SearchWithObj.
+Once the target basket is in view, the robot will move towards it (Fig.
+A.16-MoveToWithObj) and drop the cube in gripper into that basket (Fig.
+A.16-DropInto).
+H-B Visualization of Predicted Segmentation and Depth Maps
+In Fig.
+A.17, we show the robot’s prediction results for depth and object segmentation, during an arbitrary outdoor deployment session.
 We emphasize again that the robot was deployed in this zero-shot scenario never modeled by our simulator, and no model finetuning was performed.
@@
 Overall, we find that in practice, this high-quality visual information bottleneck not only greatly improves the generalization ability of our RL policy, but also facilitates our debugging of the entire robotic system as it separates the errors of the perception model from those of the policy.
+Appendix I Human Teleoperation
+We list failure modes of the Human Teleop baseline to give a more complete picture of how difficult our task is, and to show that SLIM’s performance is non-trivial.
+  1. 1.
+    The most common failure mode for teleoperation is task timeout.
+This is sometimes due to not fully grasping the cube and the cube slipping out of the fingers during movement, leaving no time for searching and grasping again.
+Failing to grasp fully is in turn due to either stopping a bit far from the cube, or incorrectly assessing the distance to the cube under monocular vision.
+  2. 2.
+    The arm can have a twisted pose due to IK near singularities.
+This pose either overloads the arm motors, or if not, flips the camera upside down, making the operator difficult to perceive the environment.
+Overall, human teleoperation is less accurate and slower than an RL policy, due to the low-level policy’s shakiness, the difficulty of simultaneously moving the end effector during locomotion, and sometimes human errors resulting from lost focus after a long time.
+Besides being less time efficient, human teleoperation seems to be less energy efficient than our RL policies as well.
+While the RL trained high-level policy can run smoothly on a particular low-level policy, human teleoperation on the same low-level policy sometimes triggers over-current protection of Go1.
+This often happens when an over-extended arm is raised, causing Go1 to lose control of its leg motors.
+In evaluation, we ignored this Go1 hardware failure, and let the human operator simply try from start again.
+Appendix J Long-Horizon Task Learning
+J-A Long-Horizon Task with Bottleneck States
+The challenges of long-horizon task for exploration and learning have been discussed in Section IV.
+Here we want to further highlight that the existence of bottleneck states, which is orthogonal to long horizon, further exacerbates the difficulties.
+[…]
+Figure A.18: Long-Horizon Task with Bottleneck States. Left: Top-down view of the end effector trajectories. Right: The same set of trajectories in 3D space.
+A bottleneck state is a milestone that must be accomplished before reaching the final task success.
+Therefore, given two tasks of the same horizon length, the one with the presence of bottleneck states will typically be more difficult than the one without.
+To visualizing the bottleneck states for the tasks considered in this work, we use spatial proximity of a key position (end effector/EE position) as a proxy of state proximity.
+Therefore, positions that are spatially close to each other can be roughly interpreted as with similar states in terms of accomplishing the mobile manipulation task under consideration.
+Under this setting, for the same task, if there is a common segment along the way with a compact spatial support, it can be regarded as a bottleneck state.
+We visualize the trajectories from three episodes in Fig.
+A.18, given a fixed object layout and different starting positions.
+As we can see from the figure, although with quite different starting positions, all the trajectories converge and intersect at the same location marked by b in the figure, corresponding to the state for grasping the target object.
+To achieve the eventual full task success (drop the target object into basket), a successful grasping of the target object has to be accomplished as a prerequisite.
+Given that grasping small objects requires high precision control (as shown by the inserts in the left part of Fig.
+A.18), this forms a clear bottleneck state, increasing the difficulties of exploration and learning, especially for the part of state space that is only available after passing the bottleneck state.
+The existence of bottleneck states coupled with long task horizon pose a great challenges to standard algorithm, as shown by the results in the subsequent section.
+Figure A.19: Standard Method v.s.
+Progressive PEX.
+We compare the learning behaviors of Standard method (without progressive policy expansion) and the Progressive PEX approach in terms of the task progress as training progresses.
+J-B The Effectiveness of Progressive Policy Expansion
+Section IV introduces the Progressive Policy Expansion approach (Progressive PEX) for long-horizon task learning.
+The comparison of Progressive PEX and the approach using a standard network without progressive policy expansion (Standard) on solving the long-horizon task are shown in Fig.
+A.19.
+We show each policy’s ability in solving the long-horizon task throughout the training process, measured by the number of subtasks that can be solved ($y$-axis, Long Horizon Task Progress).
+We order the subtasks following their respective orders appeared in the full task.
+We show the mean and standard deviation calculated across 3 seeds.
+As can be observed from the figure, Standard approach starts with a similar trend in learning to solve subtasks as Progressive PEX.
+However, after learning to solve a few tasks, Standard appears to be saturated in terms of the learning progress, likely due to the inability of continual exploration and loss of plasticity [35, 36].
+In contrast, Progressive PEX can keep pushing the frontier on the long-horizon task progress, learning to solve the subsequent new subtask that appears along the way towards solving the full task.